### v1

# Joint Demonstration and Preference Learning Improves Policy Alignment with Human Feedback

Abstract
Aligning human preference and value is an important requirement for building contemporary foundation models and embodied AI. However, popular approaches such as reinforcement learning with human feedback (RLHF) break down the task into successive stages, such as supervised fine-tuning (SFT), reward modeling (RM), and reinforcement learning (RL), each performing one specific learning task. Such a sequential approach results in serious issues such as significant under-utilization of data and distribution mismatch between the learned reward model and generated policy, which eventually lead to poor alignment performance. We develop a single stage approach named Alignment with Integrated Human Feedback (AIHF), capable of integrating both human preference and demonstration to train reward models and the policy. The proposed approach admits a suite of efficient algorithms, which can easily reduce to, and leverage, popular alignment algorithms such as RLHF and Directly Policy Optimization (DPO), and only requires minor changes to the existing alignment pipelines. We demonstrate the efficiency of the proposed solutions with extensive experiments involving alignment problems in LLMs and robotic control problems in MuJoCo. We observe that the proposed solutions outperform the existing alignment algorithms such as RLHF and DPO by large margins, especially when the amount of high-quality preference data is relatively limited.
[…]
1 Introduction
Background. As ChatGPT has taken the world by storm, it is clear that AI systems will soon become ubiquitous in our lives. For instance, Large Language Models (LLMs) have been used to solve hard problems including video gaming (Berner et al., 2019; Mnih et al., 2015), autonomous control (Bellemare et al., 2020), and robotic manipulation (Kalashnikov et al., 2018; Kober and Peters, 2008). However, for AI systems to be able to smoothly interact with humans, and reliably help perform mission-critical tasks, it is critical that the system learns, or aligns with, the human’s preferred strategy, under specific contexts, and adapts to evolving user preferences and needs (Leike et al., 2018). Achieving this level of alignment between AI and humans requires a comprehensive approach. AI systems must be capable of not only responding to instructions but also being sensitive to the nuances of human behavior and context, ensuring that their actions align with individual values and societal norms. On the other hand, various undesirable behaviors exhibited by advanced AI systems, e.g. deception (Park et al., 2023) and manipulation (Perez et al., 2022) have raised concerns about the potential hazards from AI systems.
The “alignment” task mentioned above has already proven to be useful in a few applications, such as in ChatGPT. Despite its recent popularity, the research area is still in its infancy. There are a number of critical open challenges, e.g. 1) How to solve the alignment problem leveraging diverse human feedback to further boost alignment performance? 2) How to enable AI systems to continuously learn and adapt to time-varying preferences, changing contexts, and emerging societal norms?
Contribution. In this work, we address the first question posed above. That is, we develop formulations and efficient algorithms capable of integrating different sources of human response data to achieve the best alignment performance.
Most of the existing alignment algorithms are sequential, i.e. they separate the alignment process into a number of successive stages, each utilizing one distinctive category of data. For example, the well-known RLHF approach adopted by (Ouyang et al., 2022) breaks the alignment step into the supervised finetuning (SFT), reward modeling (RM), and reinforcement learning (RL) stages, where the SFT stage uses demonstration data, the RM stage uses the preference data, and the RL stage uses a set of available prompts. Similar strategies have been used in other related works such as (Rafailov et al., 2023; Li et al., 2023; Zhu et al., 2023; Liu et al., 2023). Sequencing of learning tasks may facilitate computation at the expense of inefficient exploitation of data. To illustrate, consider the three-stage RLHF approach proposed in (Ouyang et al., 2022), in the extreme case where the amount of high-quality preference data is quite limited, the reward model trained cannot adequately reflect the preferences of the human, which may lead to unsatisfactory performance in the RL stage. This simple example underscores the importance of ensuring the reward model is consistent with the policy model. Such consistency can not be guaranteed with sequential execution of the RM and RL tasks.
The alignment problem is a learning problem with (at least) three types of input data: the demonstration data (consists of prompts and human-generated continuations), the preference data (consists of prompts and pairs of human-ranked responses), as well as prompts without any responses. The input to the learning problem, besides these data, also includes a model pre-trained using a large corpus of text data. The output of the learning process is a fine-tuned model aligned with human preferences/values which govern the generation of demonstration and preference data. To integrate these diverse data, we propose a joint learning approach, which is informed by recent advances in Inverse Reinforcement Learning (IRL) (Arora and Doshi, 2021; Zeng et al., 2022b), stochastic choice theory (Blavatskyy and Pogrebna, 2010) and bi-level optimization (Hong et al., 2020; Ji et al., 2021; Khanduri et al., 2021). More specifically, the proposed formulation integrates SFT, RM, and RL into a single stage, so that reward modeling and policy optimization can fully leverage all three kinds of data mentioned above in alignment. Several existing alignment schemes, such as RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2023), and some of their extensions can be seen as particular instances of the proposed formulation. We observe that the proposed solutions outperform the existing alignment algorithms such as RLHF by large margins, especially when the amount of high-quality preference data is relatively limited.
2 Preliminaries and Related Work
2.1 Notation
The Finite-Horizon MDP Model. Consider a Markov decision process (MDP) defined by the tuple $(\mathcal{S},\mathcal{A},P,\rho,r,\gamma)$, which consists of the state space $\mathcal{S}$, the action space $\mathcal{A}$, the transition dynamics $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$, the initial state distribution $\rho(\cdot)$, the reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ and the discounted factor $\gamma>0$. Let $s_{t}\in\mathcal{S}$, $a_{t}\in\mathcal{A}$ denote the state and action variables at time $t$ respectively, where $a_{t}$ is generated according to some policy $\pi(\cdot|s_{t}):\mathcal{S}\to\Delta_{|\mathcal{A}|}$, where $\Delta_{|\mathcal{A}|}$ denotes the probability simplex in $\mathbb{R}^{|A|}$. Note that viewing the generation process of language models as an MDP, $s_{t}$’s and $a_{t}$’s are referred to as the given sequence and next token to be generated, respectively; $\pi(\cdot)$ then represents the language model itself. Define $\tau:=\{(s_{t},a_{t})\}_{t=1}^{T}$ as a (finite time) trajectory of state and action pairs. Let $\mathcal{H}_{T}\subset\prod_{t=0}^{T-1}\big{(}S\times A\big{)}$ denote all feasible state/action pairs.
Human Feedback Generation. Assume that the human demonstration $\tau$ is generated from a ground-truth human expert policy $\pi^{\rm E}$, i.e., for each state $s_{t}$, we have $a_{t}\sim\pi^{\rm E}(\cdot|s_{t})$. Moreover, let ‘$\tau_{l}\prec\tau_{w}$’ indicate the winner trajectory $\tau_{w}$ is preferred over the trajectory $\tau_{l}$. Further, we denote $(\tau_{l}\prec\tau_{w})\sim\pi^{\rm P}$ as the preference oracle.
Note that when we present our problem formulation, we will use the above data generation distributions so quantities such as likelihood can be properly defined. Finite-sample version of demonstration and preference data will be introduced subsequently to facilitate algorithm design.
Since reward modeling is a critical task for achieving state-of-the-art alignment performance, below we review a few popular paradigms for reward learning.
2.2 Reward Learning from Preference Data
RLHF is a popular technique for finetuning AI systems to align with human preferences and values. The RLHF approach proposed in (Stiennon et al., 2020; Ouyang et al., 2022) consists of the following three-stage: 1) the supervised fine-tuning (SFT) stage, where the demonstration data is used to fine-tune the model in a supervised manner; 2) the reward modeling (RM) stage, where the preference data is used to train a reward model; 3) the reinforcement learning (RL) stage, where the SFT model is further finetuned by running RL using the trained reward model.
Furthermore, existing work such as Direct Preference optimization (DPO) (Rafailov et al., 2023) and Inverse Preference Learning (IPL) (Hejna and Sadigh, 2023) both remove the need for explicit reward modeling, and they directly extract the policy from preferences. This greatly reduces the training complexity, but it has been observed that these algorithms can be unstable in the training process (Azar et al., 2023; Xu et al., 2024). There is also a large number of works that aim to learn reward functions from rating (Daniel et al., 2014) or ranking (Yuan et al., 2023; Myers et al., 2022). However, these works often fail when the quality of the preference dataset is low Morimura et al. (2024). In our approach, the inclusion of additional demonstrations can be leveraged to counteract the adverse effects of limited preference quality.
2.3 Reward Learning using Demonstration Data
In the RL literature, a line of work referred to as Inverse Reinforcement Learning (IRL) proposes to jointly learn the reward and policy from expert demonstration data. Specifically, the target is to find the parameterized reward function $r(s,a;\theta)$ (resp. an optimal polity $\pi^{*}(s,a)$) that best explains (resp. mimics) an expert policy $\pi^{E}$ given the demonstration data $\mathcal{D}$. For example, the well-known maximum entropy IRL (MaxEnt-IRL) framework (Ziebart, 2010; Ziebart et al., 2013; Bloem and Bambos, 2014; Zhou et al., 2017) finds a policy maximizing entropy subject to the expected features that match the empirical averages in the expert’s observation dataset. However, this approach can only be used to model linear rewards.
Subsequent works such as (Levine et al., 2011; Wulfmeier et al., 2015; Zeng et al., 2022b) further improve the MaxEnt-IRL method so that nonlinear reward can be used. For example, a recent paper (Zeng et al., 2022a) proposed a maximum likelihood IRL (ML-IRL) formulation based on the Dynamic Discrete Choice (DDC) model, and nonlinear reward function can be used. It is also shown that when the reward function is linearly parameterized, then the MaxEnt is the Lagrangian dual of the ML-IRL problem. However, it is worth mentioning that to our best knowledge, almost all IRL-based methods can only utilize demonstration data, which can be problematic because a large amount of high-quality demonstration data is typically hard to obtain. Further, it is well-known that only using demonstration cannot extract precise human preference, especially in safety-related tasks where the boundaries between permissible and impermissible actions need to be precisely determined (Liao and Sun, 2024); see, e.g., (Fischer et al., 2021) which shows that insufficient demonstration dataset could lead to high generalization error.
2.4 Reward Learning from Diverse Data
For a more precise alignment of agent behavior with human intention, it is crucial to leverage diverse data source such as both expert demonstrations and human feedback to achieve alignment. In (Ibarz et al., 2018), the authors first combine two approaches to learn from human feedback: expert demonstrations and trajectory preferences. The addition of demonstrations to learning from preferences typically results in substantial performance gains compared with using either demonstrations or preferences in isolation. In (Palan et al., 2019) and (Bıyık et al., 2022), the authors integrate diverse sources of human feedback including demonstrations and pairwise comparisons in a Bayesian approach to learn reward functions that are assumed to be linear in carefully selected features and evaluate their proposed method on robot learning platform. Moreover, their proposed methods need to actively generate preference queries, which are expensive to collect in practical applications. In contrast, our proposed approach (to be presented shortly) is not Bayesian and does not include the requirement that the reward model is linear in pre-selected features. Hong et al. (2024) proposed a single-stage supervised learning algorithm ORPO that can perform supervised fine-tuning and preference alignment in one training session without maintaining multiple models simultaneously. However, Hong et al. (2024) can suffer from overoptimization due to its supervised learning pipeline Liu et al. (2024).
3 Proposed Approach: Alignment with Integrated Human Feedback (AIHF)
Figure 1: Comparison of the RLHF with the proposed AIHF.
3.1 Preliminaries
In this section, we develop our proposed formulation for the alignment problem. To flesh out our ideas, let us start with a brief discussion of the RLHF approach. First, the SFT stage maximizes the following likelihood function:
$\ell_{\rm SFT}(\pi):=\mathbb{E}_{\tau\sim\pi^{\rm E}}\Big{[}\log\prod_{t\geq 0}\Big{(}\pi(a_{t}|s_{t})\Big{)}^{\gamma^{t}}\Big{]}.$ (1)
For the RM stage, let $R(\tau;\theta):=\sum_{t\geq 0}\gamma^{t}r(s_{t},a_{t};\theta)$ denote the (parameterized) discounted cumulative reward of one trajectory $\tau$. Following the Bradley-Terry-Luce (BTL) model (Bradley and Terry, 1952), the human preference distribution for one pair of trajectories can be stipulated as (where $\sigma(\cdot)$ is the sigmoid function):
$P[\tau_{l}\prec\tau_{w}]=\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}.$ (2)
Based on (2), this stage maximizes the following:
$\ell_{\rm RM}(\theta):=\mathbb{E}_{(\tau_{l}\prec\tau_{w})\in\pi^{P}}\Big{[}\log\Big{(}\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}\Big{)}\Big{]}.$ (3)
Finally, the RL stage fine-tunes the SFT policy by solving the following policy optimization problem:
$\max_{\pi}\;\ell_{\rm RL}(\pi;R(\cdot;\theta)):=\mathbb{E}_{s_{0}\sim\rho,\tau\sim\pi}\bigg{[}R(\tau;\theta)+\sum_{t\geq 0}\gamma^{t}c(\pi)\bigg{]}$ (4)
where $c(\pi)$ is a penalty function, which can take the form of, e.g., the KL divergence $c(\pi):={D}_{\rm KL}\Big{(}\pi(\cdot|s_{t})\|\pi_{\rm SFT}(\cdot|s_{t})\Big{)}$. This penalty ensures the resulting policy is close to the original SFT policy. We have the following remark about the above problem.
Remark 3.1
(Optimal Policy for RLHF) Let $\pi_{\theta}$ denote the optimal policy for problem (4), where $\theta$ is used to parameterize the reward function. When $c(\pi):={D}_{\rm KL}\Big{(}\pi(\cdot|s_{t})\|\pi_{\rm SFT}(\cdot|s_{t})\Big{)}$, the KL-regularized MDP problem described by (4) has a closed-form solution:
$\pi_{\theta}(a|s)=\frac{\pi_{\rm SFT}(a|s)\exp(Q_{\theta}(s,a)/\beta)}{\sum_{\tilde{a}}\pi_{\rm SFT}(\tilde{a}|s)\exp(Q_{\theta}(s,a)/\beta)},$ (5)
where the corresponding value function $V_{\theta}$ and the Q-function $Q_{\theta}$ are defined as below:
$V_{\theta}(s):=\mathbb{E}_{\tau\sim\pi_{\theta}}\bigg{[}R(\tau;\theta)+\beta\sum_{t=0}^{\infty}\gamma^{t}{D}_{\rm KL}\Big{(}\pi(\cdot|s_{t})\|\pi_{\rm SFT}(\cdot|s_{t})\Big{)}\bigg{|}s_{0}=s\bigg{]}$ (6a)
$Q_{\theta}(s,a):=r(s,a;\theta)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}\big{[}V_{\theta}(s^{\prime})\big{]}.$ (6b)
Further, assuming that $T=1$, i.e., $\tau=(s_{0},a_{0})$, and considering the LLM alignment problem as a sequence-level training problem (this is a popular simplification in language models, see, e.g., Rafailov et al. (2023)), the closed-form expression of $\pi_{\theta}$ (5) can be reduced to:
$\pi_{\theta}(a_{t}|s_{t})=\frac{\pi_{\rm SFT}(a_{t}|s_{t})\exp(\frac{1}{\beta}r(a_{t},s_{t};\theta))}{\sum_{a_{t}\in A}\big{(}\pi_{\rm SFT}(a_{t}|s_{t})\exp(\frac{1}{\beta}r(a_{t},s_{t};\theta))\big{)}}.$ (7)
Remark 3.2
(Interpretation of the RLHF Policy) Let us take a closer look at (7). In this expression, the results of learning from demonstration data (i.e., $\pi_{\rm SFT}(\cdot)$) and from the preference data (i.e., $r(\cdot;\theta)$) are combined together in a product form to yield the final policy estimate. Such a form, though convenient, can result in a number of issues.
One such an issue is that, the reward $R(\tau;\theta)$ is learned only through the preference dataset $\mathcal{P}$. Despite the fact that problem (3) is relatively simple to optimize, such an approach has the following issues: 1) the high-quality demonstration data has been ignored in shaping the reward model; 2) in the situation where the preference data has limited coverage or is of low-quality, the learned reward model is likely to be of low quality as well. Here, low-quality preference data arises when, e.g., the trajectories are generated from low-quality models, or the preference annotation is noisy.
To see the impact on the noisy annotation, consider, for example, the case where a given $s_{t}$, suppose there are two actions $\hat{a}_{t}$ and $\tilde{a}_{t}$, and they are equally good. The SFT data respects the ground truth, and the learned SFT policy says $\pi_{\rm SFT}(\tilde{a}_{t}|s_{t})\approx\pi_{\rm SFT}(\hat{a}_{t}|s_{t})$. However, due to labeling noise, we have a preference data $(s_{t},\hat{a}_{t})\prec(s_{t},\tilde{a}_{t})$, therefore we have learned $R(a_{t},\tilde{a}_{t};\theta)>R(a_{t},\hat{a_{t}};\theta)$, yielding significant difference in the final policy. We refer the readers to Sec. 4 for more discussion.
3.2 A Meta-Formulation
The above discussion strongly motivates us to develop an approach that can model the entire alignment process with a common parametrization for both policy and reward models. Towards this end, let us consider the following meta-formulation:
${\bf(AIHF)}\quad\max_{\theta}$ $\quad L(\theta):=\alpha L_{1}(\pi_{\theta})+L_{2}(R(\cdot;\theta))$ (8a)
${\rm s.t.}$ $\quad\pi_{\theta}:=\arg\max_{\pi}~{}L_{3}(\pi;R(\cdot;\theta))$ (8b)
where $\theta\in\mathbb{R}^{d}$ is a parameter; $L_{1}(\pi_{\theta})$ is a measure of fit of the parameterized policy $\pi_{\theta}$ to demonstration data and $L_{2}(R(\cdot;\theta))$ is a measure of fit of the parameterized reward model $R(\cdot;\theta)$ to the preference data; $L_{3}(\pi,R(\cdot;\theta))$ is a measure of performance of policy $\pi$ with respect to reward model $R(\cdot;\theta)$; $\alpha\geq 0$ is a balancing coefficient reflecting the relative size of demonstration versus preference data. We refer the above formulation as Alignment with Integrated Human Feedback (AIHF); see Fig. 1 for an illustration of the proposed formulation. The AIHF (8) is a meta-problem that models the alignment problem. It has two levels: an upper-level problem in which the goal is to find policy and reward models that jointly maximize a measure of fit to demonstrations and preference datasets and a lower-level problem which ensures that the policy model optimizes performance with respect to the reward model. Its components can be customized to yield specific alignment formulations and algorithms. But before diving into various customizations, let us discuss the advantages of this formulation.
Generality. One can specialize the choice of loss functions and problem parameters, to yield a number of existing alignment formulations. Such generality implies that algorithms developed for (8) are easily applicable to different special formulations it covers. For more details see Sec. 3.4.
Joint optimization. The formulation jointly optimizes the reward and the policy. One benefit here is that it can strengthen the reward model through integrating both demonstrations and pairwise comparisons. Compared with the standard RLHF pipeline, through integrating additional data source such as demonstrations to train the reward model, it can further boost the policy optimization subroutine to achieve better alignment performance. See Sec. 5 for a detailed discussion on how the reward parameter $\theta$ is updated by leveraging such demonstration.
Dataset Integration. Clearly, the reward learning process leverages all the available data, therefore, we can expect that a high-quality reward model can still be obtained even under unfavorable situations where certain category of data is scarce or of low quality.
3.3 Specification of AIHF
In this section, we specify the formulation (8). Let us begin with the choice of $L_{1}$. It can be directly instantiated by using an objective similar to (1), which is the likelihood function over the collected expert demonstrations. Note that we aim to optimize the reward parameter $\theta$ to align with human feedback in (8a), thus the objective of $L_{1}$ can be specialized as a maximum likelihood function over expert demonstrations as below:
$L_{1}(\pi_{\theta}):=\mathbb{E}_{\tau\sim\pi^{\rm E}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\log\pi_{\theta}(a_{t}|s_{t})\Big{]}.$ (9)
Here $\pi_{\theta}$ optimizes the measure of performance $L_{3}(\pi;R(\cdot;\theta))$ for a reward model $R(\cdot;\theta)$ as below:
$L_{3}(\pi;R(\cdot;\theta)):=\mathbb{E}_{s_{0}\sim\rho,\tau\sim\pi}\bigg{[}R(\tau;\theta)-\beta\sum_{t\geq 0}\gamma^{t}{D}_{\rm KL}\Big{(}\pi(\cdot|s_{t})\|\pi^{0}(\cdot|s_{t})\Big{)}\bigg{]}$ (10)
where $\pi^{0}$ is some initial policy and $\beta>0$ is temperature parameter.
Next, we specify $L_{2}$. To ensure internal model consistency, we identify the likelihood function for preference data so it is in accordance with the preferences implied by the reward model $R(\cdot;\theta)$ used in the definitions of $L_{1}$ and $L_{3}$. Thus, the optimal distribution $\mu_{\theta}$ over the set of $T$-long sequence of state-action pairs is defined as follows:
$\mu_{\theta}:=\arg\max_{\mu\in\Delta^{T}}\mathbb{E}_{\tau\sim\mu}\big{[}R(\tau;\theta)-\beta\mathcal{D}_{KL}(\mu||\mu^{0})\big{]}$
where $\Delta_{T}$ denotes the simplex on $\mathcal{H}_{T}$ and $\mu^{0}$ is a prior distribution on the trajectories. It can be shown the solution of the above problem is of the form:
$\mu_{\theta}(\tau)=\frac{\mu^{0}(\tau)\exp\big{(}R(\tau;\theta)/\beta\big{)}}{\sum_{\tau^{\prime}\in\mathcal{H}_{T}}\mu^{0}(\tau^{\prime})\exp\big{(}R(\tau^{\prime};\theta)/\beta\big{)}\big{)}}.$
With this result, we can now obtain a model for the likelihood that sequence $\tau_{j}$ is preferred over $\tau_{j}$. By the independence of irrelevant alternatives property (Fudenberg et al., 2015) of the optimal choice $\mu_{\theta}$, when the set of feasible choices is reduced from $\mathcal{H}_{T}$ to just the the two-tuple $\{\tau_{l},\tau_{w}\}$, the likelihood that sequence $\tau_{w}$ is preferred over $\tau_{l}$ is given by $\mathbb{P}_{\theta}(\tau_{l}\prec\tau_{w}):=\frac{\mu_{\theta}(\tau_{w})}{\mu_{\theta}(\tau_{l})+\mu_{\theta}(\tau_{w})}.$ This motivates the choice of $L_{2}(\theta)$ as the following likelihood function:
$L_{2}(R(\cdot;\theta))=\mathbb{E}_{(\tau_{l}\prec\tau_{w})\in\pi^{P}}\Big{[}\log\frac{\mu_{\theta}(\tau_{w})}{\mu_{\theta}(\tau_{w})+\mu_{\theta}(\tau_{l})}\Big{]}=\mathbb{E}_{(\tau_{l}\prec\tau_{w})\in\pi^{P}}\Big{[}\log\frac{\mu^{0}(\tau_{w})\exp\big{(}R(\tau_{w};\theta)\big{)}}{\mu^{0}(\tau_{w})\exp\big{(}R(\tau_{w};\theta)\big{)}+\mu^{0}(\tau_{l})\exp\big{(}R(\tau_{l};\theta)\big{)}}\Big{]}.$
With $\mu^{0}$ equal to the uniform distribution on $\mathcal{H}_{T}$, this model is equivalent to the BTL model (3):
${L}^{\rm BTL}_{2}(\theta)=\ell_{\rm RM}(\theta)=\mathbb{E}_{(\tau_{l}\prec\tau_{w})\in\pi^{P}}\Big{[}\log\Big{(}\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}\Big{)}\Big{]}.$ (11)
3.4 Special Cases of AIHF
Next we discuss how formulation (8) can be specialized to some of the known alignment algorithms.
Specialization to RLHF-Type Approach. First, if we set the coefficient $\alpha=0$ in (8), we obtain the following problem:
$\begin{split}\max_{\theta}&\quad L_{2}(\theta)\;\;{\rm s.t.}\;\;\pi_{\theta}:=\arg\max_{\pi}L_{3}(\pi;R(\cdot;\theta)).\end{split}$ (12)
We note that now the upper- and lower-level problems are completely decomposable, since the upper-level problem solves for the reward parameterization $\theta$, while the lower-level problem solves for the policy (for the given reward), yielding two separate problems, which are exactly the RM and the RL problems in the typical RLHF approach.
Specialization to DPO-Type Approach. Let us consider the relationship between our formulation (8) with the DPO-type approaches. Let us set the following objective function $L_{1}=\ell_{\rm SFT}$ and $L_{2}=\ell_{\rm RM}$, and assume that $T=1$ for th generation process. Through relaxing the constraint which ensures the policy is optimal w.r.t. a certain parameterized model, we can obtain a DPO-type formulation:
$\max_{\pi}~{}L(\pi):=\alpha\cdot\mathbb{E}_{\tau^{\rm E}\sim\pi^{\rm E}}\bigg{[}\log\pi(a^{\rm E}|s^{\rm E})\bigg{]}+\mathbb{E}_{(\tau_{i}\prec\tau_{j})\sim\pi^{P}}\bigg{[}\sigma\big{(}\beta\log\frac{\pi(a_{j}|s_{j})}{\pi^{0}(a_{i}|s_{j})}-\beta\log\frac{\pi(a_{i}|s_{i})}{\pi^{0}(a_{j}|s_{i})}\big{)}\bigg{]}.$ (13)
The above formulation specializes to Liu et al. (2024), which is a slightly generalized version of DPO when both demonstration and preference data are used. Setting $\alpha=0$ reduces to the problem solved by DPO; see Rafailov et al. (2023), Eq. (2).
4 Why AIHF can outperform two-stage alignment approaches?
To understand the difference between the proposed approach and the two-stage approach of the standard alignment pipeline, let us consider the following thought process over a simplified, but informative setting.
Let us consider the simplified case where there are $N$ discrete actions, that is, the action set is $A:=\{\tau_{1},\tau_{2},\cdots\tau_{N}\}$. Assume that there is a reward function $r(\cdot):A\mapsto\mathbb{R}$ (note, that here the parameterization of the reward has been omitted for simplicity). Let us compare the optimal solutions for different alignment algorithms and get some intuition about each approach. Throughout the process, we will assume that the reward function is not parameterized, and we directly learn the optimal reward value. Indeed this is a simplification needed to understand the structure of the solution, but we argue that this is without loss of generality, since if the reward model is sufficiently parameterized by neural networks, then these optimal reward values can always be recovered by proper training.
Consider the following decision-making model:
${\pi^{*}:=\arg\max_{\pi\in\Delta^{N}}\mathbb{E}_{\tau\sim\pi(\cdot)}\Big{[}r(\tau)-\beta\log\pi(\tau)\Big{]},}$ (14)
where $\Delta^{N}$ is the probability simplex and $\beta>0$. It can be shown that the solution to (14) is:
$\pi^{*}_{i}(r)=\frac{\exp(\frac{r_{i}}{\beta})}{\sum_{j}\exp(\frac{r_{j}}{\beta})},\quad\mbox{\rm where}\quad r_{i}:=r(\tau_{i}),\quad\forall~{}i.$ (15)
Reward Estimation via Demonstration Data: When demonstration data $\mathcal{D}:=\{\tau_{n}\}$ is available where each $\tau_{n}\in A,\;\forall~{}n$, the estimation problem can be written as:
$\widehat{r}_{\mathcal{D}}:=\arg\max_{r}\ell_{\mathcal{D}}(r)$
where $\ell_{\mathcal{D}}(r):=\mathbb{E}_{\tau_{i}\sim\mathcal{D}}[\log\pi^{*}_{i}(r)]$. By (15) it follows that
$\ell_{\mathcal{D}}(r)=\frac{1}{\beta}\mathbb{E}_{\tau_{i}\sim\mathcal{D}}\big{[}r_{i}\big{]}-\log\sum_{j}\exp(\frac{r_{j}}{\beta}).$
Hence, it is easy to verify the gradient is:
$\frac{\partial\ell_{\mathcal{D}}(r)}{\partial r_{i}}$ $=\frac{1}{\beta}\Big{(}\frac{\#\tau_{i}~{}\mbox{in}~{}\mathcal{D}}{|\mathcal{D}|}-\pi^{*}_{i}(r)\Big{)},~{}\forall~{}i.$ (16)
and the second order derivatives are:
$\frac{\partial^{2}\ell_{\mathcal{D}}(r)}{\partial r_{i}^{2}}=-\frac{1}{\beta^{2}}\pi^{*}_{i}(r)(1-\pi^{*}_{i}(r))$ $~{}~{}~{}~{}~{}\frac{\partial^{2}\ell_{\mathcal{D}}(r)}{\partial r_{i}\partial r_{j}}=\frac{1}{\beta^{2}}\pi^{*}_{i}(r)\pi^{*}_{j}(r).$ (17)
Therefore, we can observe that setting the gradient to zero, the first-order stationary solution (in fact, in this case it is optimal solution since the Hessian is negative definite) indicates that the estimate of the probability of choosing action $i$ is equal to the observed frequency of action $\tau_{i}$ appears in $\mathcal{D}$.
Note that in our simplified case, the above estimator can be viewed as the counterpart for the SFT estimator since it is the best that one can get using only demonstration data.
Reward Estimation with Preferences Data: Suppose that the preference data is available, $\mathcal{P}:=\{(\tau_{i}\prec\tau_{j})\}$, where again $\tau_{i},\tau_{j}\in A$. Let us understand how to analyze the optimal policy $\pi^{*}$. By IIA (independence of irrelevant alternatives) (Fudenberg et al., 2015), when the decision maker is asked to choose between only two alternatives $\{\tau_{i},\tau_{j}\}$, we have :
$\Pr(\tau_{j}\prec\tau_{i})$ $=\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}.$
The reward estimation problem is defined as:
$\widehat{r}_{\mathcal{P}}\in\arg\max_{r}\ell_{\mathcal{P}}(r)$ (18)
where $\ell_{\mathcal{P}}(r):=\mathbb{E}_{(\tau_{j}\prec\tau_{i})\sim\mathcal{P}}\Big{[}\log\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\Big{]}$. The partial derivative of the log-likelihood is given by:
$\frac{\partial\ell_{\mathcal{P}}(r)}{\partial r_{i}}$ $=\frac{1}{\beta}\sum_{j\neq i}\big{(}\frac{\#\{\tau_{j}\prec\tau_{i}~{}\mbox{in}~{}\mathcal{P}\}}{|\mathcal{P}_{i,j}|}-\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\big{)}\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}.$ (19)
The first-order condition can be written as:
$\pi^{*}_{i}(r)$ $=\frac{\sum_{j\neq i}\{\#a_{i}\succ a_{j}~{}\mbox{in}~{}\mathcal{P}\}}{\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho^{*}_{-(i,j)}(r)}$ (20)
where $\rho^{*}_{-(i,j)}(r):=\Big{(}1-\sum_{k\in A\backslash\{i,j\}}\pi^{*}_{k}(r)\Big{)}^{-1}$ is the expected number of times an action other than $\tau_{i}$ or $\tau_{j}$ is selected when sampling $\pi^{*}(r)$ infinitely many times. The second-order derivatives are given by
$\frac{\partial^{2}\ell_{\mathcal{P}}(r)}{\partial r_{i}^{2}}=-\frac{1}{\beta^{2}}\sum_{j\neq i}\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\frac{\pi^{*}_{j}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}$ $~{}~{}\frac{\partial^{2}_{r_{i}}\ell_{\mathcal{P}}(r)}{\partial r_{j}\partial r_{i}}=\frac{1}{\beta^{2}}\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\frac{\pi^{*}_{j}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}.$
Note that compared with the previous case with only demonstration data, here the first-order solution does not imply global optimality since the Hessian is not negative semi-definite.
Next, by using the above analysis, let us understand the policies that can be obtained by different alignment algorithms.
SFT Estimator. In this case, it is easy to observe that the optimal policy is given by
$\pi^{\rm SFT}_{i}=\frac{\#\tau_{i}~{}\mbox{in}~{}\mathcal{D}}{|\mathcal{D}|}={\pi^{*}(\widehat{r}_{\mathcal{D}})},~{}\forall~{}i.$ (21)
RLHF Policy. The RLHF policy is defined as follows:
$\pi^{\rm RLHF}=\arg\max_{\pi\in\Delta^{N}}~{}~{}\mathbb{E}_{\tau_{i}\sim\pi}\big{[}{\widehat{r}_{\mathcal{P}}}(\tau_{i})\big{]}-\beta\mathrm{KL}(\pi||\pi^{\rm SFT})$
where $\widehat{r}_{\mathcal{P}}$ is the estimator obtained from preferences (i.e. solution to (20)), $\pi^{\rm SFT}$ is a prior distribution obtained from the SFT model trained with demonstration dataset $\mathcal{D}$, and $\Delta^{N}$ is the probability simplex. It follows that
$\pi^{\rm RLHF}(\tau_{i})$ $=\frac{\pi^{\rm SFT}(\tau_{i})\exp\frac{1}{\beta}\widehat{r}_{\mathcal{P}}(\tau_{i})}{\sum_{j=1}^{n}\pi^{\rm SFT}(\tau_{j})\exp\frac{1}{\beta}\widehat{r}_{\mathcal{P}}(\tau_{j})}$
$={\frac{\exp\frac{1}{\beta}\widehat{r}_{\mathcal{D}}(\tau_{i})\exp\frac{1}{\beta}\widehat{r}_{\mathcal{P}}(\tau_{i})}{\sum_{j=1}^{n}\exp\frac{1}{\beta}\widehat{r}_{\mathcal{D}}(\tau_{j})\exp{\frac{1}{\beta}\widehat{r}_{\mathcal{P}}}(\tau_{j})}}$
$={\frac{\exp\big{(}\frac{1}{\beta}(\widehat{r}_{\mathcal{D}}(\tau_{i})+\widehat{r}_{\mathcal{P}}(\tau_{i}))\big{)}}{\sum_{j=1}^{n}\exp\big{(}\frac{1}{\beta}(\widehat{r}_{\mathcal{D}}(\tau_{j})+\widehat{r}_{\mathcal{P}}(\tau_{j}))\big{)}}}$
$=\pi^{*}_{i}\Big{(}\widehat{r}_{\mathcal{D}}+\widehat{r}_{\mathcal{P}}\Big{)}$ (22)
That is, the RLHF policy can be seen as the softmax policy for the sum of reward estimators.
AIHF estimator: In AIHF, assume that the lower-level problem uses (10), and the prior $\pi^{0}$ is uniform, then the optimal $\pi^{\rm AIHF}$ follows the same expression as in (15). One can then plug in such a solution to its loss function in (8a), and obtain the following reward learning problem:
$\widehat{r}^{\rm AIHF}=\arg\max_{r}\quad\ell_{\mathcal{D}+\mathcal{P}}(r):=|\mathcal{D}|L_{1}(r)+|\mathcal{P}|L_{2}(r)$ (23)
where $L_{1}(r):=\mathbb{E}_{\tau_{i}\sim\mathcal{D}}[\log\pi^{*}_{i}(r)]$ and $L_{2}(r):=\mathbb{E}_{(\tau_{j}\prec\tau_{i})\sim\mathcal{P}}[\log\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}]$.
Using (16) and (19), we obtain:
$\frac{\partial\ell_{\mathcal{D}+\mathcal{P}}(r)}{\partial r_{i}}$ =#$=\frac{\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}}{|\mathcal{D}|}|\mathcal{D}|+\sum_{j\neq i}\#\{\tau_{j}\prec\tau_{i}~{}\mbox{in}~{}\mathcal{P}\}-\pi^{*}_{i}|\mathcal{D}|-\sum_{j\neq i}\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}|\mathcal{P}_{i,j}|.$
Hence, the first-order condition is:
#$\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}+\sum_{j\neq i}\#\{\tau_{j}\prec\tau_{i}~{}\mbox{in}~{}\mathcal{P}\}$ $=\pi^{*}_{i}(r)(|\mathcal{D}|+\sum_{j\neq i}|\mathcal{P}_{i,j}|\frac{1}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)})$
Or equivalently,
$\pi^{*}_{i}(r)$ =#$=\frac{\#\{a_{i}~{}\mbox{in}~{}\mathcal{D}\}+\sum_{j\neq i}\#\{a_{j}\prec a_{i}~{}\mbox{in}~{}\mathcal{P}\}}{|\mathcal{D}|+\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho^{*}_{-(i,j)}(r)}.$ (24)
Example 1: With $\beta=1$ and only two actions $\tau_{1}$ and $\tau_{2}$. Since $\rho^{*}(r)_{-(1,2)}=1$, it follows that:
$\pi^{AIHF}_{1}:=\pi^{*}_{1}(\widehat{r}^{AIHF})$ =#$=\frac{\#\{\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}+\#\{\tau_{2}\prec\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}}{|\mathcal{D}|+|\mathcal{P}|}$
$=\frac{|\mathcal{D}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{r}_{\mathcal{D}})+\frac{|\mathcal{P}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{r}_{\mathcal{P}}).$
That is, the policy estimate obtained by using demonstrations and preferences is a weighted average of the polices estimated by using only demonstrations and only preferences. Slightly abusing notations, let $\pi^{*}_{1}:=\pi^{*}_{1}(r^{*})$ where $r^{*}$ is the ground-truth reward. It follows that $\mbox{Var}(\pi^{*}_{1}(\widehat{r}_{\mathcal{D}}))=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{D}|}$, $\mbox{Var}(\pi^{*}_{1}(\widehat{r}_{\mathcal{P}}))=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{P}|}$ and
$\mbox{Var}(\pi^{\rm AIHF}_{1})=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{D}|+|\mathcal{P}|}.$
This shows the AIHF policy estimator has less variance than either one of the estimators obtained from demonstrations only or preferences only.
To further illustrate, suppose $r(\tau_{1})=r(\tau_{2})$ and we have the following datasets:
#$\#\{\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}=\#\{\tau_{2}~{}\mbox{in}~{}\mathcal{D}\}=50,\quad\#\{\tau_{1}\succ\tau_{2}~{}\mbox{in}~{}\mathcal{P}\}$ $=6,\quad\#\{\tau_{2}\succ\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}=4.$
With the given data, $\pi^{\rm SFT}_{1}=\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{\#\{\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}}{|\mathcal{D}|}=\frac{50}{100}$ and the solution to (20) yields
$\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}=\frac{6}{10}$
Hence, $\pi_{1}^{\rm AIHF}=\frac{100}{10+100}\pi^{*}_{1}(\widehat{r}_{\mathcal{D}})+\frac{10}{10+100}\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{56}{110}$. It follows from (22) that:
$\pi^{\rm RLHF}_{1}$ $=\frac{\pi^{\rm SFT}_{1}\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\pi^{\rm SFT}_{1}\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\pi^{\rm SFT}_{2}\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}$
$=\frac{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}=\frac{6}{10}.$
In this example, the RLHF policy estimator is the farthest from ground-truth, because it does not correctly use the information provided by the demonstration data which in this case happens by chance to be correct $\pi^{\rm SFT}_{1}=\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{1}{2}$.
As a second example, again suppose $r(\tau_{1})=r(\tau_{2})$. Consider now the datasets:
{#$\{\#\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}=6,\quad\{\#\tau_{2}~{}\mbox{in}~{}\mathcal{D}\}=4,\quad\{\#\tau_{1}\succ\tau_{2}~{}\mbox{in}~{}\mathcal{P}\}=\{\#\tau_{2}\succ\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}=50.$ (25)
In this case, $\pi^{\rm SFT}_{1}=\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{6}{10}$ and the solution to (20) yields
$\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}=\frac{50}{100}.$
It follows from (22) that:
$\pi^{\rm RLHF}_{1}$ $=\frac{\pi^{\rm SFT}_{1}\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\pi^{\rm SFT}_{1}\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\pi^{\rm SFT}_{2}\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}$
$=\frac{\pi_{1}^{SFT}}{\pi_{1}^{SFT}+\pi_{2}^{SFT}}=\frac{6}{10}.$
Hence, $\pi_{1}^{\rm AIHF}=\frac{10}{10+100}\pi^{*}_{1}(\widehat{r}_{\mathcal{D}})+\frac{100}{10+100}\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{56}{110}$. In this example, the RLHF policy estimator is again farthest from ground-truth, because it does not correctly dismiss the information provided by the demonstration data which is less informative than preference.
Numerical Example 2. Let us use an illustrative example to show that RLHF method will result in significant data under-utilization when the demonstration coverage is limited. Assume that there are 50 actions, each with a ground-truth reward defined by $r_{i}=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{(\frac{i}{50}-\mu)^{2}}{2\sigma^{2}}}$, where $\mu=0.5$ and $\sigma=2$. Assume we can sample demonstration and preference from the ground truth reward distribution: demonstrations are sampled from the multinomial distribution, while preferences are sampled from the BTL model.
In an extreme scenario, let demonstrations cover actions 1 through 45, while preferences have full coverage across all actions. In the subsequent experiment, we initially sample 2000 demonstrations using the multinomial distribution $p_{i}=\frac{\exp r_{i}}{\sum^{j=45}_{j=1}\exp r_{j}}$, and obtain 200 preferences for each preference pair with $p(i\succ j)=\frac{\exp r_{i}}{\exp r_{j}+\exp r_{i}}$. We then calculate the optimal policy by solving the previously mentioned first order conditions, yielding the following result.
Figure 2: The optimal policy of RLHF, SFT, AIHF, and Ground-truth distribution. The left region of the red dotted line is included in the demonstration, while the right region is uncovered. We report the results with 100 random repeats.
From the result shown in Figure 2, we demonstrate that both SFT and RLHF transfer the weight from uncovered action to covered action when demonstration coverage is limited, as indicated by $\pi_{SFT}(a_{uncovered})=0$. Consequently, the weight of covered actions is significantly higher than the ground truth. However, this issue does not occur when jointly optimizing the demonstration and preference in the AIHF method.
5 Proposed Algorithm for AIHF Training
We are now ready to design algorithms for the proposed AIHF formulation (8). To begin with, first note that (8) takes a hierarchical form, and it belongs to the class of problem named bi-level optimization, first developed in the 70s (Fiacco and McCormick, 1990), and recently found many applications in machine learning (Wang et al., 2021; Liu et al., 2021, 2022). Generically speaking, bi-level problems are not easy to optimize; more specifically, in (8), the upper-level problem (8a) is a function of both the lower-level optimal solution $\pi_{\theta}$ and the true parameter $\theta$. It follows that a (stochastic) first-order algorithm for $L(\theta)$ involves some (potentially non-trivial) implicit gradient computation which often involves computing the Hessian matrix for the lower-level objective function. Fortunately, as we will show shortly, with some special choices of $L_{1}$, $L_{2}$, $L_{3}$, one can design some simple and very efficient algorithms.
Throughout this section, we are searching for a good policy $\pi_{\theta}$ and a reward estimate $r(\cdot,\cdot;\theta)$ to align with human feedback, where the policy $\pi_{\theta}$ is an optimal solution w.r.t. the certain reward estimate $r(\cdot,\cdot;\theta)$ according to the policy optimization problem (3.4). Due to such optimal policy constraint w.r.t. one explicit reward estimate, we design an algorithm to solve such a single-stage, bi-level problem which is different from DPO (Rafailov et al., 2023) that simply optimizes the fixed loss function (13) directly.
On a high level, the proposed algorithm alternates between a policy alignment step (which updates $\pi$ with a fixed reward $r(\cdot,\cdot;\theta)$), and a reward alignment step (which updates $\theta$ using a stochastic gradient, a function of the demonstration and preference data). Next, we study these steps in detail.
Policy Alignment Step. From our earlier discussion, we know that when $L_{3}$ takes the form (10), the optimal policy (for a fixed reward) is given by (5). Of course, one cannot directly compute such an optimal solution due to the fact that both $Q_{\theta}$ and the normalization term are unknown. Therefore one can adopt the standard approaches such as the well-known proximal policy optimization (PPO) (Schulman et al., 2017) algorithm to obtain an approximate optimal policy. It is worth noting that, when considering $T=1$, our discussion leading to (7) indicates the optimal policy takes a much simpler form. In this case, it is possible to consider simpler method than running PPO to obtain the optimal policy. One alternative way is to use a baseline estimated reward value to perform variance reduction Li et al. (2023), thus reducing the computational complexity.
It is important to note that, the point of the above discussion is that these different choices for solving the policy alignment problem can be incorporated into our overall approach.
Reward Alignment. In this step, we use a stochastic gradient-type algorithm to optimize $L(\theta)$. Towards this end, first observe that
$\nabla L(\theta)$ $=\alpha\nabla L_{1}(\pi_{\theta})+\nabla L_{2}(\theta).$ (26)
Clearly, regardless of the choice of $L_{2}$, $\nabla L_{2}$ is relatively easy to compute because the objective is directly related to $\theta$ since $L_{2}(\theta)$ can be regarded as one supervised learning loss and do not involve the optimal policy $\pi_{\theta}$. In particular, we have the following expressions:
$\nabla{L}^{\rm BTL}_{2}(\theta)$ $=\mathbb{E}_{(\tau_{l}\prec\tau_{w})\sim\pi^{P}}\Big{[}\nabla_{\theta}\log\Big{(}\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}\Big{)}\Big{]}.$ (27)
On the contrary, the computation of $\nabla L_{1}(\pi_{\theta})$ is more involved, since $L_{1}$ depends on $\theta$ implicitly through the corresponding optimal policy $\pi_{\theta}$. Fortunately, the following lemma indicates that this gradient has a simple and intuitive form as well.
Lemma 5.1
Suppose that $L_{1}$ takes the form of the objective (1) for reward learning from demonstrations, and suppose that $L_{3}$ takes the form (10) with $c(\cdot)$ being the KL-divergence w.r.t. some initial policy $\pi^{0}$. Then we have the following expression:
$\nabla_{\theta}L_{1}(\pi_{\theta})$ $=\mathbb{E}_{\tau\sim\pi^{\rm E},\tau^{\prime}\sim\pi_{\theta}}[\nabla_{\theta}\big{(}R(\tau;\theta)-R(\tau^{\prime};\theta)\big{)}]$ (28)
where $\pi_{\theta}$ is the optimal policy given the reward model parameterized by $\theta$, with the expression (5).
Intuitively, if the current policy $\pi_{\theta}$ has not matched $\pi^{\rm E}$ yet, then the reward should be improved by going towards the direction suggested by the expert trajectories, while going away from those generated by the current policy. Similar to the BTL model, from the gradient expression (28), it is clear that the optimization is toward the direction of increasing the gap between the reward of the real samples (demonstrations) and the synthetic ones (model generated continuations).
In practice, a few approximations need to be made to obtain a stochastic gradient of $L_{1}$. First, similarly, as before, the precise expectation cannot be obtained because the ground truth policy $\pi^{\rm E}$ is unknown. Denote an offline demonstration dataset as $\mathcal{D}^{\rm E}:=\{\tau\}$, then one can replace the expectations $\mathbb{E}_{\tau\sim\pi^{E}}$ by $\mathbb{E}_{\tau\sim\mathcal{D}^{\rm E}}$. Second, in the second expectation in (28), the trajectories $\tau^{\prime}$ are sampled from $\pi_{\theta}$, the optimal policy for a fixed reward parameterization by $\theta$. This means that the policy alignment step has to identify the optimal policy $\pi_{\theta}$ first, which, due to limitations such as computational constraints, and non-linear parameterization, is generally not possible. Instead, we propose to sample from the current policy $\pi^{k+1}$ obtained from the previous policy optimization step, where index $k$ represents the iteration counter. The above two steps are summarized in Algorithm 1.
Let us remark on the gradient expression of the proposed algorithm. Here, we explain how to construct the gradient estimator $g^{k}_{1}$ for $\nabla_{\theta}L_{1}(\theta_{k})$ defined in (28) and $g^{k}_{2}$ for $\nabla_{\theta}L_{2}(\theta_{k})$ defined in (27). At iteration $k$, to estimate $\nabla_{\theta}L_{1}(\theta_{k})$, we need to sample $N$ expert trajectories $\tau^{\rm E}$ from the expert demonstration dataset $\mathcal{D}^{\rm E}$ and $N$ agent trajectories $\tau^{\prime}$ from the policy $\pi^{k+1}$. Then the gradient estimator $g_{1}^{k}$ can be constructed as: $g_{1}^{k}:=\frac{1}{N}\sum_{i=1}^{N}\Big{[}\nabla_{\theta}R(\tau_{i};\theta)-\nabla_{\theta}R(\tau_{i}^{\prime};\theta)\Big{]}$. Furthermore, through sampling $M$ pairs of preference $(\tau_{l}\prec\tau_{w})$, we can further construct the gradient estimator of $\nabla_{\theta}L_{2}(\theta_{k})$ as: $g_{2}^{k}=\sum_{i=1}^{M}\nabla_{\theta}\log\Big{(}\sigma\big{(}R(\tau_{i,w};\theta)-R(\tau_{i,l};\theta)\big{)}\Big{)}$. It is interesting to note that, although the demonstration data is used, in the final algorithm, the gradient is always in contrastive form. The difference is that for the demonstration data, it is compared with its own model generation, while for preference data, two preference data is compared.
Before we close this section, let us remark on the computational complexity of the proposed algorithm. Note that our algorithm is motivated by a class of popular algorithms in bi-level optimization, where the upper-level and lower-level problems are updated alternatingly using stochastic optimization (Hong et al., 2020).
Specifically, one only needs to perform one (or a few steps) of policy optimization step (such as PPO) to obtain an improved $\pi^{k+1}$ before moving to the next iteration.
As demonstrated in our subsequent numerical results, the computational complexity of our method is comparable or sometimes even lower as compared to their sequential optimization counterparts such as RLHF. In our numerical results, we show that the proposed algorithm AIHF can achieve a comparable performance as RLHF while performing less policy optimization steps.
  Input: Initialize reward parameter $\theta^{0}$ and policy model $\pi^{0}$, the stepsize of reward update $\eta$. Let $\mathcal{P}$, $\mathcal{D^{\rm E}}$ denote the preference and the demonstration data, respectively.   for  Iteration $k=0,1,\ldots,K-1$ do      Policy Alignment: Optimizing $L_{3}$ by RL subroutine, e.g. PPO, to obtain one improved policy $\pi^{k+1}$      Data Sample I: Sample expert trajectories $\tau\sim\mathcal{D^{\rm E}}$ and agent trajectories from $\tau^{\prime}\sim\pi^{k+1}$      Data Sample II: Sample preference pair $(\tau_{l}\prec\tau_{w})\sim\mathcal{P}$      Estimating Gradient: Calculate the estimator $g^{k}:=\alpha g_{1}^{k}+g_{2}^{k}$ of $\nabla_{\theta}L(\theta_{k})=\alpha\nabla_{\theta}L_{1}(\theta_{k})+\nabla_{\theta}L_{2}(\theta_{k})$      Reward Alignment: $\theta^{k+1}:=\theta^{k}+\eta g^{k}$   end for
Algorithm 1 Alignment with Integrated Human Feedback (AIHF)
6 Experiments
In this section, we present numerical results for the proposed method. More specifically, we intend to address the following questions: 1) How are the data efficiency and model improvement efficiency (w.r.t. the KL divergence violation) of our proposed method compared with other baselines? 2) Can we achieve the same convergence results after integrating the three stages into one stage? 3) And finally, how much does demonstration help in reward learning in AIHF? To answer these questions, we consider several applications of the alignment problem. In this section, we show the numerical results for large language model alignment in the Helpful and Harmless (HH) dataset (Bai et al., 2022) and robotics control problem in MuJoCo Todorov et al. (2012). In the appendix, we have another set of experiment where the goal is to align the language model so that it generates movie reviews with positive sentiment on IMDb dataset (Maas et al., 2011).
6.1 MuJoCo Tasks
In MuJoCo, we consider several robotic control tasks with continuous action space. We evaluate the performance of our proposed algorithm on aligning robot behaviors with provided demonstrations and preference data. After the robot training stage, we leverage the ground-truth reward function from the environment to evaluate the performance.
Data. Following the similar data generation pipeline in Brown et al. (2019), we generate the expert demonstrations and preference dataset as follows. We first train an expert agent by leveraging the ground-truth reward function and the popular Soft Actor-Critic (SAC) algorithm Haarnoja et al. (2018), which is developed to solve policy optimization problems with continuous action space. During the training process, we save the policy checkpoints and collect $30$k samples from each checkpoint. To achieve precise control of dataset quality, we categorize the data collected into three different classes: low-, medium-, and high-quality datasets according to the performance of the checkpoints. Then we combine the low- and medium-quality data as the preference dataset and use high-quality as demonstration data.
Figure 3: Left: Hopper Environment Middle: HalfCheetah Environment Right: Walker2d Environment; AIHF  (orange) vs RLHF(blue); results are averaged over 3 independent runs. We use 10k demonstrations and 20k preferences. The RLHF curve is initialized from a policy pre-trained by BC; the AIHF from a random policy. The performance is compared against the # of SAC steps performed (for AIHF each policy alignment performs 5k steps of SAC.)
Results. We show that AIHF is able to integrate (insufficient amount of) demonstration data and (not-so-high-quality) preference data to generate high-quality policy, and it significantly outperforms the RLHF. In Fig. 3, we observe that due to the limited number of demonstration data Ross and Bagnell (2010); Zeng et al. (2022b), BC fails to provide a high-performing initialization for finetuning for RLHF. Moreover, since the preference data quality is only of low-to-medium quality, the RL step based on the learned reward model fails to significantly boost the fine-tuning performance. In contrast, clearly the proposed AIHF can effectively integrate preferences and demonstrations, leading to a more robust reward function and consequently, a high-quality policy.
6.2 LLM Task
In the experiment utilizing the HH dataset, we aim to generate helpful continuations through aligning large language models with human preference. We compare our proposed algorithm AIHF which updates reward and policy iteratively, with the standard RLHF benchmark. Additionally, $L_{2}(\cdot)$ is chosen to be $L_{2}^{\rm BTL}$ defined (11), and $L_{1}$, $L_{3}$ are chosen similarly as in MuJoCo experiment. More settings are listed in the Appendix A.2. For evaluation, we use the existing reward model PKU-Alignment/beaver-7b-v3.0-reward [^2] https://huggingface.co/PKU-Alignment/beaver-7b-v3.0-reward for the helpful continuation generation; use the existing sentiment classifier lvwerra/distilbert-imdb [^3] https://huggingface.co/lvwerra/distilbert-imdb as the ground truth reward model for the movie review generation task.
Data. We divide the training dataset of HH dataset into three different subsets: demonstration dataset, preference dataset, and RL training dataset, We employed the reward model PKU-Alignment/beaver-7b-v3.0-reward Dai et al. (2023); Ji et al. (2024) to filter the datasets. Specifically, We selected the 10k responses with the highest reward score as our demonstration dataset and utilized an additional 10,000 pairwise comparisons as preference data. The remaining prompts were then used to run the RL algorithm.
Setting. For the HH dataset, We first fine-tune the language models (Pythia-160M / Pythia-1B) Biderman et al. (2023) through supervised fine-tuning over all chosen responses from the HH dataset for 1 epoch, we call it full-SFT model and use it as our base model. Moreover, we supervised fine-tuned the language model using the selected top 10k "chosen" responses and named it the demonstration-SFT model. For the reward model, we use EleutherAI/pythia-1.4B  [^4] https://huggingface.co/EleutherAI/pythia-1.4b for both 160m and 1B policy as its base model.
Figure 4: Helpfulness-controlled Generation. Left: Results on 160M models, Right: Results on 1B models. We record the average scores of AIHF and RLHF on the Anthropic-HH test dataset, reporting the results across three different trials.
Results. We observe that the proposed approach AIHF performs effectively when initiated from both the demonstration-SFT model and the full-SFT model. As shown in Fig. 4, utilizing the same data, the AIHF  algorithms can eventually outperform RLHF irrespective of the initial model. Furthermore, according to the numerical results as shown in Fig. 5, compared with the RLHF benchmark, we see that the proposed AIHF algorithm has smaller deviation from the base model. This benefit of the AIHF approach is due to the fact that we incorporate the maximum likelihood IRL objective for both reward learning and policy learning. In this case, both reward model and policy model will be trained to align with the demonstrations, which are also used in the training process of the SFT stage. Fig. 5 shows AIHF  effectively maintains the policy close to the initial model. This advantage can be partially attributed to Eq. (8a), as the inclusion of the demonstration regularization term encourages the alignment with the demonstration distribution, thus bringing better performance without significant deviation from the initial model. The win rate (ratio of samples where the reward of our model’s generation is higher than the model compared) provided by Fig 5 further shows the superiority of the proposed approach.
Figure 5: Helpfulness-controlled Generation on 1B models. Left: KL divergence to the Demonstration-SFT policy, Right: win rate of AIHF and RLHF over the Demonstration-SFT model.
7 Conclusion
In this work, we study the alignment problem when diverse data sources from human feedback are available. Furthermore, we have developed an algorithmic framework that can integrate both expert demonstration and pairwise comparison data from human feedback to learn the reward functions for further guiding policy learning/model fine-tuning in the alignment pipeline. Through extensive evaluations on robotic control tasks and large language model alignment tasks, we demonstrate that our proposed method can outperform existing benchmarks on alignment tasks and is able to recover a better reward model to guide policy learning. One limitation is that our framework only incorporates expert demonstrations and pairwise comparisons for reward learning. In fact, it is promising to extend the information sources from human feedback/behaviors to further strengthen the reward learning/policy fine-tuning stages. For example, one can utilize precise comparisons and demonstrations from different aspects to further learn a powerful reward model in order to support better alignment quality.
Appendix A Appendix / supplemental material
A.1 Proof of Lemma 5.1
Proof. Here, under a reward parameter $\theta$ and the corresponding optimal policy $\pi_{\theta}$ of (5).
Moreover, under a fixed reward parameter $\theta$, we have defined the optimal policy $\pi_{\theta}$ as below:
$\pi_{\theta}:=\arg\max_{\pi}~{}\mathbb{E}_{\tau^{\rm A}\sim\pi}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\bigg{(}r(s_{t},a_{t};\theta)+\beta\mathcal{D}_{KL}\Big{(}\pi(\cdot|s_{t})||\pi^{0}(\cdot|s_{t})\Big{)}\bigg{)}\bigg{]}.$
According to Uehara et al. [2023], the optimal policy $\pi_{\theta}$ of (10) has the closed form expression as below:
$\pi_{\theta}(a|s)=\frac{\pi^{0}(a|s)\exp\big{(}\frac{Q_{\theta}(s,a;\theta)}{\beta}\big{)}}{\sum_{\tilde{a}\in\mathcal{A}}\pi^{0}(\tilde{a}|s)\exp\big{(}\frac{Q_{\theta}(s,\tilde{a};\theta)}{\beta}\big{)}},\quad\forall s\in\mathcal{S},a\in\mathcal{A}.$ (29)
Based on the closed form of $\pi_{\theta}$, we can also obtain the closed form of $V_{\theta}$ as following:
$V_{\theta}(s):=\beta\log\Big{(}\sum_{a\in\mathcal{A}}\pi^{0}(a|s)\exp\big{(}\frac{Q_{\theta}(s,a)}{\beta}\big{)}\Big{)}.$ (30)
Then we can re-write the demonstration loss $L_{1}(\theta)$ as below:
$L_{1}(\theta)$ $=\mathbb{E}_{\tau^{\rm E}\sim\pi^{\rm E}}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\log\pi_{\theta}(a_{t}|s_{t})\bigg{]}$
$=\mathbb{E}_{\tau^{\rm E}\sim\pi^{\rm E}}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\log\bigg{(}\frac{\pi^{0}(a_{t}|s_{t})\exp\big{(}\frac{Q_{\theta}(s_{t},a_{t})}{\beta}\big{)}}{\sum_{\tilde{a}\in\mathcal{A}}\pi^{0}(\tilde{a}|s_{t})\exp\big{(}\frac{Q_{\theta}(s_{t},\tilde{a})}{\beta}\big{)}}\bigg{)}\bigg{]}$
$=\mathbb{E}_{\tau^{\rm E}\sim\pi^{\rm E}}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\bigg{(}\log\Big{(}\pi^{0}(a_{t}|s_{t})\exp\big{(}\frac{Q_{\theta}(s_{t},a_{t})}{\beta}\big{)}\Big{)}-\log\Big{(}\sum_{\tilde{a}\in\mathcal{A}}\pi^{0}(\tilde{a}|s_{t})\exp\big{(}\frac{Q_{\theta}(s_{t},\tilde{a})}{\beta}\big{)}\Big{)}\bigg{)}\bigg{]}$
$=\mathbb{E}_{\tau^{\rm E}\sim\pi^{\rm E}}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\bigg{(}\log\pi^{0}(a_{t}|s_{t})+\frac{Q_{\theta}(s_{t},a_{t})}{\beta}-\log\Big{(}\sum_{\tilde{a}\in\mathcal{A}}\pi^{0}(\tilde{a}|s_{t})\exp\big{(}\frac{Q_{\theta}(s_{t},\tilde{a})}{\beta}\big{)}\Big{)}\bigg{)}\bigg{]}$
$=\frac{1}{\beta}\mathbb{E}_{\tau^{\rm E}\sim\pi^{\rm E}}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\bigg{(}\beta\log\pi^{0}(a_{t}|s_{t})+Q_{\theta}(s_{t},a_{t})-\beta\log\Big{(}\sum_{\tilde{a}\in\mathcal{A}}\pi^{0}(\tilde{a}|s_{t})\exp\big{(}\frac{Q_{\theta}(s_{t},\tilde{a})}{\beta}\big{)}\Big{)}\bigg{)}\bigg{]}$
$=\frac{1}{\beta}\mathbb{E}_{\tau^{\rm E}\sim\pi^{\rm E}}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\bigg{(}\beta\log\pi^{0}(a_{t}|s_{t})+Q_{\theta}(s_{t},a_{t})-V_{\theta}(s_{t})\bigg{)}\bigg{]}$ (31)
Then we can take gradient of $L_{1}(\theta)$ w.r.t. the reward parameter $\theta$, we have the following expression:
$\nabla L_{1}(\theta)$ $:=\frac{1}{\beta}\mathbb{E}_{\tau^{\rm E}\sim\pi^{\rm E}}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\bigg{(}\nabla_{\theta}\beta\log\pi^{0}(a_{t}|s_{t})+\nabla_{\theta}Q_{\theta}(s_{t},a_{t})-\nabla_{\theta}V_{\theta}(s_{t})\bigg{)}\bigg{]}$
$=\frac{1}{\beta}\mathbb{E}_{\tau^{\rm E}\sim\pi^{\rm E}}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\bigg{(}\nabla_{\theta}Q_{\theta}(s_{t},a_{t})-\nabla_{\theta}V_{\theta}(s_{t})\bigg{)}\bigg{]}$
$=\frac{1}{\beta}\mathbb{E}_{\tau^{\rm E}\sim\pi^{\rm E}}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\bigg{(}\nabla_{\theta}r(s_{t},a_{t};\theta)+\gamma\nabla_{\theta}V_{\theta}(s_{t+1})-\nabla_{\theta}V_{\theta}(s_{t})\bigg{)}\bigg{]}$
$=\frac{1}{\beta}\mathbb{E}_{\tau^{\rm E}\sim\pi^{\rm E}}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta)\bigg{]}-\frac{1}{\beta}\mathbb{E}_{s_{0}\sim\rho}\bigg{[}\nabla_{\theta}V_{\theta}(s_{0})\bigg{]}$ (32)
In order to calculate the expression of $\nabla L_{1}(\theta)$, we further derive the expression of $\nabla_{\theta}V_{\theta}(s_{0})$:
$\nabla_{\theta}V_{\theta}(s_{0})$ $=\nabla_{\theta}\bigg{(}\beta\log\Big{(}\sum_{a\in\mathcal{A}}\pi^{0}(a|s_{0})\exp\big{(}\frac{Q_{\theta}(s_{0},a)}{\beta}\big{)}\Big{)}\bigg{)}$
$=\beta\sum_{a\in\mathcal{A}}\frac{\pi^{0}(a|s_{0})\exp\big{(}\frac{Q_{\theta}(s_{0},a)}{\beta}\big{)}}{\sum_{a\in\mathcal{A}}\pi^{0}(a|s_{0})\exp\big{(}\frac{Q_{\theta}(s_{0},a)}{\beta}\big{)}}\frac{\nabla_{\theta}Q_{\theta}(s,a)}{\beta}$
$=\mathbb{E}_{a\sim\pi_{\theta}(\cdot|s_{0})}\Big{[}\nabla_{\theta}Q_{\theta}(s_{0},a)\Big{]}$
$=\mathbb{E}_{a_{0}\sim\pi_{\theta}(\cdot|s_{0}),s_{1}\sim P(\cdot|s_{0},a_{0})}\Big{[}\nabla_{\theta}r(s_{0},a_{0};\theta)+\gamma\nabla_{\theta}V_{\theta}(s_{1})\Big{]}$
$=\mathbb{E}_{\tau^{\rm A}\sim\pi_{\theta}}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta)\mid s_{0}\bigg{]}$ (33)
By plugging (33) into (32), we obtain the following expression:
$\nabla L_{1}(\theta)=\frac{1}{\beta}\mathbb{E}_{\tau^{\rm E}\sim\pi^{\rm E}}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta)\bigg{]}-\frac{1}{\beta}\mathbb{E}_{\tau^{\rm A}\sim\pi_{\theta}}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta)\bigg{]}$ (34)
A.2 Experiment Setup and Additional Result
A.2.1 MuJoCo Tasks
In MuJoCo, we consider several robotic control tasks with continuous action space. We evaluate the performance of our proposed algorithm on aligning robot behaviors with provided demonstrations and preference data. After the robot training stage, we leverage the ground-truth reward function from the environment to evaluate the performance.
Data. Following the similar data generation pipeline in Brown et al. [2019], we generate the expert demonstrations and preference dataset as follows. We first train an expert agent by leveraging the ground-truth reward function and the popular Soft Actor-Critic (SAC) algorithm Haarnoja et al. [2018], which is developed to solve policy optimization problems with continuous action space. During the training process, we save the policy checkpoints and collect $30$k samples from each checkpoint. To achieve precise control of dataset quality, we categorize the data collected into three different classes: low-, medium-, and high-quality datasets according to the performance of the checkpoints. Then we combine the low- and medium-quality data as the preference dataset and use high-quality as demonstration data.
Settings. We set $L_{1}(\cdot)$ using the likelihood objective (9), $L_{2}(\cdot)$ as the BTL objective (11), and $L_{3}$ as KL-regularized objective given in (10). We implement the RLHF Brown et al. [2019] using the following 3 steps: (i) utilize the demonstration dataset to pre-train the policy using behavior cloning (BC) Pomerleau [1988]; (ii) conduct reward learning using preference data; (iii) implement the SAC for policy fine-tuning. In the proposed AIHF, the policy $\pi^{0}$ is initiated with a random policy, and the policy optimization problem is also solved using SAC.
In all Mujoco experiments, we assess the performance of benchmark algorithms on Hopper, HalfCheetah, and Walker2d from OpenAI Gym. To ensure a fair comparison, we utilize an open-source implementation of SAC [^5] https://github.com/openai/spinningup as the foundational RL algorithm for all methods. Regarding the choice of the penalty function $c(\cdot)$, since our proposed pipeline, AIHF , initiates with a random policy where $\pi(a|s)$ follows a uniform distribution, the kl penalty function is reduced to an entropy function.
In SAC, both the policy network and Q network are (64, 64) MLPs with ReLU activation function, and the step size is set to $3*10^{-3}$, we parameterize the reward function by a (64, 64) MLPs with ReLU activation function. For the reward network, we use Adam as the optimizer, and the step size is set to be $1*10^{-4}$.
The quality of the preference dataset and demonstration dataset are listed as follows Tab. 1:
Task Dataset Non-prefer Data Prefer Data Demonstration Data
Hopper-v2 $2345.20\pm 329.93$ $3024.63\pm 40.52$ $3559.61\pm 73.12$
HalfCheetah-v2 $7226.37\pm 126.88$ $9434.42\pm 1315.13$ $11635.42\pm 236.51$
Walker2d-v2 $3952.60\pm 444.45$ $5091.71\pm 291.73$ $5453.41\pm 71.07$
Table 1: The quality of preference and demonstration.
In the following sections, we include additional details about language alignment experiments.
A.2.2 Sentiment-Controlled Generation
Dataset Generation: In the IMDb sentiment completion task, we generate the demonstrations and preference datasets using the following procedure. Initially, we train a Language Model by employing the ground-truth reward function distilbert-imdb and the Proximal Policy Optimization (PPO) algorithm on 30% of the training dataset for IMDb. Throughout the training process, we save the policy checkpoint every 500 PPO steps. Subsequently, we select an additional 40% of the training dataset and generate a response for each prompt for each checkpoint. According to the evaluation score of each generation, we categorize the data collected into different classes: low-, medium-, and high-quality datasets, then we combine low-quality and medium-quality as preference datasets, and use high-quality as demonstration datasets. We show the average score of each dataset in 3,
Training: After acquiring the preference and demonstration datasets, we train the proposed algorithm AIHF and baselines on the remaining 30% of prompts from the training dataset. We evaluate the performance of each algorithm using the test datasets for IMDb and HH, along with their corresponding ground truth reward functions. For the GPU resources, we use 8$\times$ A10 for all the experiments. Please refer to Table 2 for detailed hyperparameters and model initialization setting, and Table 3 for the quality of the preference and demonstration datasets we synthesize.
Results: Policy Quality. We find that the proposed approach works well when either preference or demonstration data, or both, are limited. From the 6, we see that by using the same amount of data (10k preference, 10k demonstration), AIHF-based algorithms achieve faster convergence than their RLHF and DPO counterparts. Next, in Fig. 7, we show that policy learned using AIHF  generates higher-quality continuations (measured by higher reward scores), as compared with those generated by RLHF, across a number of different combinations of preference and demonstration data size. and the size of preference data can significantly affect the quality of the estimated policy of RLHF, while it only affects the convergence speed of AIHF.
Figure 6: IMDb Sentiment-controlled Generation. Left: AIHF  vs RLHF; Right: AIHF-DPO vs DPO; These represent limited data regimes with 1k preference, and 1k demonstration. Results are averaged over 3 independent runs. The RLHF and DPO curves are initialized from the SFT policy; the AIHF-based algorithms are initialized from a random policy. The achieved reward is compared against the # of PPO (resp, DPO) steps performed (for AIHF each policy alignment performs $20$ steps of PPO).
(a) RLHF (left) vs AIHF (right); Fix demonstration number to 1000, gradually increase preference number.
(b) RLHF (left) vs AIHF  (right); Fix preference number to 3k, gradually increase demonstration number.
Figure 7: IMDb Sentiment-controlled Generation RLHF vs AIHF ; (a) different demonstration data size; (b) preference data.
Results: Distribution Mismatch. We show that AIHF  is able to alleviate the distribution mismatch between the generated trajectories by the policy, and the data that the learned reward model is able to rank. To evaluate the extend of such mismatch, we use the following three steps: (1) use 1k preference, 1k demonstration to train policy and reward model for RLHF and AIHF ; (2) for a given set of prompts from test dataset, use RLHF and AIHF  to perform generation; (3) use the trained reward models to rate the generation; (4) compare with the score generated by the ground truth reward lvwerra/distilbert-imdb. Fig. 8 illustrates that the reward score distribution produced by AIHF  aligns closely with that of the ground truth reward, whereas that generated by RLHF exhibits a poor match. These results show that the reward model learned by AIHF is able to correctly evaluate the generation produced by the final policy.
Figure 8: Comparison of the distribution of reward score generated by the trained reward models, and the ground truth reward model. RLHF vs ground truth (left); AIHF vs ground truth (right).
hyperparameters task IMDb Secontrolled Generation H&H Alignment
policy model GPT2 PYTHIA-160M & PYTHIA-1B
reward model GPT2 PYTHIA-1.4B
reward learning rate 1e-5 1e-5
policy learning rate 1.41e-5 1e-6
reward batch size 128 8
policy batch size 256 4
sample temperature 1.00 1.00
policy update times per iteration 20 6000
reward update times per iteration 12 20
Table 2: Hyperparameters of AIHF. H&H refers to Helpful and Harmless.
Task Dataset Non-prefer Data Prefer Data Demonstration Data
IMDb $0.27\pm 0.51$ $1.13\pm 0.42$ $2.23\pm 0.13$
Helpful and Harmless $-1.61\pm 0.24$ $-1.12\pm 0.17$ $-0.82\pm 0.05$
Table 3: The quality of preference and demonstration.
From Fig.9, our proposed algorithm AIHF  could obtain higher rewards than baseline methods in the IMDb setting for almost all KL values. Although AIHF  might get a low score from the ground truth reward model in the earlier step, AIHF  would get a higher reward with more iteration and optimization steps. This indicates that with the mix of demonstration data and preference data, we could prevent the policy from known issues of reward hacking, especially when the policy learned more human-aligned features beyond base models (high KL value). Moreover, AIHF  is persistent in the number of preference data, presenting that AIHF  could still gain benefit from the limited preference data in more optimization steps as long as the demonstration data is high quality enough.
Figure 9: The frontier of expected reward vs KL to the reference policy in IMDB dataset. fix the demonstration number to 3k Left: Using 1k preference; Middle: Using 2k preferences; Right: Using 3k preference
In Fig.10, we show the converge curves of AIHF-DPO vs vanilla DPO in different combinations of preference and demonstration data size. Moreover, we observe that AIHF-DPO is stable over all different settings, while DPO is unstable and its performance slowly decreases over time.
Figure 10: DPO vs AIHF-DPO Left: 1k Demonstration 1k Preference, Middle:1k Demonstration 2k Preference Right:1k Demonstration 3k Preference.
A.2.3 Helpfulness-Controlled Generation
Results: Reward Distribution. Further, in Fig. 11, we show the overall reward distribution of the continuation, we can observe the distribution of AIHF and RLHF have some overlap in low-quality continuation, however, AIHF can generate more high-quality continuations compared to RLHF, which shows that joint optimization can more effectively align the policy model with the demonstration distribution.
Figure 11: The Reward Distribution of Helpfulness-controlled Generation. Left: Result on 160m model, Right: Results on 1B model, This figure reports the reward distribution of generation evaluated by PKU-Alignment/beaver-7b-v3.0-reward for AIHF and RLHF.

### v2

@@
 Sequencing of learning tasks may facilitate computation at the expense of inefficient exploitation of data.
-To illustrate, consider the three-stage RLHF approach proposed in (Ouyang et al., 2022), in the extreme case where the amount of high-quality preference data is quite limited, the reward model trained cannot adequately reflect the preferences of the human, which may lead to unsatisfactory performance in the RL stage.
+To illustrate, consider the three-stage RLHF approach proposed in (Ouyang et al., 2022), in the extreme case where the amount of high-quality preference data is quite limited.
+In such a situation, the reward model trained cannot adequately reflect the preferences of the human, which may lead to unsatisfactory performance in the RL stage.
 This simple example underscores the importance of ensuring the reward model is consistent with the policy model.
@@
 Furthermore, existing work such as Direct Preference optimization (DPO) (Rafailov et al., 2023) and Inverse Preference Learning (IPL) (Hejna and Sadigh, 2023) both remove the need for explicit reward modeling, and they directly extract the policy from preferences.
-This greatly reduces the training complexity, but it has been observed that these algorithms can be unstable in the training process (Azar et al., 2023; Xu et al., 2024).
+This greatly reduced the training complexity, but it has been observed that these algorithms can be unstable in the training process (Azar et al., 2023; Xu et al., 2024).
 There is also a large number of works that aim to learn reward functions from rating (Daniel et al., 2014) or ranking (Yuan et al., 2023; Myers et al., 2022).
@@
 However, due to labeling noise, we have a preference data $(s_{t},\hat{a}_{t})\prec(s_{t},\tilde{a}_{t})$, therefore we have learned $R(a_{t},\tilde{a}_{t};\theta)>R(a_{t},\hat{a_{t}};\theta)$, yielding significant difference in the final policy.
-We refer the readers to Sec.
-4 for more discussion.
 3.2 A Meta-Formulation
@@
 ${\rm s.t.}$ $\quad\pi_{\theta}:=\arg\max_{\pi}~{}L_{3}(\pi;R(\cdot;\theta))$ (8b)
-where $\theta\in\mathbb{R}^{d}$ is a parameter; $L_{1}(\pi_{\theta})$ is a measure of fit of the parameterized policy $\pi_{\theta}$ to demonstration data and $L_{2}(R(\cdot;\theta))$ is a measure of fit of the parameterized reward model $R(\cdot;\theta)$ to the preference data; $L_{3}(\pi,R(\cdot;\theta))$ is a measure of performance of policy $\pi$ with respect to reward model $R(\cdot;\theta)$; $\alpha\geq 0$ is a balancing coefficient reflecting the relative size of demonstration versus preference data.
+where $\theta\in\mathbb{R}^{d}$ is a parameter; $L_{1}(\pi_{\theta})$ is a measure of fit of the parametrized policy $\pi_{\theta}$ to demonstration data and $L_{2}(R(\cdot;\theta))$ is a measure of fit of the parametrized reward model $R(\cdot;\theta)$ to the preference data and $L_{3}(\pi,R(\cdot;\theta))$ is a measure of performance of policy $\pi$ with respect to reward model $R(\cdot;\theta)$.
+and $\alpha\geq 0$ is one balancing coefficient reflecting the relative size of demonstration versus preference data.
 We refer the above formulation as Alignment with Integrated Human Feedback (AIHF); see Fig.
@@
 Let us begin with the choice of $L_{1}$.
-It can be directly instantiated by using an objective similar to (1), which is the likelihood function over the collected expert demonstrations.
+It can be directly instantiated by using one objective similar to (1), which is the likelihood function over the collected expert demonstrations.
 Note that we aim to optimize the reward parameter $\theta$ to align with human feedback in (8a), thus the objective of $L_{1}$ can be specialized as a maximum likelihood function over expert demonstrations as below:
@@
 $\begin{split}\max_{\theta}&\quad L_{2}(\theta)\;\;{\rm s.t.}\;\;\pi_{\theta}:=\arg\max_{\pi}L_{3}(\pi;R(\cdot;\theta)).\end{split}$ (12)
-We note that now the upper- and lower-level problems are completely decomposable, since the upper-level problem solves for the reward parameterization $\theta$, while the lower-level problem solves for the policy (for the given reward), yielding two separate problems, which are exactly the RM and the RL problems in the typical RLHF approach.
+Noticed that now the upper- and lower-level problems are completely decomposable, since the upper-level problem solves for the reward parameterization $\theta$, while the lower-level problem solves for the policy (for the given reward), yielding two separate problems, which are exactly the RM and the RL problems in the typical RLHF approach.
 Specialization to DPO-Type Approach.
@@
 $\frac{\partial\ell_{\mathcal{P}}(r)}{\partial r_{i}}$ $=\frac{1}{\beta}\sum_{j\neq i}\big{(}\frac{\#\{\tau_{j}\prec\tau_{i}~{}\mbox{in}~{}\mathcal{P}\}}{|\mathcal{P}_{i,j}|}-\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\big{)}\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}.$ (19)
-The first-order condition can be written as:
+The first order condition can be written as:
 $\pi^{*}_{i}(r)$ $=\frac{\sum_{j\neq i}\{\#a_{i}\succ a_{j}~{}\mbox{in}~{}\mathcal{P}\}}{\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho^{*}_{-(i,j)}(r)}$ (20)
@@
 $\frac{\partial^{2}\ell_{\mathcal{P}}(r)}{\partial r_{i}^{2}}=-\frac{1}{\beta^{2}}\sum_{j\neq i}\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\frac{\pi^{*}_{j}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}$ $~{}~{}\frac{\partial^{2}_{r_{i}}\ell_{\mathcal{P}}(r)}{\partial r_{j}\partial r_{i}}=\frac{1}{\beta^{2}}\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\frac{\pi^{*}_{j}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}.$
-Note that compared with the previous case with only demonstration data, here the first-order solution does not imply global optimality since the Hessian is not negative semi-definite.
+Note that compared with the previous case with only demonstration data, here the first-order solution does not imply global optimality since the Hessian is negative semi-definite.
 Next, by using the above analysis, let us understand the policies that can be obtained by different alignment algorithms.
@@
 $=\pi^{*}_{i}\Big{(}\widehat{r}_{\mathcal{D}}+\widehat{r}_{\mathcal{P}}\Big{)}$ (22)
-That is, the RLHF policy can be seen as the softmax policy for the sum of reward estimators.
+That is, the RLHF policy can be seen as the softmax policy for the sum of rewards estimator.
 AIHF estimator: In AIHF, assume that the lower-level problem uses (10), and the prior $\pi^{0}$ is uniform, then the optimal $\pi^{\rm AIHF}$ follows the same expression as in (15).
@@
 Or equivalently,
-$\pi^{*}_{i}(r)$ =#$=\frac{\#\{a_{i}~{}\mbox{in}~{}\mathcal{D}\}+\sum_{j\neq i}\#\{a_{j}\prec a_{i}~{}\mbox{in}~{}\mathcal{P}\}}{|\mathcal{D}|+\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho^{*}_{-(i,j)}(r)}.$ (24)
+$\pi^{*}_{i}(r)$ =#$=\frac{\#\{a_{i}~{}\mbox{in}~{}\mathcal{D}\}+\sum_{j\neq i}\#\{a_{j}\prec a_{i}~{}\mbox{in}~{}\mathcal{P}\}}{|\mathcal{D}|+\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho^{*}_{-(i,j)}(r)}$ (24)
 Example 1: With $\beta=1$ and only two actions $\tau_{1}$ and $\tau_{2}$. Since $\rho^{*}(r)_{-(1,2)}=1$, it follows that:
 $\pi^{AIHF}_{1}:=\pi^{*}_{1}(\widehat{r}^{AIHF})$ =#$=\frac{\#\{\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}+\#\{\tau_{2}\prec\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}}{|\mathcal{D}|+|\mathcal{P}|}$
-$=\frac{|\mathcal{D}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{r}_{\mathcal{D}})+\frac{|\mathcal{P}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{r}_{\mathcal{P}}).$
-That is, the policy estimate obtained by using demonstrations and preferences is a weighted average of the polices estimated by using only demonstrations and only preferences.
-Slightly abusing notations, let $\pi^{*}_{1}:=\pi^{*}_{1}(r^{*})$ where $r^{*}$ is the ground-truth reward.
+$=\frac{|\mathcal{D}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{r}_{\mathcal{D}})+\frac{|\mathcal{P}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})$
+That is the policy estimated obtained by using demonstrations and preferences is a weighted average of the polices estimated by using only demonstrations and only preferences.
+Abusing notation let $\pi^{*}_{1}:=\pi^{*}_{1}(r^{*})$ where $r^{*}$ is the ground-truth reward.
 It follows that $\mbox{Var}(\pi^{*}_{1}(\widehat{r}_{\mathcal{D}}))=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{D}|}$, $\mbox{Var}(\pi^{*}_{1}(\widehat{r}_{\mathcal{P}}))=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{P}|}$ and
-$\mbox{Var}(\pi^{\rm AIHF}_{1})=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{D}|+|\mathcal{P}|}.$
+$\mbox{Var}(\pi^{\rm AIHF}_{1})=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{D}|+|\mathcal{P}|}$
 This shows the AIHF policy estimator has less variance than either one of the estimators obtained from demonstrations only or preferences only.
@@
 $\pi^{\rm RLHF}_{1}$ $=\frac{\pi^{\rm SFT}_{1}\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\pi^{\rm SFT}_{1}\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\pi^{\rm SFT}_{2}\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}$
-$=\frac{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}=\frac{6}{10}.$
+$=\frac{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}=\frac{6}{10}$
 In this example, the RLHF policy estimator is the farthest from ground-truth, because it does not correctly use the information provided by the demonstration data which in this case happens by chance to be correct $\pi^{\rm SFT}_{1}=\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{1}{2}$.
-As a second example, again suppose $r(\tau_{1})=r(\tau_{2})$. Consider now the datasets:
-{#$\{\#\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}=6,\quad\{\#\tau_{2}~{}\mbox{in}~{}\mathcal{D}\}=4,\quad\{\#\tau_{1}\succ\tau_{2}~{}\mbox{in}~{}\mathcal{P}\}=\{\#\tau_{2}\succ\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}=50.$ (25)
+Consider now the datasets:
+{#$\{\#\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}=6,\quad\{\#\tau_{2}~{}\mbox{in}~{}\mathcal{D}\}=4,\quad\{\#\tau_{1}\succ\tau_{2}~{}\mbox{in}~{}\mathcal{P}\}=\{\#\tau_{2}\succ\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}=50$ (25)
 In this case, $\pi^{\rm SFT}_{1}=\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{6}{10}$ and the solution to (20) yields
-$\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}=\frac{50}{100}.$
+$\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}=\frac{50}{100}$
 It follows from (22) that:
 $\pi^{\rm RLHF}_{1}$ $=\frac{\pi^{\rm SFT}_{1}\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\pi^{\rm SFT}_{1}\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\pi^{\rm SFT}_{2}\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}$
-$=\frac{\pi_{1}^{SFT}}{\pi_{1}^{SFT}+\pi_{2}^{SFT}}=\frac{6}{10}.$
+$=\frac{\pi_{1}^{SFT}}{\pi_{1}^{SFT}+\pi_{2}^{SFT}}=\frac{6}{10}$
 Hence, $\pi_{1}^{\rm AIHF}=\frac{10}{10+100}\pi^{*}_{1}(\widehat{r}_{\mathcal{D}})+\frac{100}{10+100}\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{56}{110}$.
@@
 Furthermore, through sampling $M$ pairs of preference $(\tau_{l}\prec\tau_{w})$, we can further construct the gradient estimator of $\nabla_{\theta}L_{2}(\theta_{k})$ as: $g_{2}^{k}=\sum_{i=1}^{M}\nabla_{\theta}\log\Big{(}\sigma\big{(}R(\tau_{i,w};\theta)-R(\tau_{i,l};\theta)\big{)}\Big{)}$.
-It is interesting to note that, although the demonstration data is used, in the final algorithm, the gradient is always in contrastive form.
-The difference is that for the demonstration data, it is compared with its own model generation, while for preference data, two preference data is compared.
+Here, as a remark, although the demonstration data is used, but in the final algorithm, the gradient is always in constrastive form.
+The difference is that the for the demonstration data, it is compared with its own model generation, while for preference data, two preference data is compared.
 Before we close this section, let us remark on the computational complexity of the proposed algorithm.
@@
 (2023) through supervised fine-tuning over all chosen responses from the HH dataset for 1 epoch, we call it full-SFT model and use it as our base model.
-Moreover, we supervised fine-tuned the language model using the selected top 10k "chosen" responses and named it the demonstration-SFT model.
+Moreover, we also SFT the language model using the selected top 10k chosen responses and name it as demonstration-SFT model.
 For the reward model, we use EleutherAI/pythia-1.4B  [^4] https://huggingface.co/EleutherAI/pythia-1.4b for both 160m and 1B policy as its base model.
@@
 Then we combine the low- and medium-quality data as the preference dataset and use high-quality as demonstration data.
+Other details are given in Appendix A.2.
 Settings.

### v3

@@
-Joint Demonstration and Preference Learning Improves Policy Alignment with Human Feedback
+Learning Reward and Policy Jointly from Demonstration and Preference Improves Alignment
 Abstract
-Aligning human preference and value is an important requirement for building contemporary foundation models and embodied AI.
-However, popular approaches such as reinforcement learning with human feedback (RLHF) break down the task into successive stages, such as supervised fine-tuning (SFT), reward modeling (RM), and reinforcement learning (RL), each performing one specific learning task.
-Such a sequential approach results in serious issues such as significant under-utilization of data and distribution mismatch between the learned reward model and generated policy, which eventually lead to poor alignment performance.
-We develop a single stage approach named Alignment with Integrated Human Feedback (AIHF), capable of integrating both human preference and demonstration to train reward models and the policy.
-The proposed approach admits a suite of efficient algorithms, which can easily reduce to, and leverage, popular alignment algorithms such as RLHF and Directly Policy Optimization (DPO), and only requires minor changes to the existing alignment pipelines.
-We demonstrate the efficiency of the proposed solutions with extensive experiments involving alignment problems in LLMs and robotic control problems in MuJoCo.
-We observe that the proposed solutions outperform the existing alignment algorithms such as RLHF and DPO by large margins, especially when the amount of high-quality preference data is relatively limited.
-[…]
+Aligning to human preferences and/or intentions is an important requirement for contemporary foundation models.
+To ensure alignment, popular approaches such as reinforcement learning with human feedback (RLHF) break down the task into three stages: (i) a model is computed with supervised fine-tuning (SFT) based upon large demonstrations data, (ii) a reward model (RM) is estimated based upon human feedback data, and (iii) reinforcement learning (RL) is used to further refine the SFT model by optimizing the estimated reward model.
+Demonstrations and human feedback data reflect human user preferences in different ways.
+As a result, the reward model estimate obtained from only human feedback data is likely not as accurate as a reward model estimate obtained from both demonstration and human feedback data.
+A policy model that optimizes the reward model estimate obtained from both demonstration and human feedback data will likely exhibit better alignment performance.
+We introduce a tractable algorithm for finding the reward and policy models and provide a finite-time performance guarantee.
+Additionally, we demonstrate the efficiency of the proposed solution with extensive experiments including alignment problems in LLMs and robotic control problems in MuJoCo.
+We observe that the proposed solutions outperform the existing alignment algorithm by large margins.
 1 Introduction
-Background.
 As ChatGPT has taken the world by storm, it is clear that AI systems will soon become ubiquitous in our lives.
-For instance, Large Language Models (LLMs) have been used to solve hard problems including video gaming (Berner et al., 2019; Mnih et al., 2015), autonomous control (Bellemare et al., 2020), and robotic manipulation (Kalashnikov et al., 2018; Kober and Peters, 2008).
-However, for AI systems to be able to smoothly interact with humans, and reliably help perform mission-critical tasks, it is critical that the system learns, or aligns with, the human’s preferred strategy, under specific contexts, and adapts to evolving user preferences and needs (Leike et al., 2018).
-Achieving this level of alignment between AI and humans requires a comprehensive approach.
-AI systems must be capable of not only responding to instructions but also being sensitive to the nuances of human behavior and context, ensuring that their actions align with individual values and societal norms.
-On the other hand, various undesirable behaviors exhibited by advanced AI systems, e.g.
-deception (Park et al., 2023) and manipulation (Perez et al., 2022) have raised concerns about the potential hazards from AI systems.
-The “alignment” task mentioned above has already proven to be useful in a few applications, such as in ChatGPT.
-Despite its recent popularity, the research area is still in its infancy.
-There are a number of critical open challenges, e.g.
-1) How to solve the alignment problem leveraging diverse human feedback to further boost alignment performance?
-2) How to enable AI systems to continuously learn and adapt to time-varying preferences, changing contexts, and emerging societal norms?
+For instance, Large Language Models (LLMs) have been used to solve hard problems including video gaming (Berner et al., 2019; Mnih et al., 2015), autonomous control (Bellemare et al., 2020), and robotic manipulation (Kalashnikov et al., 2018; Kober & Peters, 2008).
+In this context, the notion of alignment plays an increasingly important role in the design and training of AI systems.
+Loosely speaking, alignment refers to the performance guarantee that the AI system will generate outcomes that are intended or preferred by the human user without undesirable side effects or behaviors such as deception (Park et al., 2023) or manipulation (Perez et al., 2022).
+As human user intentions or preferences may vary under specific contexts, it is critical that the AI system adapts to evolving user preferences and/or intentions (Leike et al., 2018).
+The alignment problem is a learning problem with (at least) three types of input data: the demonstration data (consists of prompts and human-generated continuations), the preference data (consists of prompts and pairs of human-ranked responses), as well as prompts without any responses.
+Moreover, the process of aligning an LLM model is typically undertaken in successive stages.
+For example, the well-known RLHF approach adopted by Ouyang et al.
+(2022) starts with a supervised fine-tuning model (SFT) followed by reward model (RM) estimation based upon human-labeled preference data.
+The process closes with a final alignment stage in which reinforcement learning (RL) is used to optimize the estimated reward model.
+Similar strategies have been used in other related works such as Rafailov et al.
+(2023); Li et al.
+(2023); Zhu et al.
+(2023); Liu et al.
+(2023).
+The approach to alignment based on successive stages may facilitate computation, but it is at the expense of inefficient exploitation of data.
+To illustrate, consider the three-stage RLHF approach proposed in Ouyang et al.
+(2022), in the extreme case where the amount of high-quality preference data is quite limited, the reward model trained cannot adequately reflect the preferences of the human, which may lead to unsatisfactory performance in the RL stage.
+Further, the reward model estimate obtained from only the preference data fails to exploit the information about human users’ preferences that are implicit in demonstration data.
+It is therefore reasonable to expect that a policy model that is fine-tuned with the reward model estimate obtained from both demonstration and human feedback may exhibit better alignment performance.
+An alternative to the successive approach to alignment consists of jointly training the reward and policy models by leveraging demonstration and preference data.
+In contrast to the successive approach adopted in most of the current alignment approaches, the joint approach to reward and policy learning makes use of all available data, hence mitigating the risk of optimizing an inaccurate reward model.
+However, a joint approach to learning reward and policy models may improve alignment at the expense of potentially significant additional computational effort.
 Contribution.
-In this work, we address the first question posed above.
-That is, we develop formulations and efficient algorithms capable of integrating different sources of human response data to achieve the best alignment performance.
-Most of the existing alignment algorithms are sequential, i.e.
-they separate the alignment process into a number of successive stages, each utilizing one distinctive category of data.
-For example, the well-known RLHF approach adopted by (Ouyang et al., 2022) breaks the alignment step into the supervised finetuning (SFT), reward modeling (RM), and reinforcement learning (RL) stages, where the SFT stage uses demonstration data, the RM stage uses the preference data, and the RL stage uses a set of available prompts.
-Similar strategies have been used in other related works such as (Rafailov et al., 2023; Li et al., 2023; Zhu et al., 2023; Liu et al., 2023).
-Sequencing of learning tasks may facilitate computation at the expense of inefficient exploitation of data.
-To illustrate, consider the three-stage RLHF approach proposed in (Ouyang et al., 2022), in the extreme case where the amount of high-quality preference data is quite limited.
-In such a situation, the reward model trained cannot adequately reflect the preferences of the human, which may lead to unsatisfactory performance in the RL stage.
-This simple example underscores the importance of ensuring the reward model is consistent with the policy model.
-Such consistency can not be guaranteed with sequential execution of the RM and RL tasks.
-The alignment problem is a learning problem with (at least) three types of input data: the demonstration data (consists of prompts and human-generated continuations), the preference data (consists of prompts and pairs of human-ranked responses), as well as prompts without any responses.
-The input to the learning problem, besides these data, also includes a model pre-trained using a large corpus of text data.
-The output of the learning process is a fine-tuned model aligned with human preferences/values which govern the generation of demonstration and preference data.
-To integrate these diverse data, we propose a joint learning approach, which is informed by recent advances in Inverse Reinforcement Learning (IRL) (Arora and Doshi, 2021; Zeng et al., 2022b), stochastic choice theory (Blavatskyy and Pogrebna, 2010) and bi-level optimization (Hong et al., 2020; Ji et al., 2021; Khanduri et al., 2021).
-More specifically, the proposed formulation integrates SFT, RM, and RL into a single stage, so that reward modeling and policy optimization can fully leverage all three kinds of data mentioned above in alignment.
-Several existing alignment schemes, such as RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2023), and some of their extensions can be seen as particular instances of the proposed formulation.
-We observe that the proposed solutions outperform the existing alignment algorithms such as RLHF by large margins, especially when the amount of high-quality preference data is relatively limited.
+We introduce an algorithm jointly learning reward and policy models named Alignment with Integrated Human Feedback (AIHF) with a finite-time performance guarantee.
+This approach leverages recent advances in Inverse Reinforcement Learning (IRL) (Arora & Doshi, 2021; Zeng et al., 2022b), stochastic choice theory (Blavatskyy & Pogrebna, 2010) and bi-level optimization (Hong et al., 2020; Ji et al., 2021; Khanduri et al., 2021).
+The proposed formulation integrates SFT, RM, and RL into a single stage, so that reward modeling and policy optimization can fully leverage all the available human feedback data.
+More specifically, in the proposed algorithm, the policy is updated to improve alignment with the current reward model estimate and the reward model is updated to improve the fit to demonstration and human feedback data.
+As a result, upon convergence, the resulting reward and policy models are consistent in the sense that (i) the policy model is optimal with respect to the reward model and (ii) the reward model maximizes the fit to both demonstration and human feedback data.
+Several existing alignment schemes, such as RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2023) and some of their extensions can be seen as particular instances of the proposed formulation.
+We provide ample empirical evidence that the proposed AIHF solution outperforms the existing alignment algorithms by large margins, especially when the data is unbalanced, where the quality and/or quantity of one data category is worse/smaller than that of the other.
 2 Preliminaries and Related Work
@@
 The Finite-Horizon MDP Model.
-Consider a Markov decision process (MDP) defined by the tuple $(\mathcal{S},\mathcal{A},P,\rho,r,\gamma)$, which consists of the state space $\mathcal{S}$, the action space $\mathcal{A}$, the transition dynamics $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$, the initial state distribution $\rho(\cdot)$, the reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ and the discounted factor $\gamma>0$.
-Let $s_{t}\in\mathcal{S}$, $a_{t}\in\mathcal{A}$ denote the state and action variables at time $t$ respectively, where $a_{t}$ is generated according to some policy $\pi(\cdot|s_{t}):\mathcal{S}\to\Delta_{|\mathcal{A}|}$, where $\Delta_{|\mathcal{A}|}$ denotes the probability simplex in $\mathbb{R}^{|A|}$.
-Note that viewing the generation process of language models as an MDP, $s_{t}$’s and $a_{t}$’s are referred to as the given sequence and next token to be generated, respectively; $\pi(\cdot)$ then represents the language model itself.
-Define $\tau:=\{(s_{t},a_{t})\}_{t=1}^{T}$ as a (finite time) trajectory of state and action pairs.
-Let $\mathcal{H}_{T}\subset\prod_{t=0}^{T-1}\big{(}S\times A\big{)}$ denote all feasible state/action pairs.
-Human Feedback Generation.
-Assume that the human demonstration $\tau$ is generated from a ground-truth human expert policy $\pi^{\rm E}$, i.e., for each state $s_{t}$, we have $a_{t}\sim\pi^{\rm E}(\cdot|s_{t})$.
-Moreover, let ‘$\tau_{l}\prec\tau_{w}$’ indicate the winner trajectory $\tau_{w}$ is preferred over the trajectory $\tau_{l}$.
-Further, we denote $(\tau_{l}\prec\tau_{w})\sim\pi^{\rm P}$ as the preference oracle.
-Note that when we present our problem formulation, we will use the above data generation distributions so quantities such as likelihood can be properly defined.
-Finite-sample version of demonstration and preference data will be introduced subsequently to facilitate algorithm design.
-Since reward modeling is a critical task for achieving state-of-the-art alignment performance, below we review a few popular paradigms for reward learning.
-2.2 Reward Learning from Preference Data
-RLHF is a popular technique for finetuning AI systems to align with human preferences and values.
-The RLHF approach proposed in (Stiennon et al., 2020; Ouyang et al., 2022) consists of the following three-stage: 1) the supervised fine-tuning (SFT) stage, where the demonstration data is used to fine-tune the model in a supervised manner; 2) the reward modeling (RM) stage, where the preference data is used to train a reward model; 3) the reinforcement learning (RL) stage, where the SFT model is further finetuned by running RL using the trained reward model.
-Furthermore, existing work such as Direct Preference optimization (DPO) (Rafailov et al., 2023) and Inverse Preference Learning (IPL) (Hejna and Sadigh, 2023) both remove the need for explicit reward modeling, and they directly extract the policy from preferences.
-This greatly reduced the training complexity, but it has been observed that these algorithms can be unstable in the training process (Azar et al., 2023; Xu et al., 2024).
-There is also a large number of works that aim to learn reward functions from rating (Daniel et al., 2014) or ranking (Yuan et al., 2023; Myers et al., 2022).
-However, these works often fail when the quality of the preference dataset is low Morimura et al.
-(2024).
-In our approach, the inclusion of additional demonstrations can be leveraged to counteract the adverse effects of limited preference quality.
+A Markov decision process (MDP) is the tuple $(\mathcal{S},\mathcal{A},P,\rho,r,\gamma)$, wherein $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ denotes the transition probabilities, $\rho(\cdot)$ is the initial state distribution, $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ denotes the reward function and $\gamma\in(0,1)$ denotes the discount factor.
+For every $s_{t}\in\mathcal{S}$, a randomized policy $\pi(\cdot|s_{t})$ is a probability distribution in $\Delta_{|\mathcal{A}|}$, the unit simplex in $\mathbb{R}^{|A|}$.
+Define $\tau:=\{(s_{t},a_{t})\}_{t=1}^{T}$ as a (finite horizon $T$) trajectory of state and action pairs.
+Let $\mathcal{H}_{T}\subset\prod_{t=1}^{T}\big{(}\mathcal{S}\times\mathcal{A}\big{)}$ denote all feasible state/action sequence of length $T$.
+MDP Model of LLM.
+The generation of text by a language model can be seen as sampling from policies in an MDP model.
+Specifically, each state $s_{t}=(x,y_{1:t-1})$ includes the prompt $x$ and all response tokens produced up to that point $y_{1:t-1}$.
+Each action $a_{t}=y_{t}$ represents a token from the vocabulary.
+The transition kernel $P$ is deterministic, i.e.
+given tokens $s_{t}=(x,y_{1:t-1})$ and $a_{t}=y_{t}$, the environment will transition to $s_{t+1}=(x,y_{1:t})$.
+An LLM can be seen as a policy $\pi(\cdot|s_{t})$ so that a response of length $T>0$ to prompt $x$ is obtained with probability:$\pi(y_{1:T}|x):=\prod_{i=1}^{T}\pi(y_{i}|x,y_{1:i-1})$
+Human Feedback Data.
+Let $\tau:=(y_{1:T},x)$ denote a finite text produced in response to prompt $x$.
+For a pair of sequences $(\tau_{l},\tau_{w})$ (which we assume of the same length $T$ for ease of exposition) we write $\tau_{l}\prec\tau_{w}$ to indicate the sequence $\tau_{w}$ is preferred over the sequence $\tau_{l}$.
+Following the Bradley-Terry-Luce (BTL) model (Bradley & Terry, 1952), the distribution of preferences over pairs $(\tau_{l},\tau_{w})$ can be modeled as follows:
+$P\big{(}\tau_{w}\succ\tau_{l}\big{)}=\frac{\exp R(\tau_{w};\theta)}{\exp R(\tau_{w};\theta)+\exp R(\tau_{l};\theta)}=\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}$ (1)
+where $\sigma$ is the sigmoid function and $R(\tau;\theta):=\sum_{t\geq 1}^{T}\gamma^{t}r(s_{t},a_{t};\theta)$ and $r(s_{t},a_{t};\theta)$ is a reward model parametrized by $\theta\in\mathbb{R}^{d}$.
+2.2 The RLHF Pipeline
+RLHF is a popular technique for fine-tuning AI systems to align with human preferences and values.
+The RLHF approach proposed in Stiennon et al.
+(2020); Ouyang et al.
+(2022) consists of the following three-stage: 1) the supervised fine-tuning (SFT) stage, where the demonstration data is used to fine-tune the model in a supervised manner; 2) the reward modeling (RM) stage, where the preference data is used to train a reward model; 3) the reinforcement learning (RL) stage, where the SFT model is further fine-tuned by running RL using the trained reward model.
+Specifically, the RLHF pipeline can be formally described as follows:
+Supervised Fine-Tuning (SFT): Given a demonstration dataset $\mathcal{D}$ consisting of sequences of the form $\tau=\{(s_{t},a_{t})\}_{t\geq 0}$ the goal is the find the policy $\pi_{\rm SFT}(\cdot|s_{t})$ that maximizes likelihood, i.e.:
+$\pi_{\rm SFT}$ $=\arg\max_{\pi}\mathbb{E}_{\tau\sim\mathcal{D}}\Big{[}\log\prod_{t\geq 0}\Big{(}\pi(a_{t}|s_{t})\Big{)}^{\gamma^{t}}\Big{]}$ (2)
+Reward Modeling (RM): Based upon a dataset $\mathcal{P}$ of preferences over pairs $(\tau_{l},\tau_{w})$ the estimation of a reward learning problem can be formulated as the following Bradley-Terry-Luce (BTL) model (Bradley & Terry, 1952) (with $\beta>0$ a hyper-parameter):
+$\max_{\theta\in\mathbb{R}^{d}}\ell_{\rm RM}(\theta):=\mathbb{E}_{(\tau_{w}\succ\tau_{l})\in\mathcal{P}}\Big{[}\log\Big{(}\sigma\big{(}\frac{1}{\beta}\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}\Big{)}\Big{]}.$ (3)
+Reinforcement Learning (RL): Let $\hat{\theta}_{\mathcal{P}}$ denote the solution to problem (3). The last stage in the RLHF development pipeline consists of solving the problem:
+$\pi_{\rm RLHF}$ $=\arg\max_{\pi}\mathbb{E}_{\tau\sim\pi}\Big{[}\sum_{t\geq 0}\gamma^{t}\big{[}r(s_{t},a_{t};\hat{\theta}_{\mathcal{P}})-\beta{D}_{\rm KL}\Big{(}\pi(\cdot|s_{t})\|\pi_{\rm SFT}(\cdot|s_{t})\Big{)}\Big{]}$ (4)
+where ${D}_{\rm KL}\Big{(}\pi(\cdot|s_{t})\|\pi_{\rm SFT}(\cdot|s_{t})\Big{)}:=\sum_{a\in\mathcal{A}}\pi(a|s_{t})\log\frac{\pi(a|s_{t})}{\pi_{\rm SFT}(a|s_{t})}$ is the Kullback-Leibler (KL) divergence, $\pi_{\rm SFT}$ is the supervised fine-tuning model.
 2.3 Reward Learning using Demonstration Data
@@
 mimics) an expert policy $\pi^{E}$ given the demonstration data $\mathcal{D}$.
-For example, the well-known maximum entropy IRL (MaxEnt-IRL) framework (Ziebart, 2010; Ziebart et al., 2013; Bloem and Bambos, 2014; Zhou et al., 2017) finds a policy maximizing entropy subject to the expected features that match the empirical averages in the expert’s observation dataset.
+For example, the well-known maximum entropy IRL (MaxEnt-IRL) framework (Ziebart, 2010; Ziebart et al., 2013; Bloem & Bambos, 2014; Zhou et al., 2017) finds a policy maximizing entropy subject to the expected features that match the empirical averages in the expert’s observation dataset.
 However, this approach can only be used to model linear rewards.
-Subsequent works such as (Levine et al., 2011; Wulfmeier et al., 2015; Zeng et al., 2022b) further improve the MaxEnt-IRL method so that nonlinear reward can be used.
-For example, a recent paper (Zeng et al., 2022a) proposed a maximum likelihood IRL (ML-IRL) formulation based on the Dynamic Discrete Choice (DDC) model, and nonlinear reward function can be used.
-It is also shown that when the reward function is linearly parameterized, then the MaxEnt is the Lagrangian dual of the ML-IRL problem.
-However, it is worth mentioning that to our best knowledge, almost all IRL-based methods can only utilize demonstration data, which can be problematic because a large amount of high-quality demonstration data is typically hard to obtain.
-Further, it is well-known that only using demonstration cannot extract precise human preference, especially in safety-related tasks where the boundaries between permissible and impermissible actions need to be precisely determined (Liao and Sun, 2024); see, e.g., (Fischer et al., 2021) which shows that insufficient demonstration dataset could lead to high generalization error.
-2.4 Reward Learning from Diverse Data
-For a more precise alignment of agent behavior with human intention, it is crucial to leverage diverse data source such as both expert demonstrations and human feedback to achieve alignment.
-In (Ibarz et al., 2018), the authors first combine two approaches to learn from human feedback: expert demonstrations and trajectory preferences.
+Subsequent works such as Levine et al.
+(2011); Wulfmeier et al.
+(2015); Zeng et al.
+(2022b) further improve the MaxEnt-IRL method so that nonlinear reward can be used.
+For example, Zeng et al.
+(2022a) proposed a maximum likelihood IRL (ML-IRL) formulation based on the Dynamic Discrete Choice (DDC) model, and a nonlinear reward function is utilized.
+It is also shown that when the reward function is linearly parameterized, then the MaxEnt is the Lagrangian dual of the ML-IRL problem (Zeng et al., 2022a).
+On the other hand, it is worth mentioning that to our best knowledge, almost all IRL-based methods can only utilize demonstration data, which can be problematic because a large amount of high-quality demonstration data is typically hard to obtain.
+Further, it is well-known that using demonstration only cannot extract precise human preference, especially in safety-related tasks where the boundaries between permissible and impermissible actions need to be precisely determined; see, e.g., Fischer et al.
+(2021) which shows that insufficient demonstration dataset could lead to high generalization error.
+2.4 Joint Learning from demonstration and preference
+Combining data from demonstrations and human feedback to achieve alignment has also been studied in the robotics literature.
+In Ibarz et al.
+(2018), the authors first combine two approaches to learn from human feedback: expert demonstrations and trajectory preferences.
 The addition of demonstrations to learning from preferences typically results in substantial performance gains compared with using either demonstrations or preferences in isolation.
-In (Palan et al., 2019) and (Bıyık et al., 2022), the authors integrate diverse sources of human feedback including demonstrations and pairwise comparisons in a Bayesian approach to learn reward functions that are assumed to be linear in carefully selected features and evaluate their proposed method on robot learning platform.
+In Palan et al.
+(2019) and Bıyık et al.
+(2022), the authors integrate diverse sources of human feedback including demonstrations and pairwise comparisons in a Bayesian approach to learn reward functions that are assumed to be linear in carefully selected features and evaluate their proposed method on robot learning platform.
 Moreover, their proposed methods need to actively generate preference queries, which are expensive to collect in practical applications.
-In contrast, our proposed approach (to be presented shortly) is not Bayesian and does not include the requirement that the reward model is linear in pre-selected features.
+In contrast, the approach proposed in this paper is not Bayesian and does not include the requirement that the reward model is linear in pre-selected features.
+2.5 Other Approaches to Alignment
+Other approaches to alignment include Direct Preference optimization (DPO) (Rafailov et al., 2023) and Inverse Preference Learning (IPL) (Hejna & Sadigh, 2023) both remove the need for explicit reward modeling, and they directly extract the policy from preferences.
+This greatly reduced the training complexity, but it has been observed that these algorithms can be unstable in the training process (Azar et al., 2023; Xu et al., 2024).
+There is also a large number of works that aim to learn reward functions from rating (Daniel et al., 2014) or ranking (Yuan et al., 2023; Myers et al., 2022).
 Hong et al.
-(2024) proposed a single-stage supervised learning algorithm ORPO that can perform supervised fine-tuning and preference alignment in one training session without maintaining multiple models simultaneously.
-However, Hong et al.
-(2024) can suffer from overoptimization due to its supervised learning pipeline Liu et al.
-(2024).
-3 Proposed Approach: Alignment with Integrated Human Feedback (AIHF)
-Figure 1: Comparison of the RLHF with the proposed AIHF.
-3.1 Preliminaries
-In this section, we develop our proposed formulation for the alignment problem.
-To flesh out our ideas, let us start with a brief discussion of the RLHF approach.
-First, the SFT stage maximizes the following likelihood function:
-$\ell_{\rm SFT}(\pi):=\mathbb{E}_{\tau\sim\pi^{\rm E}}\Big{[}\log\prod_{t\geq 0}\Big{(}\pi(a_{t}|s_{t})\Big{)}^{\gamma^{t}}\Big{]}.$ (1)
-For the RM stage, let $R(\tau;\theta):=\sum_{t\geq 0}\gamma^{t}r(s_{t},a_{t};\theta)$ denote the (parameterized) discounted cumulative reward of one trajectory $\tau$.
-Following the Bradley-Terry-Luce (BTL) model (Bradley and Terry, 1952), the human preference distribution for one pair of trajectories can be stipulated as (where $\sigma(\cdot)$ is the sigmoid function):
-$P[\tau_{l}\prec\tau_{w}]=\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}.$ (2)
-Based on (2), this stage maximizes the following:
-$\ell_{\rm RM}(\theta):=\mathbb{E}_{(\tau_{l}\prec\tau_{w})\in\pi^{P}}\Big{[}\log\Big{(}\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}\Big{)}\Big{]}.$ (3)
-Finally, the RL stage fine-tunes the SFT policy by solving the following policy optimization problem:
-$\max_{\pi}\;\ell_{\rm RL}(\pi;R(\cdot;\theta)):=\mathbb{E}_{s_{0}\sim\rho,\tau\sim\pi}\bigg{[}R(\tau;\theta)+\sum_{t\geq 0}\gamma^{t}c(\pi)\bigg{]}$ (4)
-where $c(\pi)$ is a penalty function, which can take the form of, e.g., the KL divergence $c(\pi):={D}_{\rm KL}\Big{(}\pi(\cdot|s_{t})\|\pi_{\rm SFT}(\cdot|s_{t})\Big{)}$.
-This penalty ensures the resulting policy is close to the original SFT policy.
-We have the following remark about the above problem.
-Remark 3.1
-(Optimal Policy for RLHF) Let $\pi_{\theta}$ denote the optimal policy for problem (4), where $\theta$ is used to parameterize the reward function.
-When $c(\pi):={D}_{\rm KL}\Big{(}\pi(\cdot|s_{t})\|\pi_{\rm SFT}(\cdot|s_{t})\Big{)}$, the KL-regularized MDP problem described by (4) has a closed-form solution:
-$\pi_{\theta}(a|s)=\frac{\pi_{\rm SFT}(a|s)\exp(Q_{\theta}(s,a)/\beta)}{\sum_{\tilde{a}}\pi_{\rm SFT}(\tilde{a}|s)\exp(Q_{\theta}(s,a)/\beta)},$ (5)
-where the corresponding value function $V_{\theta}$ and the Q-function $Q_{\theta}$ are defined as below:
-$V_{\theta}(s):=\mathbb{E}_{\tau\sim\pi_{\theta}}\bigg{[}R(\tau;\theta)+\beta\sum_{t=0}^{\infty}\gamma^{t}{D}_{\rm KL}\Big{(}\pi(\cdot|s_{t})\|\pi_{\rm SFT}(\cdot|s_{t})\Big{)}\bigg{|}s_{0}=s\bigg{]}$ (6a)
-$Q_{\theta}(s,a):=r(s,a;\theta)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}\big{[}V_{\theta}(s^{\prime})\big{]}.$ (6b)
-Further, assuming that $T=1$, i.e., $\tau=(s_{0},a_{0})$, and considering the LLM alignment problem as a sequence-level training problem (this is a popular simplification in language models, see, e.g., Rafailov et al.
-(2023)), the closed-form expression of $\pi_{\theta}$ (5) can be reduced to:
-$\pi_{\theta}(a_{t}|s_{t})=\frac{\pi_{\rm SFT}(a_{t}|s_{t})\exp(\frac{1}{\beta}r(a_{t},s_{t};\theta))}{\sum_{a_{t}\in A}\big{(}\pi_{\rm SFT}(a_{t}|s_{t})\exp(\frac{1}{\beta}r(a_{t},s_{t};\theta))\big{)}}.$ (7)
-Remark 3.2
-(Interpretation of the RLHF Policy) Let us take a closer look at (7).
-In this expression, the results of learning from demonstration data (i.e., $\pi_{\rm SFT}(\cdot)$) and from the preference data (i.e., $r(\cdot;\theta)$) are combined together in a product form to yield the final policy estimate.
-Such a form, though convenient, can result in a number of issues.
-One such an issue is that, the reward $R(\tau;\theta)$ is learned only through the preference dataset $\mathcal{P}$.
-Despite the fact that problem (3) is relatively simple to optimize, such an approach has the following issues: 1) the high-quality demonstration data has been ignored in shaping the reward model; 2) in the situation where the preference data has limited coverage or is of low-quality, the learned reward model is likely to be of low quality as well.
-Here, low-quality preference data arises when, e.g., the trajectories are generated from low-quality models, or the preference annotation is noisy.
-To see the impact on the noisy annotation, consider, for example, the case where a given $s_{t}$, suppose there are two actions $\hat{a}_{t}$ and $\tilde{a}_{t}$, and they are equally good.
-The SFT data respects the ground truth, and the learned SFT policy says $\pi_{\rm SFT}(\tilde{a}_{t}|s_{t})\approx\pi_{\rm SFT}(\hat{a}_{t}|s_{t})$.
-However, due to labeling noise, we have a preference data $(s_{t},\hat{a}_{t})\prec(s_{t},\tilde{a}_{t})$, therefore we have learned $R(a_{t},\tilde{a}_{t};\theta)>R(a_{t},\hat{a_{t}};\theta)$, yielding significant difference in the final policy.
+(2024) proposed a single-stage supervised learning algorithm ORPO that can perform supervised fine-tuning and preference alignment in one training session without maintaining.
+However, all of these works highly rely on high-quality human feedback, which is often more difficult and expensive to obtain.
+3 Alignment with Integrated Human Feedback (AIHF)
+As mentioned before, the reward model obtained in (3) fails to exploit the information about human users’ preferences that are implicit in demonstration data.
+As a result, the fine-tuned model obtained with RLHF may exhibit unsatisfactory alignment performance (this phenomenon will be discussed more concretely in Sec.
+3.4).
+Below we introduce a new approach to jointly train reward and policy models by simultaneously leveraging demonstration and human feedback data.
+Figure 1: Comparison of the RLHF (left) with the proposed AIHF (right).
 3.1 A Meta-Formulation
-The above discussion strongly motivates us to develop an approach that can model the entire alignment process with a common parametrization for both policy and reward models.
-Towards this end, let us consider the following meta-formulation:
-${\bf(AIHF)}\quad\max_{\theta}$ $\quad L(\theta):=\alpha L_{1}(\pi_{\theta})+L_{2}(R(\cdot;\theta))$ (8a)
+Towards developing an approach that can model the entire alignment process with a common parametrization for both policy and reward models, consider the following meta-formulation, termed Alignment with Integrated Human Feedback (AIHF):
+${\bf(AIHF)}\quad\max_{\theta}$ $\quad L(\theta):=w_{1}L_{1}(\pi_{\theta})+L_{2}(R(\cdot;\theta))$ (5a)
 ${\rm s.t.}$ $\quad\pi_{\theta}:=\arg\max_{\pi}~{}L_{3}(\pi;R(\cdot;\theta))$ (5b)
-where $\theta\in\mathbb{R}^{d}$ is a parameter; $L_{1}(\pi_{\theta})$ is a measure of fit of the parametrized policy $\pi_{\theta}$ to demonstration data and $L_{2}(R(\cdot;\theta))$ is a measure of fit of the parametrized reward model $R(\cdot;\theta)$ to the preference data and $L_{3}(\pi,R(\cdot;\theta))$ is a measure of performance of policy $\pi$ with respect to reward model $R(\cdot;\theta)$.
-and $\alpha\geq 0$ is one balancing coefficient reflecting the relative size of demonstration versus preference data.
-We refer the above formulation as Alignment with Integrated Human Feedback (AIHF); see Fig.
-1 for an illustration of the proposed formulation.
+where $\theta\in\mathbb{R}^{d}$ is a parameter; $L_{1}(\pi_{\theta})$ is a measure of fit of the parameterized policy $\pi_{\theta}$ to demonstration data and $L_{2}(R(\cdot;\theta))$ is a measure of fit of the parameterized reward model $R(\cdot;\theta)$ to the preference data and $L_{3}(\pi,R(\cdot;\theta))$ is a measure of performance of policy $\pi$ with respect to reward model $R(\cdot;\theta)$.
+$w_{1}\geq 0$ is one balancing coefficient reflecting the relative size of demonstration versus preference data.
+Note that in the lower level policy optimization, the optimal policy corresponding to the reward model $R(\cdot;\theta)$ is actually determined by the reward parameter $\theta$, where we can denote the optimal policy under the reward model $R(\cdot;\theta)$ as $\pi^{*}_{R_{\theta}}:=\pi_{\theta}$.
+Therefore we simply put $\pi_{\theta}$ since the optimal policy under a certain reward model $R(\cdot;\theta)$ is actually determined by the reward parameters $\theta$.
+See Fig.
+1 for an illustration of AIHF.
 The AIHF (5) is a meta-problem that models the alignment problem.
-It has two levels: an upper-level problem in which the goal is to find policy and reward models that jointly maximize a measure of fit to demonstrations and preference datasets and a lower-level problem which ensures that the policy model optimizes performance with respect to the reward model.
+It has two levels: an upper-level problem in which the goal is to find policy and reward models that jointly maximize a measure of fit to demonstrations and preference datasets; and a lower-level problem which ensures that the policy model optimizes performance with respect to the reward model.
 Its components can be customized to yield specific alignment formulations and algorithms.
-But before diving into various customizations, let us discuss the advantages of this formulation.
+Before diving into various customizations, let us discuss the advantages of this formulation.
 Generality.
-One can specialize the choice of loss functions and problem parameters, to yield a number of existing alignment formulations.
+One can specialize the loss functions and problem parameters to yield a number of existing alignment formulations.
 Such generality implies that algorithms developed for (5) are easily applicable to different special formulations it covers.
 For more details see Sec.
-3.4.
+3.3.
 Joint optimization.
@@
 See Sec.
-5 for a detailed discussion on how the reward parameter $\theta$ is updated by leveraging such demonstration.
+4 for a detailed discussion on how the reward parameter $\theta$ is updated by leveraging such demonstration, and see Tab.
+5 for the experimental comparison between the reward model learned by RLHF and by our AIHF (5).
 Dataset Integration.
-Clearly, the reward learning process leverages all the available data, therefore, we can expect that a high-quality reward model can still be obtained even under unfavorable situations where certain category of data is scarce or of low quality.
+Clearly, the reward learning process leverages all the available data, therefore, we can expect that a high-quality reward model and its induced optimal policy can still be obtained even under unfavorable situations where the preference data is not sufficient.
+See Sec.
+5 for experimental evidences.
 3.2 Specification of AIHF
@@
 Note that we aim to optimize the reward parameter $\theta$ to align with human feedback in (5a), thus the objective of $L_{1}$ can be specialized as a maximum likelihood function over expert demonstrations as below:
-$L_{1}(\pi_{\theta}):=\mathbb{E}_{\tau\sim\pi^{\rm E}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\log\pi_{\theta}(a_{t}|s_{t})\Big{]}.$ (9)
-Here $\pi_{\theta}$ optimizes the measure of performance $L_{3}(\pi;R(\cdot;\theta))$ for a reward model $R(\cdot;\theta)$ as below:
+$L_{1}(\pi_{\theta}):=\mathbb{E}_{\tau\sim\mathcal{D}}\Big{[}\log\prod_{t\geq 0}\Big{(}\pi_{\theta}(a_{t}|s_{t})\Big{)}^{\gamma^{t}}\Big{]}=\mathbb{E}_{\tau\sim\mathcal{D}}\Big{[}\sum_{t\geq 0}\gamma^{t}\log\pi_{\theta}(a_{t}|s_{t})\Big{]}.$ (6)
+Here $\pi_{\theta}$ optimizes the measure of performance $L_{3}(\pi;R(\cdot;\theta))$ for a reward model $R(\cdot;\theta)$ as:
 $L_{3}(\pi;R(\cdot;\theta)):=\mathbb{E}_{s_{0}\sim\rho,\tau\sim\pi}\bigg{[}R(\tau;\theta)-\beta\sum_{t\geq 0}\gamma^{t}{D}_{\rm KL}\Big{(}\pi(\cdot|s_{t})\|\pi^{0}(\cdot|s_{t})\Big{)}\bigg{]}$ (7)
@@
 Thus, the optimal distribution $\mu_{\theta}$ over the set of $T$-long sequence of state-action pairs is defined as follows:
-$\mu_{\theta}:=\arg\max_{\mu\in\Delta^{T}}\mathbb{E}_{\tau\sim\mu}\big{[}R(\tau;\theta)-\beta\mathcal{D}_{KL}(\mu||\mu^{0})\big{]}$
+$\mu_{\theta}:=\arg\max_{\mu\in\Delta^{T}}\mathbb{E}_{\tau\sim\mu}\big{[}R(\tau;\theta)-\beta\mathcal{D}_{\rm KL}(\mu||\mu^{0})\big{]}$
 where $\Delta_{T}$ denotes the simplex on $\mathcal{H}_{T}$ and $\mu^{0}$ is a prior distribution on the trajectories. It can be shown the solution of the above problem is of the form:
@@
 With this result, we can now obtain a model for the likelihood that sequence $\tau_{j}$ is preferred over $\tau_{j}$.
-By the independence of irrelevant alternatives property (Fudenberg et al., 2015) of the optimal choice $\mu_{\theta}$, when the set of feasible choices is reduced from $\mathcal{H}_{T}$ to just the the two-tuple $\{\tau_{l},\tau_{w}\}$, the likelihood that sequence $\tau_{w}$ is preferred over $\tau_{l}$ is given by $\mathbb{P}_{\theta}(\tau_{l}\prec\tau_{w}):=\frac{\mu_{\theta}(\tau_{w})}{\mu_{\theta}(\tau_{l})+\mu_{\theta}(\tau_{w})}.$ This motivates the choice of $L_{2}(\theta)$ as the following likelihood function:
-$L_{2}(R(\cdot;\theta))=\mathbb{E}_{(\tau_{l}\prec\tau_{w})\in\pi^{P}}\Big{[}\log\frac{\mu_{\theta}(\tau_{w})}{\mu_{\theta}(\tau_{w})+\mu_{\theta}(\tau_{l})}\Big{]}=\mathbb{E}_{(\tau_{l}\prec\tau_{w})\in\pi^{P}}\Big{[}\log\frac{\mu^{0}(\tau_{w})\exp\big{(}R(\tau_{w};\theta)\big{)}}{\mu^{0}(\tau_{w})\exp\big{(}R(\tau_{w};\theta)\big{)}+\mu^{0}(\tau_{l})\exp\big{(}R(\tau_{l};\theta)\big{)}}\Big{]}.$
+By the independence of irrelevant alternatives property (Fudenberg et al., 2015) of the optimal choice $\mu_{\theta}$, when the set of feasible choices is reduced from $\mathcal{H}_{T}$ to just the the two-tuple $\{\tau_{l},\tau_{w}\}$, the likelihood that sequence $\tau_{w}$ is preferred over $\tau_{l}$ is given by $\mathbb{P}_{\theta}(\tau_{w}\succ\tau_{l}):=\frac{\mu_{\theta}(\tau_{w})}{\mu_{\theta}(\tau_{l})+\mu_{\theta}(\tau_{w})}.$ This motivates the choice of $L_{2}(\theta)$ as the following likelihood function:
+$L_{2}(R(\cdot;\theta))$ $=\mathbb{E}_{(\tau_{w}\succ\tau_{l})\in\mathcal{P}}\Big{[}\log\frac{\mu_{\theta}(\tau_{w})}{\mu_{\theta}(\tau_{w})+\mu_{\theta}(\tau_{l})}\Big{]}$
+$=\mathbb{E}_{(\tau_{w}\succ\tau_{l})\in\mathcal{P}}\Big{[}\log\frac{\mu^{0}(\tau_{w})\exp\big{(}R(\tau_{w};\theta)\big{)}}{\mu^{0}(\tau_{w})\exp\big{(}R(\tau_{w};\theta)\big{)}+\mu^{0}(\tau_{l})\exp\big{(}R(\tau_{l};\theta)\big{)}}\Big{]}.$
 With $\mu^{0}$ equal to the uniform distribution on $\mathcal{H}_{T}$, this model is equivalent to the BTL model (3):
-${L}^{\rm BTL}_{2}(\theta)=\ell_{\rm RM}(\theta)=\mathbb{E}_{(\tau_{l}\prec\tau_{w})\in\pi^{P}}\Big{[}\log\Big{(}\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}\Big{)}\Big{]}.$ (11)
+${L}^{\rm BTL}_{2}(\theta)=\ell_{\rm RM}(\theta)=\mathbb{E}_{(\tau_{w}\succ\tau_{l})\in\mathcal{P}}\Big{[}\log\Big{(}\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}\Big{)}\Big{]}.$ (8)
 3.3 Special Cases of AIHF
-Next we discuss how formulation (8) can be specialized to some of the known alignment algorithms.
-Specialization to RLHF-Type Approach. First, if we set the coefficient $\alpha=0$ in (8), we obtain the following problem:
+Next, we discuss how formulation (5) can be specialized to some of the known alignment algorithms.
+Specialization to RLHF-Type Approach. First, if we set the coefficient $w_{1}=0$ in (5), we obtain:
 $\begin{split}\max_{\theta}&\quad L_{2}(\theta)\;\;{\rm s.t.}\;\;\pi_{\theta}:=\arg\max_{\pi}L_{3}(\pi;R(\cdot;\theta)).\end{split}$ (9)
@@
 Specialization to DPO-Type Approach.
-Let us consider the relationship between our formulation (8) with the DPO-type approaches.
-Let us set the following objective function $L_{1}=\ell_{\rm SFT}$ and $L_{2}=\ell_{\rm RM}$, and assume that $T=1$ for th generation process.
-Through relaxing the constraint which ensures the policy is optimal w.r.t.
+Consider the relationship between formulation (5) with the DPO-type approaches.
+Let us set the following objective function $L_{1}=\ell_{\rm SFT}$ and $L_{2}=\ell_{\rm RM}$, and assume that $T=1$ for the generation process.
+Relaxing the constraint (5b) which ensures the policy is optimal w.r.t.
 a certain parameterized model, we can obtain a DPO-type formulation:
-$\max_{\pi}~{}L(\pi):=\alpha\cdot\mathbb{E}_{\tau^{\rm E}\sim\pi^{\rm E}}\bigg{[}\log\pi(a^{\rm E}|s^{\rm E})\bigg{]}+\mathbb{E}_{(\tau_{i}\prec\tau_{j})\sim\pi^{P}}\bigg{[}\sigma\big{(}\beta\log\frac{\pi(a_{j}|s_{j})}{\pi^{0}(a_{i}|s_{j})}-\beta\log\frac{\pi(a_{i}|s_{i})}{\pi^{0}(a_{j}|s_{i})}\big{)}\bigg{]}.$ (13)
+$\max_{\pi}~{}L(\pi):=w_{1}\cdot\mathbb{E}_{\tau^{\rm E}\sim\pi^{\rm E}}\bigg{[}\log\pi(a^{\rm E}|s^{\rm E})\bigg{]}+\mathbb{E}_{(\tau_{j}\succ\tau_{i})\sim\pi^{P}}\bigg{[}\log\bigg{(}\sigma\big{(}\beta\log\frac{\pi(a_{j}|s_{j})}{\pi^{0}(a_{i}|s_{j})}-\beta\log\frac{\pi(a_{i}|s_{i})}{\pi^{0}(a_{j}|s_{i})}\big{)}\bigg{)}\bigg{]}.$
 The above formulation specializes to Liu et al.
 (2024), which is a slightly generalized version of DPO when both demonstration and preference data are used.
-Setting $\alpha=0$ reduces to the problem solved by DPO; see Rafailov et al.
-(2023), Eq.
-(2).
-4 Why AIHF can outperform two-stage alignment approaches?
-To understand the difference between the proposed approach and the two-stage approach of the standard alignment pipeline, let us consider the following thought process over a simplified, but informative setting.
-Let us consider the simplified case where there are $N$ discrete actions, that is, the action set is $A:=\{\tau_{1},\tau_{2},\cdots\tau_{N}\}$.
-Assume that there is a reward function $r(\cdot):A\mapsto\mathbb{R}$ (note, that here the parameterization of the reward has been omitted for simplicity).
-Let us compare the optimal solutions for different alignment algorithms and get some intuition about each approach.
-Throughout the process, we will assume that the reward function is not parameterized, and we directly learn the optimal reward value.
-Indeed this is a simplification needed to understand the structure of the solution, but we argue that this is without loss of generality, since if the reward model is sufficiently parameterized by neural networks, then these optimal reward values can always be recovered by proper training.
-Consider the following decision-making model:
-${\pi^{*}:=\arg\max_{\pi\in\Delta^{N}}\mathbb{E}_{\tau\sim\pi(\cdot)}\Big{[}r(\tau)-\beta\log\pi(\tau)\Big{]},}$ (14)
-where $\Delta^{N}$ is the probability simplex and $\beta>0$. It can be shown that the solution to (14) is:
-$\pi^{*}_{i}(r)=\frac{\exp(\frac{r_{i}}{\beta})}{\sum_{j}\exp(\frac{r_{j}}{\beta})},\quad\mbox{\rm where}\quad r_{i}:=r(\tau_{i}),\quad\forall~{}i.$ (15)
-Reward Estimation via Demonstration Data: When demonstration data $\mathcal{D}:=\{\tau_{n}\}$ is available where each $\tau_{n}\in A,\;\forall~{}n$, the estimation problem can be written as:
-$\widehat{r}_{\mathcal{D}}:=\arg\max_{r}\ell_{\mathcal{D}}(r)$
-where $\ell_{\mathcal{D}}(r):=\mathbb{E}_{\tau_{i}\sim\mathcal{D}}[\log\pi^{*}_{i}(r)]$. By (15) it follows that
-$\ell_{\mathcal{D}}(r)=\frac{1}{\beta}\mathbb{E}_{\tau_{i}\sim\mathcal{D}}\big{[}r_{i}\big{]}-\log\sum_{j}\exp(\frac{r_{j}}{\beta}).$
-Hence, it is easy to verify the gradient is:
-$\frac{\partial\ell_{\mathcal{D}}(r)}{\partial r_{i}}$ $=\frac{1}{\beta}\Big{(}\frac{\#\tau_{i}~{}\mbox{in}~{}\mathcal{D}}{|\mathcal{D}|}-\pi^{*}_{i}(r)\Big{)},~{}\forall~{}i.$ (16)
-and the second order derivatives are:
-$\frac{\partial^{2}\ell_{\mathcal{D}}(r)}{\partial r_{i}^{2}}=-\frac{1}{\beta^{2}}\pi^{*}_{i}(r)(1-\pi^{*}_{i}(r))$ $~{}~{}~{}~{}~{}\frac{\partial^{2}\ell_{\mathcal{D}}(r)}{\partial r_{i}\partial r_{j}}=\frac{1}{\beta^{2}}\pi^{*}_{i}(r)\pi^{*}_{j}(r).$ (17)
-Therefore, we can observe that setting the gradient to zero, the first-order stationary solution (in fact, in this case it is optimal solution since the Hessian is negative definite) indicates that the estimate of the probability of choosing action $i$ is equal to the observed frequency of action $\tau_{i}$ appears in $\mathcal{D}$.
-Note that in our simplified case, the above estimator can be viewed as the counterpart for the SFT estimator since it is the best that one can get using only demonstration data.
-Reward Estimation with Preferences Data: Suppose that the preference data is available, $\mathcal{P}:=\{(\tau_{i}\prec\tau_{j})\}$, where again $\tau_{i},\tau_{j}\in A$.
-Let us understand how to analyze the optimal policy $\pi^{*}$.
-By IIA (independence of irrelevant alternatives) (Fudenberg et al., 2015), when the decision maker is asked to choose between only two alternatives $\{\tau_{i},\tau_{j}\}$, we have :
-$\Pr(\tau_{j}\prec\tau_{i})$ $=\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}.$
-The reward estimation problem is defined as:
-$\widehat{r}_{\mathcal{P}}\in\arg\max_{r}\ell_{\mathcal{P}}(r)$ (18)
-where $\ell_{\mathcal{P}}(r):=\mathbb{E}_{(\tau_{j}\prec\tau_{i})\sim\mathcal{P}}\Big{[}\log\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\Big{]}$.
-The partial derivative of the log-likelihood is given by:
-$\frac{\partial\ell_{\mathcal{P}}(r)}{\partial r_{i}}$ $=\frac{1}{\beta}\sum_{j\neq i}\big{(}\frac{\#\{\tau_{j}\prec\tau_{i}~{}\mbox{in}~{}\mathcal{P}\}}{|\mathcal{P}_{i,j}|}-\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\big{)}\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}.$ (19)
-The first order condition can be written as:
-$\pi^{*}_{i}(r)$ $=\frac{\sum_{j\neq i}\{\#a_{i}\succ a_{j}~{}\mbox{in}~{}\mathcal{P}\}}{\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho^{*}_{-(i,j)}(r)}$ (20)
-where $\rho^{*}_{-(i,j)}(r):=\Big{(}1-\sum_{k\in A\backslash\{i,j\}}\pi^{*}_{k}(r)\Big{)}^{-1}$ is the expected number of times an action other than $\tau_{i}$ or $\tau_{j}$ is selected when sampling $\pi^{*}(r)$ infinitely many times.
-The second-order derivatives are given by
-$\frac{\partial^{2}\ell_{\mathcal{P}}(r)}{\partial r_{i}^{2}}=-\frac{1}{\beta^{2}}\sum_{j\neq i}\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\frac{\pi^{*}_{j}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}$ $~{}~{}\frac{\partial^{2}_{r_{i}}\ell_{\mathcal{P}}(r)}{\partial r_{j}\partial r_{i}}=\frac{1}{\beta^{2}}\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\frac{\pi^{*}_{j}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}.$
-Note that compared with the previous case with only demonstration data, here the first-order solution does not imply global optimality since the Hessian is negative semi-definite.
-Next, by using the above analysis, let us understand the policies that can be obtained by different alignment algorithms.
-SFT Estimator. In this case, it is easy to observe that the optimal policy is given by
-$\pi^{\rm SFT}_{i}=\frac{\#\tau_{i}~{}\mbox{in}~{}\mathcal{D}}{|\mathcal{D}|}={\pi^{*}(\widehat{r}_{\mathcal{D}})},~{}\forall~{}i.$ (21)
-RLHF Policy. The RLHF policy is defined as follows:
-$\pi^{\rm RLHF}=\arg\max_{\pi\in\Delta^{N}}~{}~{}\mathbb{E}_{\tau_{i}\sim\pi}\big{[}{\widehat{r}_{\mathcal{P}}}(\tau_{i})\big{]}-\beta\mathrm{KL}(\pi||\pi^{\rm SFT})$
-where $\widehat{r}_{\mathcal{P}}$ is the estimator obtained from preferences (i.e.
-solution to (20)), $\pi^{\rm SFT}$ is a prior distribution obtained from the SFT model trained with demonstration dataset $\mathcal{D}$, and $\Delta^{N}$ is the probability simplex.
-It follows that
-$\pi^{\rm RLHF}(\tau_{i})$ $=\frac{\pi^{\rm SFT}(\tau_{i})\exp\frac{1}{\beta}\widehat{r}_{\mathcal{P}}(\tau_{i})}{\sum_{j=1}^{n}\pi^{\rm SFT}(\tau_{j})\exp\frac{1}{\beta}\widehat{r}_{\mathcal{P}}(\tau_{j})}$
-$={\frac{\exp\frac{1}{\beta}\widehat{r}_{\mathcal{D}}(\tau_{i})\exp\frac{1}{\beta}\widehat{r}_{\mathcal{P}}(\tau_{i})}{\sum_{j=1}^{n}\exp\frac{1}{\beta}\widehat{r}_{\mathcal{D}}(\tau_{j})\exp{\frac{1}{\beta}\widehat{r}_{\mathcal{P}}}(\tau_{j})}}$
-$={\frac{\exp\big{(}\frac{1}{\beta}(\widehat{r}_{\mathcal{D}}(\tau_{i})+\widehat{r}_{\mathcal{P}}(\tau_{i}))\big{)}}{\sum_{j=1}^{n}\exp\big{(}\frac{1}{\beta}(\widehat{r}_{\mathcal{D}}(\tau_{j})+\widehat{r}_{\mathcal{P}}(\tau_{j}))\big{)}}}$
-$=\pi^{*}_{i}\Big{(}\widehat{r}_{\mathcal{D}}+\widehat{r}_{\mathcal{P}}\Big{)}$ (22)
-That is, the RLHF policy can be seen as the softmax policy for the sum of rewards estimator.
-AIHF estimator: In AIHF, assume that the lower-level problem uses (10), and the prior $\pi^{0}$ is uniform, then the optimal $\pi^{\rm AIHF}$ follows the same expression as in (15).
-One can then plug in such a solution to its loss function in (8a), and obtain the following reward learning problem:
-$\widehat{r}^{\rm AIHF}=\arg\max_{r}\quad\ell_{\mathcal{D}+\mathcal{P}}(r):=|\mathcal{D}|L_{1}(r)+|\mathcal{P}|L_{2}(r)$ (23)
-where $L_{1}(r):=\mathbb{E}_{\tau_{i}\sim\mathcal{D}}[\log\pi^{*}_{i}(r)]$ and $L_{2}(r):=\mathbb{E}_{(\tau_{j}\prec\tau_{i})\sim\mathcal{P}}[\log\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}]$.
-Using (16) and (19), we obtain:
-$\frac{\partial\ell_{\mathcal{D}+\mathcal{P}}(r)}{\partial r_{i}}$ =#$=\frac{\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}}{|\mathcal{D}|}|\mathcal{D}|+\sum_{j\neq i}\#\{\tau_{j}\prec\tau_{i}~{}\mbox{in}~{}\mathcal{P}\}-\pi^{*}_{i}|\mathcal{D}|-\sum_{j\neq i}\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}|\mathcal{P}_{i,j}|.$
-Hence, the first-order condition is:
-#$\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}+\sum_{j\neq i}\#\{\tau_{j}\prec\tau_{i}~{}\mbox{in}~{}\mathcal{P}\}$ $=\pi^{*}_{i}(r)(|\mathcal{D}|+\sum_{j\neq i}|\mathcal{P}_{i,j}|\frac{1}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)})$
-Or equivalently,
-$\pi^{*}_{i}(r)$ =#$=\frac{\#\{a_{i}~{}\mbox{in}~{}\mathcal{D}\}+\sum_{j\neq i}\#\{a_{j}\prec a_{i}~{}\mbox{in}~{}\mathcal{P}\}}{|\mathcal{D}|+\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho^{*}_{-(i,j)}(r)}$ (24)
-Example 1: With $\beta=1$ and only two actions $\tau_{1}$ and $\tau_{2}$. Since $\rho^{*}(r)_{-(1,2)}=1$, it follows that:
-$\pi^{AIHF}_{1}:=\pi^{*}_{1}(\widehat{r}^{AIHF})$ =#$=\frac{\#\{\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}+\#\{\tau_{2}\prec\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}}{|\mathcal{D}|+|\mathcal{P}|}$
-$=\frac{|\mathcal{D}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{r}_{\mathcal{D}})+\frac{|\mathcal{P}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})$
-That is the policy estimated obtained by using demonstrations and preferences is a weighted average of the polices estimated by using only demonstrations and only preferences.
-Abusing notation let $\pi^{*}_{1}:=\pi^{*}_{1}(r^{*})$ where $r^{*}$ is the ground-truth reward.
-It follows that $\mbox{Var}(\pi^{*}_{1}(\widehat{r}_{\mathcal{D}}))=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{D}|}$, $\mbox{Var}(\pi^{*}_{1}(\widehat{r}_{\mathcal{P}}))=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{P}|}$ and
-$\mbox{Var}(\pi^{\rm AIHF}_{1})=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{D}|+|\mathcal{P}|}$
-This shows the AIHF policy estimator has less variance than either one of the estimators obtained from demonstrations only or preferences only.
-To further illustrate, suppose $r(\tau_{1})=r(\tau_{2})$ and we have the following datasets:
-#$\#\{\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}=\#\{\tau_{2}~{}\mbox{in}~{}\mathcal{D}\}=50,\quad\#\{\tau_{1}\succ\tau_{2}~{}\mbox{in}~{}\mathcal{P}\}$ $=6,\quad\#\{\tau_{2}\succ\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}=4.$
-With the given data, $\pi^{\rm SFT}_{1}=\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{\#\{\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}}{|\mathcal{D}|}=\frac{50}{100}$ and the solution to (20) yields
-$\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}=\frac{6}{10}$
-Hence, $\pi_{1}^{\rm AIHF}=\frac{100}{10+100}\pi^{*}_{1}(\widehat{r}_{\mathcal{D}})+\frac{10}{10+100}\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{56}{110}$. It follows from (22) that:
-$\pi^{\rm RLHF}_{1}$ $=\frac{\pi^{\rm SFT}_{1}\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\pi^{\rm SFT}_{1}\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\pi^{\rm SFT}_{2}\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}$
-$=\frac{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}=\frac{6}{10}$
-In this example, the RLHF policy estimator is the farthest from ground-truth, because it does not correctly use the information provided by the demonstration data which in this case happens by chance to be correct $\pi^{\rm SFT}_{1}=\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{1}{2}$.
-Consider now the datasets:
-{#$\{\#\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}=6,\quad\{\#\tau_{2}~{}\mbox{in}~{}\mathcal{D}\}=4,\quad\{\#\tau_{1}\succ\tau_{2}~{}\mbox{in}~{}\mathcal{P}\}=\{\#\tau_{2}\succ\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}=50$ (25)
-In this case, $\pi^{\rm SFT}_{1}=\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{6}{10}$ and the solution to (20) yields
-$\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}=\frac{50}{100}$
-It follows from (22) that:
-$\pi^{\rm RLHF}_{1}$ $=\frac{\pi^{\rm SFT}_{1}\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\pi^{\rm SFT}_{1}\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\pi^{\rm SFT}_{2}\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}$
-$=\frac{\pi_{1}^{SFT}}{\pi_{1}^{SFT}+\pi_{2}^{SFT}}=\frac{6}{10}$
-Hence, $\pi_{1}^{\rm AIHF}=\frac{10}{10+100}\pi^{*}_{1}(\widehat{r}_{\mathcal{D}})+\frac{100}{10+100}\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{56}{110}$.
-In this example, the RLHF policy estimator is again farthest from ground-truth, because it does not correctly dismiss the information provided by the demonstration data which is less informative than preference.
-Numerical Example 2.
-Let us use an illustrative example to show that RLHF method will result in significant data under-utilization when the demonstration coverage is limited.
-Assume that there are 50 actions, each with a ground-truth reward defined by $r_{i}=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{(\frac{i}{50}-\mu)^{2}}{2\sigma^{2}}}$, where $\mu=0.5$ and $\sigma=2$.
-Assume we can sample demonstration and preference from the ground truth reward distribution: demonstrations are sampled from the multinomial distribution, while preferences are sampled from the BTL model.
-In an extreme scenario, let demonstrations cover actions 1 through 45, while preferences have full coverage across all actions.
-In the subsequent experiment, we initially sample 2000 demonstrations using the multinomial distribution $p_{i}=\frac{\exp r_{i}}{\sum^{j=45}_{j=1}\exp r_{j}}$, and obtain 200 preferences for each preference pair with $p(i\succ j)=\frac{\exp r_{i}}{\exp r_{j}+\exp r_{i}}$.
-We then calculate the optimal policy by solving the previously mentioned first order conditions, yielding the following result.
-Figure 2: The optimal policy of RLHF, SFT, AIHF, and Ground-truth distribution.
-The left region of the red dotted line is included in the demonstration, while the right region is uncovered.
-We report the results with 100 random repeats.
-From the result shown in Figure 2, we demonstrate that both SFT and RLHF transfer the weight from uncovered action to covered action when demonstration coverage is limited, as indicated by $\pi_{SFT}(a_{uncovered})=0$.
-Consequently, the weight of covered actions is significantly higher than the ground truth.
-However, this issue does not occur when jointly optimizing the demonstration and preference in the AIHF method.
+Setting $w_{1}=0$ reduces to the problem solved by DPO; see Rafailov et al.
+(2023, Eq.
+(2)), We refer to the algorithm defined by Eq.
+(5) as Direct AIHF.
+Specialization to Self-Play Approach.
+Define $\ell(\cdot)$ as a monotonic and convex loss function, consider setting $L_{1}:=w_{1}\cdot\mathbb{E}_{\tau^{E}\in\pi^{E},\alpha\in\pi(.|s^{E})}\ell\left(R(\tau^{E};\theta)-R(\tau;\theta)\right)$, and setting $L_{2}$ and $L_{3}$ according to (8) and (7), respectively.
+Note that the choice of $L_{1}$ means that given demonstration data, we will find a policy that generates trajectories that match the rewards of the demonstration data.
+Again using the DPO type of reformulation, by substituting the reward expression obtained from the optimal policy (42) to $L_{1}$ and selecting the $\sigma(\cdot)$ as $\ell(\cdot)$, then the AIHF problem in this case becomes:
+$\max_{\pi}L(\pi):=$ $w_{1}\mathbb{E}_{\tau^{\rm E}\sim\pi^{\mathrm{E}},\tilde{a}\sim\pi(\cdot|s)}\left[\log\sigma\left(\beta\log\frac{\pi(a^{\rm E}|s^{\rm E})}{\pi^{0}(a^{\rm E}|s^{\rm E})}-\beta\log\frac{\pi(\tilde{a}|s^{\rm E})}{\pi^{0}(\tilde{a}|s^{\rm E})}\right)\right]$
+$+\mathbb{E}_{(\tau_{j}\succ\tau_{i})\sim\pi^{P}}\bigg{[}\log\sigma\left(\beta\log\frac{\pi(a_{j}|s_{j})}{\pi^{0}(a_{j}|s_{j})}-\beta\log\frac{\pi(a_{i}|s_{i})}{\pi^{0}(a_{i}|s_{i})}\right)\bigg{]}.$ (11)
+Note that the first part of the above formulation is similar to what has been proposed in SPIN (Chen et al., 2024), which only utilizes the SFT data.
+We refer to the algorithm defined by Eq.
+(3.3) as Self-Play AIHF.
+3.4 Why AIHF can outperform two-stage alignment approaches?
+To understand the difference between the proposed approach and the successive stages approach of the standard alignment pipeline, let us consider the a static setting with action set is $A:=\{\tau_{1},\tau_{2},\cdots\tau_{N}\}$, reward function $R(\cdot):A\mapsto\mathbb{R}$, and demonstration $\mathcal{D}$ and preference dataset $\mathcal{P}$.
+In what follows, we compare the optimal solutions for policies obtained by different alignment approaches.
+Due to space limitation, all derivation in this section is relegated to Appendix A.2.
+Policy with Demonstration Data.
+It can be easily shown that when only the demonstration data $\mathcal{D}$ is available, the probability of generating $i$-th data equals its empirical probability, i.e., $\pi_{\rm SFT}(\tau_{i})=\frac{\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}}{|\mathcal{D}|}$.
+Assume that such a policy is parameterized by an implicit reward function $R_{D}$, using the following softmax choice model where $\tau_{i}\in A$ is selected with probability $\pi^{*}_{i}(R)=\frac{\exp(R_{i}/\beta)}{\sum_{j=1}^{N}\exp(R_{j}/\beta)}$ where $R_{i}:=R(\tau_{i})$.
+Assuming a reference value $\widehat{R}_{\mathcal{D}}(\tau_{1})=\bar{R}_{1}$, then the optimal rewards satisfies (See Sec.
+A.2.1):
+#$\frac{\#\{\tau_{i}\in\mathcal{D}\}}{|\mathcal{D}|}=\pi^{*}_{i}(\widehat{R}_{\mathcal{D}})=\frac{\exp(\widehat{R}_{\mathcal{D}}(\tau_{i})/\beta)}{\sum_{j=1}^{N}\exp(\widehat{R}_{\mathcal{D}}(\tau_{j})/\beta)}~{}~{}~{}~{}i\in\{2,\dots,N\}.$ (12)
+This implicit reward will be used shortly to characterize the RLHF policy.
+Policy with Preference Only Data. Next, it can be shown that when only the preference data $\mathcal{P}$ is available, the reward estimation problem is defined as:
+$\widehat{R}_{\mathcal{P}}=\arg\max_{R}\ell_{RM}(R):=\mathbb{E}_{(\tau_{i}\succ\tau_{j})\sim\mathcal{P}}\Big{[}\log\frac{\pi^{*}_{i}(R)}{\pi^{*}_{i}(R)+\pi^{*}_{j}(R)}\Big{]}.$ (13)
+Again with a fixed reference value $\widehat{R}_{\mathcal{P}}(\tau_{1})=\bar{R}_{1}$, the solution is (see Sec. A.2.1):
+$\pi^{*}_{i}(\widehat{R}_{\mathcal{P}})$ $=\frac{\sum_{j:j\neq i}|\mathcal{P}_{i\succ j}|}{\sum_{j:j\neq i}|\mathcal{P}_{i,j}|\rho_{-(i,j)}(\pi^{*}(\widehat{R}_{\mathcal{P}}))}$ (14)
+where $|\mathcal{P}_{i\succ j}|:=\#\{\tau_{i}\succ\tau_{j}~{}\mbox{in}~{}\mathcal{P}\}$ and $|\mathcal{P}_{i,j}|:=|\mathcal{P}_{i\succ j}|+|\mathcal{P}_{j\succ i}|$ and $\rho_{-(i,j)}(\pi):=\Big{(}1-\sum_{k\in A\backslash\{i,j\}}\pi_{k}\Big{)}^{-1}$ is the expected number of times an action other than $\tau_{i}$ or $\tau_{j}$ is selected when sampling actions from $\pi$ infinitely many times.
+RLHF Policy.
+Based on the above results, in Sec.
+A.2.2 we show that the RLHF approach has the optimal policy $\pi^{\rm RLHF}(\tau_{i})=\pi^{*}_{i}\Big{(}\widehat{R}_{\mathcal{D}}+\widehat{R}_{\mathcal{P}}\Big{)}$.
+That is, the RLHF policy can be seen as the softmax policy for the sum of reward estimators obtained from demonstrations and preferences separately.
+AIHF Policy. Finally, we also find that the AIHF policy is of the form:
+$\pi^{\rm AIHF}(\tau_{i})=\frac{\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}+\sum_{j\neq i}|\mathcal{P}_{i\succ j}|}{|\mathcal{D}|+\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho_{-(i,j)}\big{(}\pi^{*}(\widehat{R}^{\rm AIHF})\big{)}}.$ (15)
+Discussion.
+Let us summarize our findings.
+First, $\pi^{\rm RLHF}$, which takes the form of the softmax of the sum of two rewards—one learned during the SFT stage and another during the reward training stage—offers some interesting insights into this popular approach.
+Second, $\pi^{\rm AIHF}$ is more robust than $\pi^{\rm RLHF}$.
+To see this, suppose that $|\mathcal{D}|\gg|\mathcal{P}|$, i.e.
+there is more demonstration than preference data.
+In this case, the policy estimator in (15) will be largely defined by the demonstration data (which is reasonable) whereas the RLHF policy can be noisy since it (soft) maximizes the sum of two reward estimators: one that is more accurate (i.e.
+the one based on demonstrations, $\widehat{R}_{\mathcal{D}}$) and one that is less accurate (i.e.
+the one based on preferences $\widehat{R}_{\mathcal{P}}$).
+A similar observation can be made when $|\mathcal{D}|\ll|\mathcal{P}|$; see Sec.
+A.2.2 for more details.
+Third, $\pi^{\rm AIHF}$ can have less variance as compared with $\pi^{\rm RLHF}$.
+Indeed, $\pi^{\rm AIHF}$ takes a form of a weighted average of the policies estimated separately with demonstration and preference data, as by using (12) and (14), we can re-write (15) as:
+$\pi^{*}_{i}(\widehat{R}^{\rm AIHF})=\frac{|\mathcal{D}|}{|\mathcal{D}|+\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho_{-(i,j)}\big{(}\pi^{*}(\widehat{R}^{\rm AIHF})\big{)}}\pi^{*}_{i}(\widehat{R}_{\mathcal{D}})+\frac{\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho_{-(i,j)}\big{(}\pi^{*}(\widehat{R}_{\mathcal{P}})\big{)}}{|\mathcal{D}|+\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho_{-(i,j)}\big{(}\pi^{*}(\widehat{R}^{\rm AIHF})\big{)}}\pi^{*}_{i}(\widehat{R}_{\mathcal{P}})$
+Such averaging entails reduced variance. We also include simple numerical examples in the Appendix A.2.3 to further illustrate this point.
 4 Proposed Algorithm for AIHF Training
 We are now ready to design algorithms for the proposed AIHF formulation (5).
-To begin with, first note that (8) takes a hierarchical form, and it belongs to the class of problem named bi-level optimization, first developed in the 70s (Fiacco and McCormick, 1990), and recently found many applications in machine learning (Wang et al., 2021; Liu et al., 2021, 2022).
+To begin with, first note that (5) takes a hierarchical form, and it belongs to the class of problem named bi-level optimization, first developed in the 70s (Fiacco & McCormick, 1990), and recently found many applications in machine learning (Wang et al., 2021; Liu et al., 2021, 2022).
 Generically speaking, bi-level problems are not easy to optimize; more specifically, in (5), the upper-level problem (5a) is a function of both the lower-level optimal solution $\pi_{\theta}$ and the true parameter $\theta$.
 It follows that a (stochastic) first-order algorithm for $L(\theta)$ involves some (potentially non-trivial) implicit gradient computation which often involves computing the Hessian matrix for the lower-level objective function.
-Fortunately, as we will show shortly, with some special choices of $L_{1}$, $L_{2}$, $L_{3}$, one can design some simple and very efficient algorithms.
-Throughout this section, we are searching for a good policy $\pi_{\theta}$ and a reward estimate $r(\cdot,\cdot;\theta)$ to align with human feedback, where the policy $\pi_{\theta}$ is an optimal solution w.r.t.
-the certain reward estimate $r(\cdot,\cdot;\theta)$ according to the policy optimization problem (3.4).
+Fortunately, as we will show shortly, with some special choices of $L_{1}$, $L_{2}$, and $L_{3}$, one can design some simple and very efficient algorithms.
+Before we go to details, we note that throughout this section, we assume that we are searching for a good policy $\pi_{\theta}$ and a reward estimate $r(\cdot,\cdot;\theta)$ to align with human feedback, where the policy $\pi_{\theta}$ is an optimal solution w.r.t.
+the certain reward estimate $r(\cdot,\cdot;\theta)$ according to the policy optimization problem (5b).
 Due to such optimal policy constraint w.r.t.
@@
 Policy Alignment Step.
-From our earlier discussion, we know that when $L_{3}$ takes the form (10), the optimal policy (for a fixed reward) is given by (5).
-Of course, one cannot directly compute such an optimal solution due to the fact that both $Q_{\theta}$ and the normalization term are unknown.
-Therefore one can adopt the standard approaches such as the well-known proximal policy optimization (PPO) (Schulman et al., 2017) algorithm to obtain an approximate optimal policy.
+One can adopt the standard approaches such as the well-known proximal policy optimization (PPO) (Schulman et al., 2017) algorithm to obtain an approximate optimal policy that solves (7).
 It is worth noting that, when considering $T=1$, our discussion leading to (42) indicates the optimal policy takes a much simpler form.
-In this case, it is possible to consider simpler method than running PPO to obtain the optimal policy.
-One alternative way is to use a baseline estimated reward value to perform variance reduction Li et al.
-(2023), thus reducing the computational complexity.
+In this case, it is possible to consider a simpler method than running PPO to obtain the optimal policy.
+One alternative way is to use a baseline estimated reward value to perform variance reduction (Li et al., 2023), thus reducing the computational complexity.
 It is important to note that, the point of the above discussion is that these different choices for solving the policy alignment problem can be incorporated into our overall approach.
-Reward Alignment. In this step, we use a stochastic gradient-type algorithm to optimize $L(\theta)$. Towards this end, first observe that
-$\nabla L(\theta)$ $=\alpha\nabla L_{1}(\pi_{\theta})+\nabla L_{2}(\theta).$ (26)
+Reward Alignment. In this step, we use a stochastic gradient-type algorithm to optimize $L(\theta)$. Towards this end, first, observe that
+$\nabla L(\theta)$ $=w_{1}\nabla L_{1}(\pi_{\theta})+\nabla L_{2}(\theta).$ (16)
 Clearly, regardless of the choice of $L_{2}$, $\nabla L_{2}$ is relatively easy to compute because the objective is directly related to $\theta$ since $L_{2}(\theta)$ can be regarded as one supervised learning loss and do not involve the optimal policy $\pi_{\theta}$.
 In particular, we have the following expressions:
-$\nabla{L}^{\rm BTL}_{2}(\theta)$ $=\mathbb{E}_{(\tau_{l}\prec\tau_{w})\sim\pi^{P}}\Big{[}\nabla_{\theta}\log\Big{(}\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}\Big{)}\Big{]}.$ (27)
+$\nabla{L}^{\rm BTL}_{2}(\theta)$ $=\mathbb{E}_{(\tau_{w}\succ\tau_{l})\sim\pi^{P}}\Big{[}\nabla_{\theta}\log\Big{(}\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}\Big{)}\Big{]}.$ (17a)
 On the contrary, the computation of $\nabla L_{1}(\pi_{\theta})$ is more involved, since $L_{1}$ depends on $\theta$ implicitly through the corresponding optimal policy $\pi_{\theta}$.
-Fortunately, the following lemma indicates that this gradient has a simple and intuitive form as well.
+Fortunately, the following lemma indicates that this gradient has a simple and intuitive form as well, and the proof can be found in Appendix A.3.2.
 Lemma 4.1
@@
 Instead, we propose to sample from the current policy $\pi^{k+1}$ obtained from the previous policy optimization step, where index $k$ represents the iteration counter.
+Following the approximation steps mentioned above, we construct a stochastic estimator $g_{k}$ to approximate the exact gradient $\nabla L(\theta_{k})$ in (16) as follows:
+$g_{k}:=w_{1}g^{k}_{1}+g^{k}_{2}:=$ $w_{1}\big{(}\nabla_{\theta}R(\tau^{E}_{k},\theta_{k})-\nabla_{\theta}R(\tau^{A}_{k},\theta_{k})\big{)}+\big{(}1-\sigma(R(\tau^{W}_{k},\theta_{k})-R(\tau^{L}_{k},\theta_{k}))\big{)}$
+×$\quad{\times}\big{(}\nabla_{\theta}R(\tau^{W}_{k},\theta_{k})-\nabla_{\theta}R(\tau^{L}_{k},\theta_{k})\big{)}.$ (19)
 The above two steps are summarized in Algorithm 1.
-Let us remark on the gradient expression of the proposed algorithm.
-Here, we explain how to construct the gradient estimator $g^{k}_{1}$ for $\nabla_{\theta}L_{1}(\theta_{k})$ defined in (28) and $g^{k}_{2}$ for $\nabla_{\theta}L_{2}(\theta_{k})$ defined in (27).
-At iteration $k$, to estimate $\nabla_{\theta}L_{1}(\theta_{k})$, we need to sample $N$ expert trajectories $\tau^{\rm E}$ from the expert demonstration dataset $\mathcal{D}^{\rm E}$ and $N$ agent trajectories $\tau^{\prime}$ from the policy $\pi^{k+1}$.
-Then the gradient estimator $g_{1}^{k}$ can be constructed as: $g_{1}^{k}:=\frac{1}{N}\sum_{i=1}^{N}\Big{[}\nabla_{\theta}R(\tau_{i};\theta)-\nabla_{\theta}R(\tau_{i}^{\prime};\theta)\Big{]}$.
-Furthermore, through sampling $M$ pairs of preference $(\tau_{l}\prec\tau_{w})$, we can further construct the gradient estimator of $\nabla_{\theta}L_{2}(\theta_{k})$ as: $g_{2}^{k}=\sum_{i=1}^{M}\nabla_{\theta}\log\Big{(}\sigma\big{(}R(\tau_{i,w};\theta)-R(\tau_{i,l};\theta)\big{)}\Big{)}$.
-Here, as a remark, although the demonstration data is used, but in the final algorithm, the gradient is always in constrastive form.
-The difference is that the for the demonstration data, it is compared with its own model generation, while for preference data, two preference data is compared.
-Before we close this section, let us remark on the computational complexity of the proposed algorithm.
-Note that our algorithm is motivated by a class of popular algorithms in bi-level optimization, where the upper-level and lower-level problems are updated alternatingly using stochastic optimization (Hong et al., 2020).
-Specifically, one only needs to perform one (or a few steps) of policy optimization step (such as PPO) to obtain an improved $\pi^{k+1}$ before moving to the next iteration.
-As demonstrated in our subsequent numerical results, the computational complexity of our method is comparable or sometimes even lower as compared to their sequential optimization counterparts such as RLHF.
-In our numerical results, we show that the proposed algorithm AIHF can achieve a comparable performance as RLHF while performing less policy optimization steps.
+let us remark on the computational complexity of the proposed algorithm.
+Note that our algorithm is motivated by a class of popular algorithms in bi-level optimization, where the upper-level and lower-level problems are updated alternatingly using stochastic optimization (Hong et al., 2023).
+We conclude the section by theoretically inspecting the proposed algorithms.
   Input: Initialize reward parameter $\theta^{0}$ and policy model $\pi^{0}$, the stepsize of reward update $\eta$.
@@
 for  Iteration $k=0,1,\ldots,K-1$ do      Policy Alignment: Optimizing $L_{3}$ by RL subroutine, e.g.
-PPO, to obtain one improved policy $\pi^{k+1}$      Data Sample I: Sample expert trajectories $\tau\sim\mathcal{D^{\rm E}}$ and agent trajectories from $\tau^{\prime}\sim\pi^{k+1}$      Data Sample II: Sample preference pair $(\tau_{l}\prec\tau_{w})\sim\mathcal{P}$      Estimating Gradient: Calculate the estimator $g^{k}:=\alpha g_{1}^{k}+g_{2}^{k}$ of $\nabla_{\theta}L(\theta_{k})=\alpha\nabla_{\theta}L_{1}(\theta_{k})+\nabla_{\theta}L_{2}(\theta_{k})$      Reward Alignment: $\theta^{k+1}:=\theta^{k}+\eta g^{k}$   end for
+PPO, to obtain one improved policy $\pi^{k+1}$      Data Sample I: Sample an expert trajectory $\tau\sim\mathcal{D^{\rm E}}$ and agent trajectory from $\tau^{\prime}\sim\pi^{k+1}$      Data Sample II: Sample preference pair $(\tau_{w}\succ\tau_{l})\sim\mathcal{P}$      Estimating Gradient: Calculate one gradient estimator $g^{k}:=w_{1}g_{1}^{k}+g_{2}^{k}$ of $\nabla_{\theta}L(\theta)=w_{1}\nabla_{\theta}L_{1}(\theta)+\nabla_{\theta}L_{2}(\theta)$      Reward Alignment: $\theta^{k+1}:=\theta^{k}+\eta g^{k}$   end for
 Algorithm 1 Alignment with Integrated Human Feedback (AIHF)
+Theorem 4.1
+Suppose Assumptions 1 - 2 hold.
+Selecting stepsize $\alpha:=\frac{\alpha_{0}}{K^{\sigma}}$ for the reward update step (4) where $\alpha_{0}>0$ and $\sigma\in(0,1)$ are some fixed constants, and $K$ is the total number of iterations to be run by the algorithm.
+Then the following result holds:
+$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\big{\|}\log\pi_{k+1}-\log\pi_{\theta_{k}}\big{\|}_{\infty}\right]=\mathcal{O}(K^{-1})+\mathcal{O}(K^{-\sigma})$ (20a)
+$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\|\nabla L(\theta_{k})\|^{2}\right]=\mathcal{O}(K^{-\sigma})+\mathcal{O}(K^{-1+\sigma})+\mathcal{O}(K^{-1})$ (20b)
+where $\|\log\pi_{k+1}-\log\pi_{\theta_{k}}\|_{\infty}:=\max_{s\in\mathcal{S},a\in\mathcal{A}}\big{|}\log\pi_{k+1}(a|s)-\log\pi_{\theta_{k}}(a|s)\big{|}$.
+In particular, setting $\sigma=1/2$, then both quantities in (20a) and (20b) converge with the rate $\mathcal{O}(K^{-1/2}).$
+The above theorem shows that Alg. 1 could converge to the stationary point if we take a large loop number $K$. Note that details and proofs of the result above are delegated to Appendix A.3.
 5 Experiments
-In this section, we present numerical results for the proposed method.
-More specifically, we intend to address the following questions: 1) How are the data efficiency and model improvement efficiency (w.r.t.
-the KL divergence violation) of our proposed method compared with other baselines?
-2) Can we achieve the same convergence results after integrating the three stages into one stage?
-3) And finally, how much does demonstration help in reward learning in AIHF?
-To answer these questions, we consider several applications of the alignment problem.
-In this section, we show the numerical results for large language model alignment in the Helpful and Harmless (HH) dataset (Bai et al., 2022) and robotics control problem in MuJoCo Todorov et al.
-(2012).
-In the appendix, we have another set of experiment where the goal is to align the language model so that it generates movie reviews with positive sentiment on IMDb dataset (Maas et al., 2011).
-6.1 MuJoCo Tasks
+In this section, we provide numerical evaluations of the proposed method (5) (Alg.
+1) and its variants (10) and (3.3), and comparing them with state-of-the-art methods RLHF (Ouyang et al., 2022), DPO (Rafailov et al., 2023), IPO (Calandriello et al., 2024) and SPIN (Chen et al., 2024).
+Our experiments demonstrate the advantages of the proposed methods in the following aspects: (1) Reward learning from demonstration and preference is the key to improving over standard RLHF.
+(2) Using demonstration in reward learning could increase model improvement efficiency (w.r.t.
+the KL divergence violation) (3) AIHF could reduce the effect of distribution mismatch caused by the sequential alignment method, thus breaking the performance limits of the state-of-the-art methods.
+Models and datasets.
+In the first setting, we test Alg.
+1 on Anthropic-HH (Bai et al., 2022) dataset [^2] Dataset available at https://huggingface.co/datasets/Anthropic/hh-rlhf.
+with (relatively small) Pythia (Biderman et al., 2023) models [^3] Models available at https://huggingface.co/EleutherAI.
+as policy models.
+Anthropic-HH is a preference dataset collected from 52B LLMs that provide two continuations based on helpfulness and harmlessness, and we pick 10k chosen/preferred continuation data to form the demonstration dataset, while others serve as preference dataset and RL prompt dataset.
+For the HH dataset, We first fine-tune the language models (Pythia-160M/1B/2.8B) through supervised fine-tuning over all chosen responses from the HH dataset for 1 epoch, we call it full-SFT model and use it as our base model.
+Moreover, we also SFT the language model using the selected top 10k chosen responses and name it as demonstration-SFT model.
+For each policy model, we use the exact same model Pythia-1.4B as the reward model.
+The other setting we test is on 7B models.
+We use Ultrafeedback [^4] Available at https://huggingface.co/datasets/HuggingFaceH4/ultrafeedback_binarized.
+as our preference dataset (61.1k preference data) and Ultrachat200k [^5] Available at https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k.
+as the demonstration dataset (208k demonstration data), with mistral-7b-sft-beta  [^6] Available at https://huggingface.co/HuggingFaceH4/mistral-7b-sft-beta.
+(Jiang et al., 2023) as our base model.
+We use the same mistral-7b-sft-beta model as the initialization of the reward model.
+Evaluation.
+For the Anthropic-HH dataset, we present the reward evaluated by the PKU-Alignment/beaver-7b-v3.0-reward model(Ji et al., 2024).
+In our 7B model experiments, we adopt the widely recognized HuggingFace Open LLM Leaderboard framework (Beeching et al., 2023).
+This evaluation suite measures LLM performance across six tasks: commonsense reasoning (Arc (Clark et al., 2018), HellaSwag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021)), multi-task language understanding (MMLU (Hendrycks et al., 2020)), mimicking human falsehoods (TruthfulQA (Lin et al., 2021)), and math problem-solving (GSM8K (Cobbe et al., 2021)).
+Additional implementation details can be found in Appendix A.1.
+(a) Pythia-160M
+(b) Pythia-1B
+(c) Pythia-2.8B
+Figure 2: Experiment results of Pythia-160M/1B/2.8B policy models, with the reward model trained from Pythia-1.4B.
+We record the average scores (across three trials) of AIHF and RLHF on the Anthropic-HH test dataset (See Tab.
+2 in Appendix for more comparisons with other algorithms)
+Results of small model (160M, 1B, and 2.8B) experiments.
+We observe that the proposed AIHF performs effectively when initiated from both the demonstration-SFT model and the full-SFT model.
+As shown in Fig.2, utilizing the same data, AIHF  algorithm can eventually outperform RLHF irrespective of the initial model.
+Furthermore, according to the numerical results as shown in Fig.3(a), we see that the proposed AIHF algorithm has a smaller deviation from the base model compared with the RLHF algorithm.
+This benefit of the AIHF approach is due to the fact that we incorporate the maximum likelihood IRL objective for both reward learning and policy learning.
+In this case, both the reward model and policy model will be trained to align with the demonstrations, which are also used in the training process of the SFT stage.
+We also conducted a study on the demonstration/preference data ratio in Fig.
+3(b) and 3(c).
+We observe that AIHF consistently outperforms RLHF across different demonstration/preference data ratios.
+Furthermore, we also record the performance of the two variants, namely Direct AIHF (10) and Self-Play-AIHF (3.3), also RLHF, DPO, IPO, and SPIN in Tab.
+2.
+Our proposed AIHF still outperforms all these methods in this experiment setting.
+(a) KL divergence to demonstration-SFT policy
+(b) AIHF vs RLHF with 10k demonstration, 5k preference
+(c) AIHF vs RLHF with 10k demonstration, 10k preference
+Figure 3: Experiment results on Pythia-1B policy models, where the reward model is trained from Pythia-1.4B models.
+We record the average scores of AIHF and RLHF on the Anthropic-HH test dataset, reporting the results across three different trials.
+Results of large model (7B) experiments.
+We run AIHF, Direct AIHF, and Self-Play AIHF along with other methods on the 7B experiment setting.
+The results are presented in Fig.
+4 (the numbers are recorded in Tab.
+4), where we can see that similar to the 1B setting, AIHF is consistently outperforming other methods.
+Additionally, both Direct AIHF and Self-Play AIHF effectively outperform the RLHF model (zephyr-7b-beta).
+The success of AIHF, as well as Self-Play AIHF and Direct AIHF further suggests that joint learning from demonstration and preference is indeed beneficial for the alignment.
+We also conduct an ablation study with different choices of $w_{1}$ in (10), as shown in Tab.
+6, the improvement of joint learning methods over baseline is robust.
+Furthermore, We also evaluate the reward models estimated using different methods (DPO, standard preference learning, and AIHF) over the widely used RewardBench (Lambert et al., 2024).
+The results, illustrated in Tab.
+5 in the Appendix, show that the reward model trained through the AIHF can achieve significant improvement (especially on reasoning tasks) compared to both the standard BTL reward model in (3) and the implicit reward model in DPO.
+Figure 4: Performance comparison between Direct AIHF, Self-Play AIHF training across the six benchmark datasets (See also Table 4 in the Appendix).
+Other Results.
+Due to the page limits, we leave two additional experiments in the appendix: 1) movie review generation with positive sentiment on IMDb dataset (Maas et al., 2011), 2) experiment on Robotics control tasks in MuJoCo (Todorov et al., 2012).
+For the result of MuJoCo Experiment A.1.1, we observe that even though Behavior Cloning (BC)/SFT could provide a high-performing initialization, RLHF still fails to improve policy quality in the following RL stage.
+In the contract, AIHF can effectively integrate preferences and demonstrations, leading to a more robust reward function and consequently, a high-quality policy.
+For the IMDB result (Fig.
+6), We show that AIHF is able to alleviate the distribution mismatch between the generated trajectories by the policy, and the data that the learned reward model is able to rank.
+6 Conclusion
+In this work, we study the alignment problem when diverse data sources from human feedback are available.
+Furthermore, we have developed an algorithmic framework that can integrate both expert demonstration and pairwise comparison data from human feedback to learn the reward functions for further guiding policy learning/model fine-tuning in the alignment pipeline.
+Through extensive evaluations on robotic control tasks and large language model alignment tasks, we demonstrate that our proposed method can outperform existing benchmarks on alignment tasks and is able to recover a better reward model to guide policy learning.
+Appendix A Appendix
+A.1 Experiment Setup and Additional Result
+A.1.1 MuJoCo Tasks
 In MuJoCo, we consider several robotic control tasks with continuous action space.
@@
 Then we combine the low- and medium-quality data as the preference dataset and use high-quality as demonstration data.
-Figure 3: Left: Hopper Environment Middle: HalfCheetah Environment Right: Walker2d Environment; AIHF  (orange) vs RLHF(blue); results are averaged over 3 independent runs.
+Figure 5: Top-Left: Hopper Environment; Top-Right: HalfCheetah Environment; Bottom: Walker2d Environment; AIHF  (orange) vs RLHF (blue) vs IPL (purple) (Hejna & Sadigh, 2024); results are averaged over 3 independent runs.
 We use 10k demonstrations and 20k preferences.
-The RLHF curve is initialized from a policy pre-trained by BC; the AIHF from a random policy.
+The RLHF and IPL curve is initialized from a policy pre-trained by BC; the AIHF from a random policy.
 The performance is compared against the # of SAC steps performed (for AIHF each policy alignment performs 5k steps of SAC.)
 Results.
-We show that AIHF is able to integrate (insufficient amount of) demonstration data and (not-so-high-quality) preference data to generate high-quality policy, and it significantly outperforms the RLHF.
-In Fig.
-3, we observe that due to the limited number of demonstration data Ross and Bagnell (2010); Zeng et al.
-(2022b), BC fails to provide a high-performing initialization for finetuning for RLHF.
+We show that AIHF is able to integrate (insufficient amount of) demonstration data and (not-so-high-quality) preference data to generate high-quality policy, and it significantly outperforms the RLHF and IPL.
+Note that IPL (Hejna & Sadigh, 2024) is actually the DPO-type algorithm applied to multi-horizon MDP, therefore it is an ideal baseline for the MuJoCo setting (since the underlying problem is a multi-horizon MDP).
+In Figure 5, We can see that AIHF outperforms both RLHF and IPL algorithms.
+We also observe that due to the limited number of demonstration data, even BC could provide a high-performing initialization, RLHF still fails to improve policy quality in the following RL stage Ross & Bagnell (2010); Zeng et al.
+(2022b).
 Moreover, since the preference data quality is only of low-to-medium quality, the RL step based on the learned reward model fails to significantly boost the fine-tuning performance.
 In contrast, clearly the proposed AIHF can effectively integrate preferences and demonstrations, leading to a more robust reward function and consequently, a high-quality policy.
-6.2 LLM Task
-In the experiment utilizing the HH dataset, we aim to generate helpful continuations through aligning large language models with human preference.
-We compare our proposed algorithm AIHF which updates reward and policy iteratively, with the standard RLHF benchmark.
-Additionally, $L_{2}(\cdot)$ is chosen to be $L_{2}^{\rm BTL}$ defined (11), and $L_{1}$, $L_{3}$ are chosen similarly as in MuJoCo experiment.
-More settings are listed in the Appendix A.2.
-For evaluation, we use the existing reward model PKU-Alignment/beaver-7b-v3.0-reward [^2] https://huggingface.co/PKU-Alignment/beaver-7b-v3.0-reward for the helpful continuation generation; use the existing sentiment classifier lvwerra/distilbert-imdb [^3] https://huggingface.co/lvwerra/distilbert-imdb as the ground truth reward model for the movie review generation task.
-Data.
-We divide the training dataset of HH dataset into three different subsets: demonstration dataset, preference dataset, and RL training dataset, We employed the reward model PKU-Alignment/beaver-7b-v3.0-reward Dai et al.
-(2023); Ji et al.
-(2024) to filter the datasets.
-Specifically, We selected the 10k responses with the highest reward score as our demonstration dataset and utilized an additional 10,000 pairwise comparisons as preference data.
-The remaining prompts were then used to run the RL algorithm.
-Setting.
-For the HH dataset, We first fine-tune the language models (Pythia-160M / Pythia-1B) Biderman et al.
-(2023) through supervised fine-tuning over all chosen responses from the HH dataset for 1 epoch, we call it full-SFT model and use it as our base model.
-Moreover, we also SFT the language model using the selected top 10k chosen responses and name it as demonstration-SFT model.
-For the reward model, we use EleutherAI/pythia-1.4B  [^4] https://huggingface.co/EleutherAI/pythia-1.4b for both 160m and 1B policy as its base model.
-Figure 4: Helpfulness-controlled Generation.
-Left: Results on 160M models, Right: Results on 1B models.
-We record the average scores of AIHF and RLHF on the Anthropic-HH test dataset, reporting the results across three different trials.
-Results.
-We observe that the proposed approach AIHF performs effectively when initiated from both the demonstration-SFT model and the full-SFT model.
-As shown in Fig.
-4, utilizing the same data, the AIHF  algorithms can eventually outperform RLHF irrespective of the initial model.
-Furthermore, according to the numerical results as shown in Fig.
-5, compared with the RLHF benchmark, we see that the proposed AIHF algorithm has smaller deviation from the base model.
-This benefit of the AIHF approach is due to the fact that we incorporate the maximum likelihood IRL objective for both reward learning and policy learning.
-In this case, both reward model and policy model will be trained to align with the demonstrations, which are also used in the training process of the SFT stage.
+In SAC, both the policy network and Q network are (64, 64) MLPs with ReLU activation function, and the step size is set to $3*10^{-3}$, we parameterize the reward function by a (64, 64) MLPs with ReLU activation function.
+For the reward network, we use Adam as the optimizer, and the step size is set to be $1*10^{-4}$.
+The quality of the preference dataset and demonstration dataset are listed in Tab. 1.
+Task Dataset Non-prefer Data Prefer Data Demonstration Data
+Hopper-v2 $2345.20\pm 329.93$ $3024.63\pm 40.52$ $3559.61\pm 73.12$
+HalfCheetah-v2 $7226.37\pm 126.88$ $9434.42\pm 1315.13$ $11635.42\pm 236.51$
+Walker2d-v2 $3952.60\pm 444.45$ $5091.71\pm 291.73$ $5453.41\pm 71.07$
+Table 1: The quality of preference and demonstration.
+A.1.2 Sentiment-Controlled Generation
+Dataset Generation: In the IMDb sentiment completion task, we generate the demonstrations and preference datasets using the following procedure.
+Initially, we train a Language Model by employing the ground-truth reward function distilbert-imdb and the Proximal Policy Optimization (PPO) algorithm on 30% of the training dataset for IMDb.
+Throughout the training process, we save the policy checkpoint every 500 PPO steps.
+Subsequently, we select an additional 40% of the training dataset and generate a response for each prompt for each checkpoint.
+According to the evaluation score of each generation, we categorize the data collected into different classes: low-, medium-, and high-quality datasets, then we combine low-quality and medium-quality as preference datasets, and use high-quality as demonstration datasets.
+Training: After acquiring the preference and demonstration datasets, we train the proposed algorithm AIHF and baselines on the remaining 30% of prompts from the training dataset.
+We evaluate the performance of each algorithm using the test datasets for IMDb and HH, along with their corresponding ground truth reward functions.
+For the GPU resources, we use 8$\times$ A100 40G for all the experiments.
+Results: Policy Quality.
+We find that the proposed approach works well when either preference or demonstration data, or both, are limited.
+From the 7, we see that by using the same amount of data (10k preference, 10k demonstration), AIHF-based algorithms achieve faster convergence than their RLHF and DPO counterparts.
+Results: Distribution Mismatch.
+We show that AIHF  is able to alleviate the distribution mismatch between the generated trajectories by the policy, and the data that the learned reward model is able to rank.
+To evaluate the extend of such mismatch, we use the following three steps: (1) use 1k preference, 1k demonstration to train policy and reward model for RLHF and AIHF ; (2) for a given set of prompts from test dataset, use RLHF and AIHF  to perform generation; (3) use the trained reward models to rate the generation; (4) compare with the score generated by the ground truth reward lvwerra/distilbert-imdb.
 Fig.
-5 shows AIHF  effectively maintains the policy close to the initial model.
-This advantage can be partially attributed to Eq.
-(8a), as the inclusion of the demonstration regularization term encourages the alignment with the demonstration distribution, thus bringing better performance without significant deviation from the initial model.
-The win rate (ratio of samples where the reward of our model’s generation is higher than the model compared) provided by Fig 5 further shows the superiority of the proposed approach.
-Figure 5: Helpfulness-controlled Generation on 1B models. Left: KL divergence to the Demonstration-SFT policy, Right: win rate of AIHF and RLHF over the Demonstration-SFT model.
-7 Conclusion
-In this work, we study the alignment problem when diverse data sources from human feedback are available.
-Furthermore, we have developed an algorithmic framework that can integrate both expert demonstration and pairwise comparison data from human feedback to learn the reward functions for further guiding policy learning/model fine-tuning in the alignment pipeline.
-Through extensive evaluations on robotic control tasks and large language model alignment tasks, we demonstrate that our proposed method can outperform existing benchmarks on alignment tasks and is able to recover a better reward model to guide policy learning.
-One limitation is that our framework only incorporates expert demonstrations and pairwise comparisons for reward learning.
-In fact, it is promising to extend the information sources from human feedback/behaviors to further strengthen the reward learning/policy fine-tuning stages.
-For example, one can utilize precise comparisons and demonstrations from different aspects to further learn a powerful reward model in order to support better alignment quality.
-Appendix A Appendix / supplemental material
-A.1 Proof of Lemma 5.1
-Proof. Here, under a reward parameter $\theta$ and the corresponding optimal policy $\pi_{\theta}$ of (5).
+6 illustrates that the reward score distribution produced by AIHF  aligns closely with that of the ground truth reward, whereas that generated by RLHF exhibits a poor match.
+These results show that the reward model learned by AIHF is able to correctly evaluate the generation produced by the final policy.
+Figure 6: Comparison of the distribution of reward score generated by the trained reward models, and the ground truth reward model. RLHF vs ground truth (left); AIHF vs ground truth (right).
+From Fig.7, our proposed algorithm AIHF  could obtain higher rewards than baseline methods in the IMDb setting for almost all KL values.
+Although AIHF  might get a low score from the ground truth reward model in the earlier step, AIHF  would get a higher reward with more iteration and optimization steps.
+This indicates that with the mix of demonstration data and preference data, we could prevent the policy from known issues of reward hacking, especially when the policy learned more human-aligned features beyond base models (high KL value).
+Moreover, AIHF  is persistent in the number of preference data, presenting that AIHF  could still gain benefit from the limited preference data in more optimization steps as long as the demonstration data is high quality enough.
+Figure 7: The frontier of expected reward vs KL to the reference policy in IMDB dataset.
+fix the demonstration number to 3k Left: Using 1k preference; Middle: Using 2k preferences; Right: Using 3k preference
+A.1.3 The result of 1B and 2.8B experiments
+Additional Results.
+We show here the distribution of the reward values of the continuation generated by the reward models trained by AIHF and RLHF.
+In Fig.
+8, we can observe the distribution of AIHF and RLHF have overlaps in low-quality continuation, however, AIHF can generate more high-quality continuations compared to RLHF, which shows that joint optimization can more effectively align the policy model with the demonstration distribution.
+Figure 8: The Reward Distribution of Helpfulness-controlled Generation.
+Left: Result on 160m model, Right: Results on 1B model, This figure reports the reward distribution of generation evaluated by PKU-Alignment/beaver-7b-v3.0-reward for AIHF and RLHF.
+We record the performance of the two variants, namely Direct AIHF (10) and Self-Play-AIHF (3.3), also RLHF, DPO, IPO and SPIN in Tab.
+2.
+Our proposed AIHF still outperforms all these methods in this experiment setting.
+Dataset DPO SPIN IPO RLHF Direct AIHF Self-Play-AIHF AIHF
+10k Demonstrations, 5k preferences $0.463\pm 0.093$ $0.625\pm 0.048$ $0.616\pm 0.076$ $0.710\pm 0.085$ $0.752\pm 0.036$ $0.640\pm 0.102$ $1.167\pm 0.157$
+10k Demonstrations, 10k preferences $0.474\pm 0.052$ $0.625\pm 0.048$ $0.650\pm 0.017$ $0.896\pm 0.095$ $0.798\pm 0.026$ $0.674\pm 0.096$ $1.190\pm 0.069$
+Table 2: Evaluated reward scores of fine-tuned 1B Pythia models in HH dataset, evaluated by PKU-Alignment/beaver-7b-v3.0-reward.
+A.1.4 The result of 7B experiments
+We utilize DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) to reduce the memory cost, and we use VLLM (Kwon et al., 2023) for accelerate data generation/inference.
+We use eight NVIDIA A100-40G to do the training with a per-device batch size of 1 for 7B model.
+We train all models with bfloat16 precision.
+We set the learning rate to be 5e-7 for the 7B model with the cosine learning rate scheduler, and the max sequence length is set to be 512.
+Our device cannot conduct the implementation of the standard AIHF (5) using PPO.
+Therefore, we employed an online DPO approach (Xiong et al., 2024) for the lower-level policy optimization: 1) at each policy optimization iteration, we first generate multiple continuations from the current policy model over the given prompt dataset, 2) then we utilize our estimated explicit reward models to score each continuation, 3) over each prompt, choose the generated continuation with the highest reward score and the one with lowest reward score as a generated preference pair to run DPO algorithm to fine-tune the policy model.
+For Self-Play AIHF (3.3) implementation, we adopt the same strategy as Chen et al.
+(2024), at each epoch, we generate samples with picked 50k data and generate continuation $\tilde{a}\sim\pi(\cdot|s)$ using the current model $\pi$, then optimize (3.3) with the sampled $\tilde{a}$.
+To evaluate the effectiveness of our approach, we assessed our fine-tuned models on the widely used HuggingFace Open LLM Leaderboard (Beeching et al., 2023) in Table 3.
+We also list the metric and number of shots used for LLM evaluation on each dataset in Table 4.
+Moreover, in Tab.
+5, we also utilize the Reward-Bench (Lambert et al., 2024) to evaluate the explicit reward models (estimated by standard preference learning in (3) and AIHF) and the implicit reward model (extracted from DPO).
+The implicit reward is calculated by the policy model similar to what the DPO paper Rafailov et al.
+(2023) proposed, i.e.
+$r(a|s):=\log(\frac{\pi(a|s)}{\pi^{0}(a|s)})$, the first explicit reward is estimated by BTL reward model (3), and second explicit reward is estimated by AIHF.
+The evaluation results show that AIHF can outperform the benchmark methods in terms of performance for both policy models and reward models.
+We also conduct an ablation study with different choices of $w_{1}$ in (10), as shown in Tab.
+6, the improvement of joint learning methods over baseline is robust.
+Dataset Arc Challenge TruthfulQA MC2 Winogrande GSM8K HellaSwag MMLU
+Metric acc acc acc strict-match acc_norm acc
+Num. of Shots 25 0 5 5 10 5
+Table 3: A summarization of the benchmarks we use in this work. We list the metric and number of shots used for LLM evaluation on each dataset.
+Tasks Arc Challenge TruthfulQA MC2 Winogrande GSM8k HellaSwag MMLU Avg
+mistral-7b-sft-beta 54.69% 42.96% 77.27% 39.88% 82.23% 59.72% 59.46%
+zephyr-7b-beta 59.64% 55.18% 77.82% 33.51% 84.19% 59.76% 61.68%
+SPIN 58.45% 43.66% 78.30% 39.50% 83.59% 58.60% 60.35%
+DPO 62.80% 53.17% 79.40% 39.20% 85.13% 59.41% 63.19%
+IPO 58.02% 48.29% 79.24% 42.91% 83.93% 60.07% 62.08%
+Direct AIHF 61.17% 60.03% 79.00% 39.80% 85.71% 60.02% 64.29%
+Self-play AIHF 61.77% 58.29% 78.53% 44.20% 85.53% 58.66% 64.50%
+AIHF 63.90% 58.38% 79.24% 40.56% 86.23% 60.18% 64.75%
+Table 4: Test performance of Direct AIHF and Self-Play AIHF based on mistral-7b-sft-beta across HuggingFace Open LLM Leaderboard datasets.
+Reward Model Chat Chat Hard Safety Reasoning Average
+DPO Reward Model 37.43% 55.92% 64.14% 47.33% 51.21%
+BTL Reward Model 95.11% 56.58% 63.69% 69.22% 71.15%
+AIHF Reward Model 94.41% 55.37% 63.98% 76.75% 72.63%
+Table 5: Evaluation of Reward Models in Reward-Bench.
+Tasks Arc Challenge TruthfulQA MC2 Winogrande GSM8k HellaSwag MMLU Avg
+Direct AIHF($w_{1}=0.01$) 61.86% 57.55% 79.08% 36.61% 85.58% 60.09% 63.46%
+Direct AIHF($w_{1}=0.001$) 63.25% 58.73% 79.16% 36.84% 85.59% 59.26% 63.80%
+Direct AIHF($w_{1}=0.0001$) 61.17% 60.03% 79.00% 39.80% 85.71% 60.02% 64.28%
+Table 6: Test performance of Direct AIHF over different choices of the balancing coefficient across HuggingFace Open LLM Leaderboard datasets.
+A.2 Why AIHF Can Outperform Two-Stage Alignment Approaches
+In this section, we provide detailed analysis for Sec. 3.4.
+A.2.1 SFT and RLHF Policy
+We revisit the RLHF pipeline in the context of a simple softmax choice model where $\tau_{i}\in A$ is selected with probability
+$\pi^{*}_{i}(R)=\frac{\exp(R_{i}/\beta)}{\sum_{j=1}^{N}\exp(R_{j}/\beta)}.$ (21)
+For simplicity, we have defined $R_{i}:=R(\tau_{i})$. Using this policy model, below let us analyze different policies according to different ways that the rewards are learned.
+Supervised Fine-Tuning (SFT): Given a demonstration dataset $\mathcal{D}$ the goal is the find the policy $\pi_{\rm SFT}$ that maximizes likelihood, i.e.:
+$\pi_{\rm SFT}$ $:=\arg\max_{\pi}\sum_{\ell=1}^{N}\mathbb{E}_{\tau_{\ell}\sim\mathcal{D}}\left[\log\pi^{*}_{\ell}(R)\right],\quad\mbox{\rm s.t.}\;\sum_{\ell=1}^{N}\pi^{*}_{\ell}(R)=1.$ (22)
+Next let us identify the optimal reward and its corresponding policy in this case. First, let us find the reward function. Write (22) in terms of reward optimization, we have:
+$R^{*}=\arg\max_{R}\sum_{\ell=1}^{N}\mathbb{E}_{\tau_{\ell}\sim D}\log\frac{\exp(R_{\ell}/\beta)}{\sum_{j=1}^{N}\exp(R_{j}/\beta)}=:L_{1}(R)$ (23)
+The partial derivative of the objective function w.r.t. a component $R_{i},i\in[1,\cdots,N]$ is given by:
+$\frac{\partial L_{1}(R)}{\partial R_{i}}$ $=\sum_{\ell:\ell\neq i}\mathbb{E}_{\tau_{\ell}\sim D}\left[-\frac{1}{\beta}\frac{\exp(R_{i}/\beta)}{\sum_{j=1}^{N}\exp(R_{j}/\beta)}\right]+\mathbb{E}_{\tau_{i}\sim D}\left[\frac{1}{\beta}-\frac{1}{\beta}\frac{\exp(R_{i}/\beta)}{\sum_{j=1}^{N}\exp(R_{j}/\beta)}\right]$
+$=\frac{1}{\beta}\left(\frac{\#\{\tau_{i}\in\mathcal{D}\}}{|P|}-\sum_{\ell}\mathbb{E}_{\tau_{\ell}\in D}\left[\frac{\exp(R_{i}/\beta)}{\sum_{j=1}^{N}\exp(R_{j}/\beta)}\right]\right),\quad i\in[1:N].$ (24)
+Setting the above partial gradient to zero, we obtain
+#$\frac{\#\{\tau_{i}\in\mathcal{D}\}}{|\mathcal{D}|}=\frac{\exp(\widehat{R}_{\mathcal{D}}(\tau_{i})/\beta)}{\sum_{j=1}^{N}\exp(\widehat{R}_{\mathcal{D}}(\tau_{j})/\beta)}~{}~{}~{}~{}i\in[1:N],$ (25)
+where we used $\widehat{R}_{D}(\tau_{i})$’s to denote the optimal reward estimated from the set of SFT data $\mathcal{D}$. Then according to the definition of the policy given earlier, we obtain
+$\pi^{*}_{i}(\widehat{R}_{\mathcal{D}})=\frac{\#\{\tau_{i}\in\mathcal{D}\}}{|\mathcal{D}|},\;\forall~{}i\in[1,\cdots N].$
+Note that the optimal reward learned from the SFT data $\widehat{R}_{D}$ should satisfy the system (25). It turns out that there is a unique solution to the following system of equations:
+#$\frac{\#\{\tau_{i}\in\mathcal{D}\}}{|\mathcal{D}|}=\frac{\exp(\widehat{R}_{\mathcal{D}}(\tau_{i})/\beta)}{\sum_{j=1}^{N}\exp(\widehat{R}_{\mathcal{D}}(\tau_{j})/\beta)}~{}~{}~{}~{}i\in\{2,\dots,N\}$ (26)
+with $\widehat{R}_{\mathcal{D}}(\tau_{1})=\bar{R}_{1}$ a fixed reference value. A simple argument is provided below. The system (26) can be re-written as:
+$\log\frac{\#\{\tau_{i}\in\mathcal{D}\}}{|\mathcal{D}|}-\log\frac{\#\{\tau_{1}\in\mathcal{D}\}}{|\mathcal{D}|}$ $=\frac{1}{\beta}\big{(}\widehat{R}_{\mathcal{D}}(\tau_{i})-\bar{R}_{1}\big{)}~{}~{}~{}~{}~{}i\in\{2,\dots,N\}$
+Hence, rewards are uniquely determined by demonstration data with a fixed reference value $\widehat{R}_{\mathcal{D}}(\tau_{1})=\bar{R}_{1}$.
+It is important to note that, despite the fact that in the SFT we may not directly learn a reward function, the above analysis says that we are implicitly learning a reward $\widehat{R}_{D}$, assuming that the policy is parameterized as (21).
+This implicit reward will be used later to analyze the RLHF policy.
+Indeed, if we solve the following SFT policy optimization problem directly
+$\max\sum_{\ell}\mathbb{E}_{\tau_{\ell}\in\mathcal{D}}\log(\pi_{\ell}),\quad{\rm s.t.}\quad\sum_{i\in\mathcal{D}}\pi_{i}=1,$ (27)
+one can easily obtain that the optimal policy is given by $\pi^{*}_{i}=\frac{\#\{\tau_{i}\in\mathcal{D}\}}{|\mathcal{D}|},\;\forall~{}i\in[1,\cdots N]$.
+Reward Modeling with Preference Data: Now let us analyze the case where the reward is only learned from a preference dataset.
+With preference data $\mathcal{P}:=\{(\tau_{i}\succ\tau_{j})\}$, the BTL model is:
+$P(\tau_{i}\succ\tau_{j})$ $=\sigma\big{(}\frac{1}{\beta}\big{(}R(\tau_{i})-R(\tau_{j})\big{)}=\frac{\pi^{*}_{i}(R)}{\pi^{*}_{i}(R)+\pi^{*}_{j}(R)}.$
+The reward estimation problem is defined as:
+$\widehat{R}_{\mathcal{P}}=\arg\max_{R}L_{2}(R)$ $:=\sum_{\ell,j:\ell\neq j}\mathbb{E}_{(\tau_{\ell}\succ\tau_{j})\sim\mathcal{P}}\Big{[}\log\frac{\pi^{*}_{\ell}(R)}{\pi^{*}_{\ell}(R)+\pi^{*}_{j}(R)}\Big{]}$ (28)
+$=\sum_{\ell,j:\ell\neq j}\mathbb{E}_{(\tau_{\ell}\succ\tau_{j})\sim\mathcal{P}}\left[\log\left(\frac{\exp(R_{\ell}/\beta)}{\exp(R_{\ell}/\beta)+\exp(R_{j}/\beta)}\right)\right]$ (29)
+Take the gradient w.r.t. $R_{i}$, we obtain
+$\frac{\partial\ell_{RM}({R})}{\partial R_{i}}$ $=\sum_{j:j\neq i}\mathbb{E}_{(\tau_{i}\succ\tau_{j})}\left(\frac{1}{\beta}-\frac{1}{\beta}\left(\frac{\exp(R_{\ell}/\beta)}{\exp(R_{\ell}/\beta)+\exp(R_{j}/\beta)}\right)\right)$
+$\quad+\sum_{j:j\neq i}\mathbb{E}_{(\tau_{j}\succ\tau_{i})}\left[-\frac{1}{\beta}\left(\frac{\exp(R_{\ell}/\beta)}{\exp(R_{\ell}/\beta)+\exp(R_{j}/\beta)}\right)\right]$
+$=\frac{1}{\beta}\sum_{j:j\neq i}\frac{|\mathcal{P}_{i\succ j}|}{|\mathcal{P}|}\left(1-\left(\frac{\exp(R_{\ell}/\beta)}{\exp(R_{\ell}/\beta)+\exp(R_{j}/\beta)}\right)\right)$
+$\quad+\frac{1}{\beta}\sum_{j:j\neq i}\frac{|\mathcal{P}_{j\succ i}|}{|\mathcal{P}|}\left(-\left(\frac{\exp(R_{\ell}/\beta)}{\exp(R_{\ell}/\beta)+\exp(R_{j}/\beta)}\right)\right)$
+$=\frac{1}{\beta}\left(\sum_{j:j\neq i}\frac{|\mathcal{P}_{i\succ j}|}{|\mathcal{P}|}-\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}\frac{\exp(R_{\ell}/\beta)}{\exp(R_{\ell}/\beta)+\exp(R_{j}/\beta)}\right)$
+$=\frac{1}{\beta}\sum_{j:j\neq i}\Big{(}\frac{|\mathcal{P}_{i\succ j}|}{|\mathcal{P}_{i,j}|}-\frac{\pi^{*}_{i}(R)}{\pi^{*}_{i}(R)+\pi^{*}_{j}(R)}\Big{)}\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}$ (30)
+where in the above derivation we have defined $|\mathcal{P}_{i\succ j}|:=\#\{\tau_{i}\succ\tau_{j}~{}\mbox{in}~{}\mathcal{P}\}$ and $|\mathcal{P}_{i,j}|:=|\mathcal{P}_{i\succ j}|+|\mathcal{P}_{j\succ i}|$.
+Setting the gradient to zero, we obtain
+$\sum_{j:j\neq i}\Big{(}\frac{|\mathcal{P}_{i\succ j}|}{|\mathcal{P}_{i,j}|}-\frac{\pi^{*}_{i}(\widehat{R}_{P})}{\pi^{*}_{i}(\widehat{R}_{P})+\pi^{*}_{j}(\widehat{R}_{P})}\Big{)}\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}=0.$ (31)
+We shall denote by $\widehat{R}_{\mathcal{P}}$ as a solution to the above system of equations. The first-order condition (31) can be written in implicit form as:
+$\pi^{*}_{i}(\widehat{R}_{P})=\frac{\sum_{j:j\neq i}|\mathcal{P}_{i\succ j}|}{\sum_{j:j\neq i}\frac{|\mathcal{P}_{ij}|}{\pi_{i}^{*}(\widehat{R}_{P})+\pi_{j}^{*}(\widehat{R}_{P})}}=\frac{\sum_{j:j\neq i}|\mathcal{P}_{i\succ j}|}{\sum_{j:j\neq i}\frac{|\mathcal{P}_{ij}|}{1-\sum_{\ell\in A\backslash\{i,j\}}\pi^{*}_{\ell}(\widehat{R}_{P})}}:=\frac{\sum_{j:j\neq i}|\mathcal{P}_{i\succ j}|}{\sum_{j:j\neq i}{|\mathcal{P}_{ij}|}\rho_{-(i,j)}(\pi^{*}(\widehat{R}_{P}))}$ (32)
+where we have defined $\rho_{-(i,j)}(\pi):=\Big{(}1-\sum_{\ell\in A\backslash\{i,j\}}\pi_{\ell}\Big{)}^{-1}$ as the expected number of times an action other than $\tau_{i}$ or $\tau_{j}$ is selected when sampling actions from $\pi$ infinitely many times.
+RLHF Policy. Based on the above two analysis, we are ready to analyze the RLHF policy. The RLHF policy is defined as follows:
+$\pi^{\rm RLHF}=\arg\max_{\pi\in\Delta^{N}}~{}~{}\mathbb{E}_{\tau_{i}\sim\pi}\big{[}{\widehat{R}_{\mathcal{P}}}(\tau_{i})\big{]}-\beta\mathrm{KL}(\pi||\pi^{\rm SFT})$
+where $\widehat{R}_{\mathcal{P}}$ is the estimator obtained from preference data, $\pi^{\rm SFT}$ is the SFT model trained with demonstration dataset $\mathcal{D}$, and $\Delta^{N}$ is the probability simplex.
+It can be easily shown that the solution $\pi^{\rm RLHF}$ is of the form:
+$\pi^{\rm RLHF}(\tau_{i})$ $=\frac{\pi^{\rm SFT}(\tau_{i})\exp\left(\frac{1}{\beta}\widehat{R}_{\mathcal{P}}(\tau_{i})\right)}{\sum_{j=1}^{N}\pi^{\rm SFT}(\tau_{j})\exp\left(\frac{1}{\beta}\widehat{R}_{\mathcal{P}}(\tau_{j})\right)}$
+[…]
+$={\frac{\exp\big{(}\frac{1}{\beta}(\widehat{R}_{\mathcal{D}}(\tau_{i})+\widehat{R}_{\mathcal{P}}(\tau_{i}))\big{)}}{\sum_{j=1}^{N}\exp\big{(}\frac{1}{\beta}(\widehat{R}_{\mathcal{D}}(\tau_{j})+\widehat{R}_{\mathcal{P}}(\tau_{j}))\big{)}}}$
+$=\pi^{*}_{i}\Big{(}\widehat{R}_{\mathcal{D}}+\widehat{R}_{\mathcal{P}}\Big{)}.$ (33)
+A.2.2 The AIHF Policy
+The AIHF Policy. The proposed AIHF estimation problem is
+$\widehat{R}^{\rm AIHF}=\arg\max_{R}\ell_{\mathcal{D}+\mathcal{P}}(R):=|\mathcal{D}|L_{1}(R)+|\mathcal{P}|L_{2}(R)$ (34)
+where $L_{1}(R):=\mathbb{E}_{\tau_{i}\sim\mathcal{D}}[\log\pi^{*}_{i}(R)]$ and $L_{2}(R):=\mathbb{E}_{(\tau_{j}\prec\tau_{i})\sim\mathcal{P}}[\log\frac{\pi^{*}_{i}(R)}{\pi^{*}_{i}(R)+\pi^{*}_{j}(R)}]$, and $\pi^{*}_{i}(R)$ is given by (21).
+Similarly as before, taking gradient w.r.t. $R_{i}$, and leverage (A.2.1) and (A.2.1), we obtain
+$\frac{\partial\ell_{\mathcal{D}+\mathcal{P}}(R)}{\partial R_{i}}$ =#$=\frac{\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}}{|\mathcal{D}|}|\mathcal{D}|-\pi^{*}_{i}(R)|\mathcal{D}|+\sum_{j:j\neq i}\left(|\mathcal{P}_{i\succ j}|-\sum_{j\neq i}\frac{\pi^{*}_{i}(R)}{\pi^{*}_{i}(R)+\pi^{*}_{j}(R)}|\mathcal{P}_{i,j}|\right)$ (35)
+Setting the above condition to zero, we obtain that the AIHF reward should satisfy the following
+#$\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}+\sum_{j:j\neq i}|\mathcal{P}_{i\succ j}|$ $=\pi^{*}_{i}(\widehat{R}^{\rm AIHF})\Big{(}|\mathcal{D}|+\sum_{j:j\neq i}\frac{|\mathcal{P}_{i,j}|}{\pi^{*}_{i}(\widehat{R}^{\rm AIHF})+\pi^{*}_{j}(\widehat{R}^{\rm AIHF})}\Big{)}$
+$=\pi^{*}_{i}(\widehat{R}^{\rm AIHF})\Big{(}|\mathcal{D}|+\sum_{j:j\neq i}|\mathcal{P}_{i,j}|\rho_{-(i,j)}\big{(}\pi^{*}(\widehat{R}^{\rm AIHF})\big{)}\Big{)}$
+Or equivalently,
+$\pi^{*}_{i}(\widehat{R}^{\rm AIHF})$ =#$=\frac{\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}+\sum_{j\neq i}|\mathcal{P}_{i\succ j}|}{|\mathcal{D}|+\sum_{j:j\neq i}|\mathcal{P}_{i,j}|\rho_{-(i,j)}\big{(}\pi^{*}(\widehat{R}^{\rm AIHF})\big{)}}$ (36)
+The system (36) has a unique solution $\widehat{R}^{\rm AIHF}$ with a fixed reference value $\widehat{R}^{\rm AIHF}(\tau_{1})=\bar{R}_{1}$.
+Discussion.
+Now let us compare the RLHF policy (33) and AIHF policy (36).
+First, observe that the RLHF policy is a function of the sum of two reward functions, one learned from the demonstration data and one learned from the preference data, while the AIHF policy takes a more complex form.
+To have a better understanding of the two policies, let us consider the extreme cases where the SFT and the preference datasets are unbalanced, where either $|\mathcal{D}|\gg|\mathcal{P}|$ or $|\mathcal{P}|\gg|\mathcal{D}|$.
+  • Case $|\mathcal{D}|\gg|\mathcal{P}|$.
+In this case, $\pi^{*}_{i}(\widehat{R}^{\rm AIHF})\approx\frac{\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}}{|\mathcal{D}|}=\pi^{*}_{i}(\widehat{R}_{\mathcal{D}})$, so the AIHF policy will ignore the preference dataset, while focusing on the SFT policy; however, for the RLHF policy, one needs to first separately performs the reward estimation, but in this case $\hat{R}_{P}$ can be very noisy due to lack of data; Then the RLHF policy will also be noisy since it will be influenced by the noisy reward function $\hat{R}_{P}$.
+  • Case $|\mathcal{P}|\gg|\mathcal{D}|$.
+In this case, let’s further assume that $\sum_{j:j\neq i}|\mathcal{P}_{i\succ j}|\gg\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\},\;\forall~{}i$, that is, for each trajectory $\tau_{i}$, the number of times it appears in the preference data is much larger than its appearance in the demonstration data.
+In this case, it is easy to see that $\pi^{*}_{i}(\widehat{R}^{\rm AIHF})\approx\pi^{*}_{i}(\widehat{R}_{\mathcal{P}})$, while the RLHF policy will still be noisy as in the first case.
+A.2.3 Numerical Examples
+Example 1: With $\beta=1$ and only two actions $\tau_{1}$ and $\tau_{2}$. Since $\rho_{-(1,2)}(\pi)=\rho_{-(2,1)}(\pi)=1$, it follows from equations (26), (32), (36) that:
+$\pi^{\rm AIHF}_{1}:=\pi^{*}_{1}(\widehat{R}^{\rm AIHF})$ =#$=\frac{\#\{\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}+\#\{\tau_{2}\prec\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}}{|\mathcal{D}|+|\mathcal{P}|}$
+$=\frac{|\mathcal{D}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{R}_{\mathcal{D}})+\frac{|\mathcal{P}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{R}_{\mathcal{P}}).$
+Slightly abusing notations, let $\pi^{*}_{1}:=\pi^{*}_{1}(R^{*})$ where $R^{*}$ is the ground-truth reward.
+It follows that $\mbox{Var}(\pi^{*}_{1}(\widehat{R}_{\mathcal{D}}))=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{D}|}$, $\mbox{Var}(\pi^{*}_{1}(\widehat{R}_{\mathcal{P}}))=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{P}|}$ and
+$\mbox{Var}(\pi^{\rm AIHF}_{1})=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{D}|+|\mathcal{P}|}<\min\{\mbox{Var}(\pi^{*}_{1}(\widehat{R}_{\mathcal{D}})),\mbox{Var}(\pi^{*}_{1}(\widehat{R}_{\mathcal{P}}))\}$
+Hence, the AIHF policy estimate has less variance than either the policy obtained from only demonstrations or preferences.
+To further illustrate, suppose the ground truth is $R^{*}(\tau_{1})=R^{*}(\tau_{2})$ and in the dataset there are more demonstrations than preferences, i.e. $|\mathcal{D}|\gg|\mathcal{P}|$:
+#$\#\{\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}=\#\{\tau_{2}~{}\mbox{in}~{}\mathcal{D}\}=50,\quad\#\{\tau_{1}\succ\tau_{2}~{}\mbox{in}~{}\mathcal{P}\}$ $=6,\quad\#\{\tau_{2}\succ\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}=4.$
+With the given data, $\pi^{\rm SFT}_{1}=\pi^{*}_{1}(\widehat{R}_{\mathcal{D}})=\frac{\#\{\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}}{|\mathcal{D}|}=\frac{50}{100}$ and the solution to (32) yields
+$\pi^{*}_{1}(\widehat{R}_{\mathcal{P}})=\frac{\exp\widehat{R}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{R}_{\mathcal{P}}(\tau_{1})+\exp\widehat{R}_{\mathcal{P}}(\tau_{2})}=\frac{6}{10}.$
+Hence, $\pi_{1}^{\rm AIHF}=\frac{100}{10+100}\pi^{*}_{1}(\widehat{R}_{\mathcal{D}})+\frac{10}{10+100}\pi^{*}_{1}(\widehat{R}_{\mathcal{P}})=\frac{56}{110}$. It follows from (33) that:
+$\pi^{\rm RLHF}_{1}$ $=\frac{\pi^{\rm SFT}_{1}\exp\widehat{R}_{\mathcal{P}}(\tau_{1})}{\pi^{\rm SFT}_{1}\exp\widehat{R}_{\mathcal{P}}(\tau_{1})+\pi^{\rm SFT}_{2}\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}$
+$=\frac{\exp\widehat{R}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{R}_{\mathcal{P}}(\tau_{1})+\exp\widehat{R}_{\mathcal{P}}(\tau_{2})}=\frac{6}{10}.$
+In this example, the RLHF policy estimator is the furthest away from ground-truth, because it does not correctly use the information provided by the demonstration data which in this case happens by chance to be correct $\pi^{\rm SFT}_{1}=\pi^{*}_{1}(\widehat{R}_{\mathcal{P}})=\frac{1}{2}$.
+As a second example, again suppose the ground truth is $R^{*}(\tau_{1})=R^{*}(\tau_{2})$ and in the dataset there are more preferences than demonstrations, $|\mathcal{P}|\gg|\mathcal{D}|$:
+{#$\{\#\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}=6,\quad\{\#\tau_{2}~{}\mbox{in}~{}\mathcal{D}\}=4,\quad\{\#\tau_{1}\succ\tau_{2}~{}\mbox{in}~{}\mathcal{P}\}=\{\#\tau_{2}\succ\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}=50.$ (37)
+In this case, $\pi^{\rm SFT}_{1}=\pi^{*}_{1}(\widehat{R}_{\mathcal{D}})=\frac{6}{10}$ and the solution to (32) yields
+$\pi^{*}_{1}(\widehat{R}_{\mathcal{P}})=\frac{\exp\widehat{R}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{R}_{\mathcal{P}}(\tau_{1})+\exp\widehat{R}_{\mathcal{P}}(\tau_{2})}=\frac{50}{100}.$
+It follows from (33) that:
+$\pi^{\rm RLHF}_{1}$ $=\frac{\pi^{\rm SFT}_{1}\exp\widehat{R}_{\mathcal{P}}(\tau_{1})}{\pi^{\rm SFT}_{1}\exp\widehat{R}_{\mathcal{P}}(\tau_{1})+\pi^{\rm SFT}_{2}\exp\widehat{R}_{\mathcal{P}}(\tau_{2})}$
+$=\frac{\pi_{1}^{SFT}}{\pi_{1}^{SFT}+\pi_{2}^{SFT}}=\frac{6}{10}.$
+Hence, $\pi_{1}^{\rm AIHF}=\frac{10}{10+100}\pi^{*}_{1}(\widehat{R}_{\mathcal{D}})+\frac{100}{10+100}\pi^{*}_{1}(\widehat{R}_{\mathcal{P}})=\frac{56}{110}$.
+In this example, the RLHF policy estimator is again farthest from ground-truth, because it does not correctly dismiss the information provided by the demonstration data which is less informative than preference.
+Example 2: Let us use an illustrative example to show that RLHF method will result in significant data under-utilization when the demonstration coverage is limited.
+With $\beta=1$, assume that there are 50 actions, i.e.
+$A=\{1,2,\dots,50\}$ and each with a ground-truth reward defined by $R^{*}(\tau_{i})=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{(\frac{i}{50}-\mu)^{2}}{2\sigma^{2}}}$, where $\mu=0.5$ and $\sigma=2$.
+Assume we can sample demonstration and preference from the ground truth reward distribution: demonstrations are sampled from the multinomial distribution, while preferences are sampled from the BTL model.
+In an extreme scenario, let demonstrations only include actions 1 through 45, i.e.
+$\mathcal{D}\cap\{45,46,\dots,50\}=\varnothing$, while preferences have full coverage across all actions.
+In the subsequent experiment, we initially sample 2000 demonstrations using the multinomial distribution $\pi^{*}_{i}=\frac{\exp R^{*}_{i}}{\sum^{j=45}_{j=1}\exp R^{*}_{j}}$, and obtain 200 preferences for each preference pair with $P(i\succ j)=\frac{\exp R^{*}_{i}}{\exp R^{*}_{j}+\exp R^{*}_{i}}$.
+We then calculate the RLHF and AIHF policies as in in (33) and (36) to obtain the result depicted in Figure 9:
+Figure 9: The optimal policy of RLHF, SFT, AIHF, and Ground-truth distribution.
+The left region of the red dotted line is included in the demonstration, while the right region is uncovered.
+We report the results with 100 random repeats.
+From the result shown in Figure 9, we demonstrate that both SFT and RLHF transfer the weight from uncovered action to covered actions when demonstration coverage is limited, as indicated by $\pi_{SFT}(\tau_{i})=0,\tau_{i}\in\{45,46,\dots,50\}$.
+Consequently, the weight of covered actions is significantly higher than the ground truth.
+However, this issue does not occur when jointly optimizing the demonstration and preference in the AIHF method.
+A.3 Proofs
+A.3.1 Useful Lemmas
+Assumption 1 (Ergodicity)
+For any policy $\pi$, assume the Markov chain with transition kernel $\mathcal{P}$ is irreducible and aperiodic under policy $\pi$. Then there exist constants $\kappa>0$ and $\rho\in(0,1)$ such that
+$\sup_{s\in\mathcal{S}}~{}\|\mathbb{P}(s_{t}\in\cdot|s_{0}=s,\pi)-\mu_{\pi}(\cdot)\|_{TV}\leq\kappa\rho^{t},\quad\forall~{}t\geq 0$
+where $\|\cdot\|_{TV}$ is the total variation (TV) norm; $\mu_{\pi}$ is the stationary state distribution under $\pi$.
+Assumption 1 assumes the Markov chain mixes at a geometric rate.
+It is a common assumption in the literature of RL, which holds for any time-homogeneous Markov chain with finite-state space or any uniformly ergodic Markov chain with general-state space.
+Assumption 2
+For any $s\in\mathcal{S}$, $a\in\mathcal{A}$ and any reward parameter $\theta$, the following holds:
+$\big{\|}\nabla_{\theta}r(s,a;\theta)\big{\|}$ $\leq L_{r},$ (38a)
+$\big{\|}\nabla_{\theta}r(s,a;\theta_{1})-\nabla_{\theta}r(s,a;\theta_{2})\big{\|}$ $\leq L_{g}\|\theta_{1}-\theta_{2}\|$ (38b)
+where $L_{r}$ and $L_{g}$ are positive constants.
+2, we next provide the following Lipschitz properties:
+Lemma A.1
+Suppose Assumptions 1 - 2 hold. For any reward parameter $\theta_{1}$ and $\theta_{2}$, the following results hold:
+$|Q^{soft}_{r_{\theta_{1}},\pi_{\theta_{1}}}(s,a)-Q^{soft}_{r_{\theta_{2}},\pi_{\theta_{2}}}(s,a)|$ $\leq L_{q}\|\theta_{1}-\theta_{2}\|,\quad\forall s\in\mathcal{S},a\in\mathcal{A}$ (39a)
+$\|\nabla L(\theta_{1})-\nabla L(\theta_{2})\|$ $\leq L_{c}\|\theta_{1}-\theta_{2}\|$ (39b)
+where $Q^{soft}_{r_{\theta},\pi_{\theta}}(\cdot,\cdot)$ denotes the soft Q-function under the reward function $r(\cdot,\cdot;\theta)$ and the policy $\pi_{\theta}$.
+The positive constants $L_{q}$ and $L_{c}$ are defined in Appendix A.3.3.
+A.3.2 Proof of Lemma 4.1
+Before we proceed to the proof of Lemma 4.1, we have the following remark.
+Remark: The KL-regularized MDP problem described by (5b) and (7) has a closed-form solution:
+$\pi_{\theta}(a|s)=\frac{\pi^{0}(a|s)\exp(Q_{\theta}(s,a)/\beta)}{\sum_{\tilde{a}}\pi^{0}(\tilde{a}|s)\exp(Q_{\theta}(s,a)/\beta)},$ (40)
+where the corresponding value function $V_{\theta}$ and the Q-function $Q_{\theta}$ are defined as below:
+$V_{\theta}(s):=\mathbb{E}_{\tau\sim\pi_{\theta}}\bigg{[}R(\tau;\theta)-\beta\sum_{t=0}^{\infty}\gamma^{t}{D}_{\rm KL}\Big{(}\pi(\cdot|s_{t})\|\pi^{0}(\cdot|s_{t})\Big{)}\bigg{|}s_{0}=s\bigg{]}$ (41a)
+$Q_{\theta}(s,a):=r(s,a;\theta)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}\big{[}V_{\theta}(s^{\prime})\big{]}.$ (41b)
+Further, assuming that $T=1$, i.e., $\tau=(s_{0},a_{0})$, and considering the LLM alignment problem as a sequence-level training problem (this is a popular simplification in language models, see, e.g., Rafailov et al.
+(2023)), the closed-form expression of $\pi_{\theta}$ in (40) can be reduced to:
+$\pi_{\theta}(a|s)=\frac{\pi^{0}(a|s)\exp(\frac{1}{\beta}r(s,a;\theta))}{\sum_{a\in A}\big{(}\pi^{0}(a|s)\exp(\frac{1}{\beta}r(s,a;\theta))\big{)}}.$ (42)
+Here, under a reward parameter $\theta$ and the corresponding optimal policy $\pi_{\theta}$ of (40).
 Moreover, under a fixed reward parameter $\theta$, we have defined the optimal policy $\pi_{\theta}$ as below:
-$\pi_{\theta}:=\arg\max_{\pi}~{}\mathbb{E}_{\tau^{\rm A}\sim\pi}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\bigg{(}r(s_{t},a_{t};\theta)+\beta\mathcal{D}_{KL}\Big{(}\pi(\cdot|s_{t})||\pi^{0}(\cdot|s_{t})\Big{)}\bigg{)}\bigg{]}.$
-According to Uehara et al. [2023], the optimal policy $\pi_{\theta}$ of (10) has the closed form expression as below:
+$\pi_{\theta}:=\arg\max_{\pi}~{}\mathbb{E}_{\tau^{\rm A}\sim\pi}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\bigg{(}r(s_{t},a_{t};\theta)-\beta\mathcal{D}_{KL}\Big{(}\pi(\cdot|s_{t})||\pi^{0}(\cdot|s_{t})\Big{)}\bigg{)}\bigg{]}.$
+According to Uehara et al. (2023), the optimal policy $\pi_{\theta}$ of (7) has the closed form expression as below:
 $\pi_{\theta}(a|s)=\frac{\pi^{0}(a|s)\exp\big{(}\frac{Q_{\theta}(s,a;\theta)}{\beta}\big{)}}{\sum_{\tilde{a}\in\mathcal{A}}\pi^{0}(\tilde{a}|s)\exp\big{(}\frac{Q_{\theta}(s,\tilde{a};\theta)}{\beta}\big{)}},\quad\forall s\in\mathcal{S},a\in\mathcal{A}.$ (43)
@@
 $\nabla L_{1}(\theta)=\frac{1}{\beta}\mathbb{E}_{\tau^{\rm E}\sim\pi^{\rm E}}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta)\bigg{]}-\frac{1}{\beta}\mathbb{E}_{\tau^{\rm A}\sim\pi_{\theta}}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta)\bigg{]}$ (48)
-A.2 Experiment Setup and Additional Result
-A.2.1 MuJoCo Tasks
-In MuJoCo, we consider several robotic control tasks with continuous action space.
-We evaluate the performance of our proposed algorithm on aligning robot behaviors with provided demonstrations and preference data.
-After the robot training stage, we leverage the ground-truth reward function from the environment to evaluate the performance.
-Data.
-Following the similar data generation pipeline in Brown et al.
-[2019], we generate the expert demonstrations and preference dataset as follows.
-We first train an expert agent by leveraging the ground-truth reward function and the popular Soft Actor-Critic (SAC) algorithm Haarnoja et al.
-[2018], which is developed to solve policy optimization problems with continuous action space.
-During the training process, we save the policy checkpoints and collect $30$k samples from each checkpoint.
-To achieve precise control of dataset quality, we categorize the data collected into three different classes: low-, medium-, and high-quality datasets according to the performance of the checkpoints.
-Then we combine the low- and medium-quality data as the preference dataset and use high-quality as demonstration data.
-Other details are given in Appendix A.2.
-Settings.
-We set $L_{1}(\cdot)$ using the likelihood objective (9), $L_{2}(\cdot)$ as the BTL objective (11), and $L_{3}$ as KL-regularized objective given in (10).
-We implement the RLHF Brown et al.
-[2019] using the following 3 steps: (i) utilize the demonstration dataset to pre-train the policy using behavior cloning (BC) Pomerleau [1988]; (ii) conduct reward learning using preference data; (iii) implement the SAC for policy fine-tuning.
-In the proposed AIHF, the policy $\pi^{0}$ is initiated with a random policy, and the policy optimization problem is also solved using SAC.
-In all Mujoco experiments, we assess the performance of benchmark algorithms on Hopper, HalfCheetah, and Walker2d from OpenAI Gym.
-To ensure a fair comparison, we utilize an open-source implementation of SAC [^5] https://github.com/openai/spinningup as the foundational RL algorithm for all methods.
-Regarding the choice of the penalty function $c(\cdot)$, since our proposed pipeline, AIHF , initiates with a random policy where $\pi(a|s)$ follows a uniform distribution, the kl penalty function is reduced to an entropy function.
-In SAC, both the policy network and Q network are (64, 64) MLPs with ReLU activation function, and the step size is set to $3*10^{-3}$, we parameterize the reward function by a (64, 64) MLPs with ReLU activation function.
-For the reward network, we use Adam as the optimizer, and the step size is set to be $1*10^{-4}$.
-The quality of the preference dataset and demonstration dataset are listed as follows Tab. 1:
-Task Dataset Non-prefer Data Prefer Data Demonstration Data
-Hopper-v2 $2345.20\pm 329.93$ $3024.63\pm 40.52$ $3559.61\pm 73.12$
-HalfCheetah-v2 $7226.37\pm 126.88$ $9434.42\pm 1315.13$ $11635.42\pm 236.51$
-Walker2d-v2 $3952.60\pm 444.45$ $5091.71\pm 291.73$ $5453.41\pm 71.07$
-Table 1: The quality of preference and demonstration.
-In the following sections, we include additional details about language alignment experiments.
-A.2.2 Sentiment-Controlled Generation
-Dataset Generation: In the IMDb sentiment completion task, we generate the demonstrations and preference datasets using the following procedure.
-Initially, we train a Language Model by employing the ground-truth reward function distilbert-imdb and the Proximal Policy Optimization (PPO) algorithm on 30% of the training dataset for IMDb.
-Throughout the training process, we save the policy checkpoint every 500 PPO steps.
-Subsequently, we select an additional 40% of the training dataset and generate a response for each prompt for each checkpoint.
-According to the evaluation score of each generation, we categorize the data collected into different classes: low-, medium-, and high-quality datasets, then we combine low-quality and medium-quality as preference datasets, and use high-quality as demonstration datasets.
-We show the average score of each dataset in 3,
-Training: After acquiring the preference and demonstration datasets, we train the proposed algorithm AIHF and baselines on the remaining 30% of prompts from the training dataset.
-We evaluate the performance of each algorithm using the test datasets for IMDb and HH, along with their corresponding ground truth reward functions.
-For the GPU resources, we use 8$\times$ A10 for all the experiments.
-Please refer to Table 2 for detailed hyperparameters and model initialization setting, and Table 3 for the quality of the preference and demonstration datasets we synthesize.
-Results: Policy Quality.
-We find that the proposed approach works well when either preference or demonstration data, or both, are limited.
-From the 6, we see that by using the same amount of data (10k preference, 10k demonstration), AIHF-based algorithms achieve faster convergence than their RLHF and DPO counterparts.
-Next, in Fig.
-7, we show that policy learned using AIHF  generates higher-quality continuations (measured by higher reward scores), as compared with those generated by RLHF, across a number of different combinations of preference and demonstration data size.
-and the size of preference data can significantly affect the quality of the estimated policy of RLHF, while it only affects the convergence speed of AIHF.
-Figure 6: IMDb Sentiment-controlled Generation.
-Left: AIHF  vs RLHF; Right: AIHF-DPO vs DPO; These represent limited data regimes with 1k preference, and 1k demonstration.
-Results are averaged over 3 independent runs.
-The RLHF and DPO curves are initialized from the SFT policy; the AIHF-based algorithms are initialized from a random policy.
-The achieved reward is compared against the # of PPO (resp, DPO) steps performed (for AIHF each policy alignment performs $20$ steps of PPO).
-(a) RLHF (left) vs AIHF (right); Fix demonstration number to 1000, gradually increase preference number.
-(b) RLHF (left) vs AIHF  (right); Fix preference number to 3k, gradually increase demonstration number.
-Figure 7: IMDb Sentiment-controlled Generation RLHF vs AIHF ; (a) different demonstration data size; (b) preference data.
-Results: Distribution Mismatch.
-We show that AIHF  is able to alleviate the distribution mismatch between the generated trajectories by the policy, and the data that the learned reward model is able to rank.
-To evaluate the extend of such mismatch, we use the following three steps: (1) use 1k preference, 1k demonstration to train policy and reward model for RLHF and AIHF ; (2) for a given set of prompts from test dataset, use RLHF and AIHF  to perform generation; (3) use the trained reward models to rate the generation; (4) compare with the score generated by the ground truth reward lvwerra/distilbert-imdb.
-Fig.
-8 illustrates that the reward score distribution produced by AIHF  aligns closely with that of the ground truth reward, whereas that generated by RLHF exhibits a poor match.
-These results show that the reward model learned by AIHF is able to correctly evaluate the generation produced by the final policy.
-Figure 8: Comparison of the distribution of reward score generated by the trained reward models, and the ground truth reward model. RLHF vs ground truth (left); AIHF vs ground truth (right).
-hyperparameters task IMDb Secontrolled Generation H&H Alignment
-policy model GPT2 PYTHIA-160M & PYTHIA-1B
-reward model GPT2 PYTHIA-1.4B
-reward learning rate 1e-5 1e-5
-policy learning rate 1.41e-5 1e-6
-reward batch size 128 8
-policy batch size 256 4
-sample temperature 1.00 1.00
-policy update times per iteration 20 6000
-reward update times per iteration 12 20
-Table 2: Hyperparameters of AIHF. H&H refers to Helpful and Harmless.
-Task Dataset Non-prefer Data Prefer Data Demonstration Data
-IMDb $0.27\pm 0.51$ $1.13\pm 0.42$ $2.23\pm 0.13$
-Helpful and Harmless $-1.61\pm 0.24$ $-1.12\pm 0.17$ $-0.82\pm 0.05$
-Table 3: The quality of preference and demonstration.
-From Fig.9, our proposed algorithm AIHF  could obtain higher rewards than baseline methods in the IMDb setting for almost all KL values.
-Although AIHF  might get a low score from the ground truth reward model in the earlier step, AIHF  would get a higher reward with more iteration and optimization steps.
-This indicates that with the mix of demonstration data and preference data, we could prevent the policy from known issues of reward hacking, especially when the policy learned more human-aligned features beyond base models (high KL value).
-Moreover, AIHF  is persistent in the number of preference data, presenting that AIHF  could still gain benefit from the limited preference data in more optimization steps as long as the demonstration data is high quality enough.
-Figure 9: The frontier of expected reward vs KL to the reference policy in IMDB dataset.
-fix the demonstration number to 3k Left: Using 1k preference; Middle: Using 2k preferences; Right: Using 3k preference
-In Fig.10, we show the converge curves of AIHF-DPO vs vanilla DPO in different combinations of preference and demonstration data size.
-Moreover, we observe that AIHF-DPO is stable over all different settings, while DPO is unstable and its performance slowly decreases over time.
-Figure 10: DPO vs AIHF-DPO Left: 1k Demonstration 1k Preference, Middle:1k Demonstration 2k Preference Right:1k Demonstration 3k Preference.
-A.2.3 Helpfulness-Controlled Generation
-Results: Reward Distribution.
-Further, in Fig.
-11, we show the overall reward distribution of the continuation, we can observe the distribution of AIHF and RLHF have some overlap in low-quality continuation, however, AIHF can generate more high-quality continuations compared to RLHF, which shows that joint optimization can more effectively align the policy model with the demonstration distribution.
-Figure 11: The Reward Distribution of Helpfulness-controlled Generation.
-Left: Result on 160m model, Right: Results on 1B model, This figure reports the reward distribution of generation evaluated by PKU-Alignment/beaver-7b-v3.0-reward for AIHF and RLHF.
+A.3.3 Proof of Lemma A.1
+To prove Lemma A.1, we will prove the equality (39a) and the equality (39b) respectively. The constants $L_{q}$ and $L_{c}$ in Lemma A.1 has the expression:
+$L_{q}:=\frac{L_{r}}{1-\gamma},\quad L_{c}:=\frac{2L_{q}L_{r}C_{d}\sqrt{|\mathcal{S}|\cdot|\mathcal{A}|}}{1-\gamma}+\frac{2L_{g}}{1-\gamma}.$
+A.3.4 Proof of Inequality (39a)
+In this subsection, we prove the inequality (39a) in Lemma A.1.
+We show that $Q^{\text{soft}}_{r_{\theta},\pi_{\theta}}$ has a bounded gradient with respect to any reward parameter $\theta$, then the inequality (39a) holds due to the mean value theorem.
+According to the soft Bellman equation, we have shown the explicit expression of $\nabla_{\theta}Q^{\text{soft}}_{r_{\theta},\pi_{\theta}}(s,a)$ for any $s\in\mathcal{S}$ and $a\in\mathcal{A}$.
+Using this expression, we have the following series of relations:
+$\|\nabla_{\theta}Q^{\text{soft}}_{r_{\theta},\pi_{\theta}}(s,a)\|$ $=\bigg{\|}\mathbb{E}_{a_{0}\sim\pi_{\theta}(\cdot|s_{0}),s_{1}\sim P(\cdot|s_{0},a_{0})}\Big{[}\nabla_{\theta}r(s_{0},a_{0};\theta)+\gamma\nabla_{\theta}V_{\theta}(s_{1})\Big{]}\bigg{\|}$
+=$\overset{(i)}{=}\bigg{\|}\mathbb{E}_{\tau\sim\pi_{\theta}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta)~{}\bigg{|}(s_{0},a_{0})=(s,a)\bigg{]}\bigg{\|}$
+$\overset{(ii)}{\leq}\mathbb{E}_{\tau\sim\pi_{\theta}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\bigg{\|}\nabla_{\theta}r(s_{t},a_{t};\theta)\bigg{\|}~{}\bigg{|}(s_{0},a_{0})=(s,a)\bigg{]}$
+$\overset{(iii)}{\leq}\mathbb{E}_{\tau\sim\pi_{\theta}}\bigg{[}\sum_{t\geq 0}\gamma^{t}L_{r}~{}\bigg{|}(s_{0},a_{0})=(s,a)\bigg{]}$
+$=\frac{L_{r}}{1-\gamma}$ (49)
+where (i) is from the equality (47) in the proof of Lemma A.1, (ii) follows Jensen’s inequality and (iii) follows the inequality (38a) in Assumption 2.
+To complete this proof, we use the mean value theorem to show that
+$|Q^{\text{soft}}_{r_{\theta_{1}},\pi_{\theta_{1}}}(s,a)-Q^{\text{soft}}_{r_{\theta_{2}},\pi_{\theta_{2}}}(s,a)|\leq\|\max_{\theta}\nabla_{\theta}Q^{\text{soft}}_{r_{\theta},\pi_{\theta}}(s,a)\|\cdot\|\theta_{1}-\theta_{2}\|\leq L_{q}\|\theta_{1}-\theta_{2}\|$ (50)
+where the last inequality follows (49) and we denote $L_{q}:=\frac{L_{r}}{1-\gamma}$. Therefore, we have proved the Lipschitz continuous inequality in (39a).
+A.3.5 Proof of Inequality (39b)
+In this section, we prove the inequality (39b) in Lemma A.1.
+According to Lemma A.1, the gradient $\nabla L_{1}(\theta)$ is expressed as:
+$\nabla L_{1}(\theta)=\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta)\bigg{]}-\mathbb{E}_{\tau\sim\pi_{\theta}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta)\bigg{]}.$ (51)
+Using the above relation, we have
+$\|\nabla L_{1}(\theta_{1})-\nabla L_{1}(\theta_{2})\|$
+=$\overset{(i)}{=}\bigg{\|}\bigg{(}\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}-\mathbb{E}_{\tau\sim\pi_{\theta_{1}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}\bigg{)}-$
+$\quad\quad\bigg{(}\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\bigg{]}-\mathbb{E}_{\tau\sim\pi_{\theta_{2}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\bigg{]}\bigg{)}\bigg{\|}$
+$\leq\underbrace{\bigg{\|}\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}-\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\bigg{]}\bigg{\|}}_{\rm:=term~{}A}+$
+$\quad\quad\underbrace{\bigg{\|}\mathbb{E}_{\tau\sim\pi_{\theta_{1}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}-\mathbb{E}_{\tau\sim\pi_{\theta_{2}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\bigg{]}\bigg{\|}}_{\rm:=term~{}B}$ (52)
+where (i) follows the exact gradient expression in equation (51). Then we separately analyze term A and term B in (52).
+For term A, it follows that
+$\bigg{\|}\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}-\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\bigg{]}\bigg{\|}$
+$\overset{(i)}{\leq}\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\big{\|}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})-\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\big{\|}\bigg{]}$
+$\overset{(ii)}{\leq}\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}L_{g}\|\theta_{1}-\theta_{2}\|\bigg{]}$
+$=\frac{L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|$ (53)
+where (i) follows Jensen’s inequality and (ii) is from (38b) in Assumption 2.
+For the term B, it holds that
+$\bigg{\|}\mathbb{E}_{\tau\sim\pi_{\theta_{1}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}-\mathbb{E}_{\tau\sim\pi_{\theta_{2}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\bigg{]}\bigg{\|}$
+$\overset{(i)}{\leq}\bigg{\|}\mathbb{E}_{\tau\sim\pi_{\theta_{1}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}-\mathbb{E}_{\tau\sim\pi_{\theta_{2}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}\bigg{\|}$
+$\quad+\bigg{\|}\mathbb{E}_{\tau\sim\pi_{\theta_{2}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}-\mathbb{E}_{\tau\sim\pi_{\theta_{2}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\bigg{]}\bigg{\|}$
+$\overset{(ii)}{\leq}\frac{1}{1-\gamma}\bigg{\|}\mathbb{E}_{(s,a)\sim d(\cdot,\cdot;\pi_{\theta_{1}})}\bigg{[}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}-\mathbb{E}_{(s,a)\sim d(\cdot,\cdot;\pi_{\theta_{2}})}\bigg{[}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}\bigg{\|}$
+$\quad+\mathbb{E}_{\tau\sim\pi_{\theta_{2}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\bigg{\|}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})-\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\bigg{\|}\bigg{]}$
+$\overset{(iii)}{\leq}\frac{1}{1-\gamma}\bigg{\|}\sum_{s\in\mathcal{S},a\in\mathcal{A}}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{(}d(s,a;\pi_{\theta_{1}})-d(s,a;\pi_{\theta_{2}})\bigg{)}\bigg{\|}+\mathbb{E}_{\tau\sim\pi_{\theta_{2}}}\bigg{[}\sum_{k\geq 0}\gamma^{k}L_{g}\|\theta_{1}-\theta_{2}\|\bigg{]}$
+$\overset{(iv)}{\leq}\frac{2L_{r}}{1-\gamma}\|d(\cdot,\cdot;\pi_{\theta_{1}})-d(\cdot,\cdot;\pi_{\theta_{2}})\|_{TV}+\frac{L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|$ (54)
+where (i) follows the triangle inequality, (ii) is from Jensen’s inequality and the definition of the discounted state-action visitation measure $d(s,a;\pi):=(1-\gamma)\pi(a|s)\sum_{t\geq 0}\gamma^{t}\mathcal{P}^{\pi}(s_{t}=s|s_{0}\sim\eta)$; (iii) is from (38b) in Assumption 2;(iv) is from (38a) and the definition of the total variation norm.
+Consider the $L_{2}$ term:
+$L_{2}(\theta):=\mathbb{E}_{(\tau_{i},\tau_{w})\sim\pi^{P}}\left[\log\left(\sigma\left(R(\tau_{w};\theta)-R(\tau_{i};\theta)\right)\right)\right]$
+where $\sigma(x)$ is sigmoid function defined by: $\sigma(x)=\frac{1}{1+e^{-x}}$. We have
+$\nabla_{\theta}L_{2}(\theta)=\mathbb{E}_{(\tau_{i},\tau_{w})\sim\pi^{P}}\left[(1-\sigma(R(\tau_{w};\theta)-R(\tau_{i};\theta)))\cdot\left(\nabla_{\theta}R(\tau_{w};\theta)-\nabla_{\theta}R(\tau_{i};\theta)\right)\right].$
+$=\mathbb{E}_{(\tau_{i},\tau_{w})\sim\pi^{P}}\big{[}\left(\nabla_{\theta}R(\tau_{w};\theta)-\nabla_{\theta}R(\tau_{i};\theta)\right)-\sigma(R(\tau_{w};\theta)-R(\tau_{i};\theta))(\nabla_{\theta}R(\tau_{w};\theta)-\nabla_{\theta}R(\tau_{i};\theta))\big{]}$
+Using the triangle inequality, we obtain the following equation:
+$\|\nabla L_{2}(\tau_{w},\tau_{l};\theta_{1})-\nabla L_{2}(\tau_{w},\tau_{l};\theta_{2})\|$
+$\leq$ $\underbrace{\bigg{\|}\mathbb{E}_{(\tau_{i},\tau_{w})\sim\pi^{P}}\big{[}\left(\nabla_{\theta}R(\tau_{w};\theta_{1})-\nabla_{\theta}R(\tau_{i};\theta_{1})\right)-\left(\nabla_{\theta}R(\tau_{w};\theta_{2})-\nabla_{\theta}R(\tau_{i};\theta_{2})\right)\big{]}\bigg{\|}}_{\rm:=term~{}A}$
+$+\underbrace{\bigg{\|}\mathbb{E}_{(\tau_{i},\tau_{w})\sim\pi^{P}}\big{[}\sigma(R(\tau_{w};\theta_{1})-R(\tau_{i};\theta_{1}))(\nabla_{\theta}R(\tau_{w};\theta_{1})-\nabla_{\theta}R(\tau_{i};\theta_{1}))}$
+$\underbrace{-\sigma(R(\tau_{w};\theta_{2})-R(\tau_{i};\theta_{2}))(\nabla_{\theta}R(\tau_{w};\theta_{2})-\nabla_{\theta}R(\tau_{i};\theta_{2}))\big{]}\bigg{\|}}_{\rm:=term~{}B}$ (55)
+First we bound the term A of (A.3.5)
+termA=([$\rm term~{}A=\bigg{\|}\bigg{(}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{]}\bigg{)}-\bigg{(}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{2})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{]}\bigg{)}\bigg{\|}$
+$\leq\bigg{\|}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{2})\bigg{\|}+\bigg{\|}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{\|}$
+$\leq\frac{2L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|$ (56)
+Then we bounded term B of (A.3.5):
+$\rm term~{}B=$
+${=}\bigg{\|}\sigma\bigg{(}\bigg{[}\sum_{t\geq 0}\gamma^{t}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{]}\bigg{)}\bigg{(}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{]}\bigg{)}$
+$-\sigma\bigg{(}\bigg{[}\sum_{t\geq 0}\gamma^{t}r(s^{w}_{t},a^{w}_{t};\theta_{2})-\gamma^{t}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{]}\bigg{)}\bigg{(}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{2})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{]}\bigg{)}\bigg{\|}$
+$=\bigg{\|}\sigma\bigg{(}\sum_{t\geq 0}\gamma^{t}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{)}\bigg{(}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{)}$
+$-\sigma\bigg{(}\sum_{t\geq 0}\gamma^{t}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{)}\bigg{(}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{2})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{)}$
+$+\sigma\bigg{(}\sum_{t\geq 0}\gamma^{t}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{)}\bigg{(}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{2})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{)}$
+$\sigma\bigg{(}\sum_{t\geq 0}\gamma^{t}r(s^{w}_{t},a^{w}_{t};\theta_{2})-\gamma^{t}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{)}\bigg{(}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{2})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{)}\bigg{\|}$
+$\leq\bigg{\|}\sigma\bigg{(}\sum_{t\geq 0}\gamma^{t}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{)}\bigg{(}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{1})$
+$\quad+\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{2})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{)}\bigg{\|}+\bigg{\|}\bigg{(}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{)}$
+$\quad\bigg{[}\sigma\bigg{(}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{)}-\sigma\bigg{(}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{)}\bigg{]}\bigg{\|}$
+$\leq\frac{2L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|+\frac{L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|=\frac{3L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|$ (57)
+Plugging the inequalities (53), (54) to (52), it holds that
+$\|\nabla L(\theta_{1})-\nabla L(\theta_{2})\|$
+$\leq\frac{2L_{r}}{1-\gamma}\|d(\cdot,\cdot;\pi_{\theta_{1}})-d(\cdot,\cdot;\pi_{\theta_{2}})\|_{TV}+\frac{6L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|$
+$\overset{(i)}{\leq}\frac{2L_{r}C_{d}}{1-\gamma}\|Q^{\text{soft}}_{r_{\theta_{1}},\pi_{\theta_{1}}}-Q^{\text{soft}}_{r_{\theta_{2}},\pi_{\theta_{2}}}\|+\frac{6L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|$
+$\overset{(ii)}{\leq}\frac{2L_{r}C_{d}\sqrt{|\mathcal{S}|\cdot|\mathcal{A}|}}{1-\gamma}\|Q^{\text{soft}}_{r_{\theta_{1}},\pi_{\theta_{1}}}-Q^{\text{soft}}_{r_{\theta_{2}},\pi_{\theta_{2}}}\|_{\infty}+\frac{6L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|$
+$\overset{(iii)}{\leq}\bigg{(}\frac{2L_{q}L_{r}C_{d}\sqrt{|\mathcal{S}|\cdot|\mathcal{A}|}}{1-\gamma}+\frac{6L_{g}}{1-\gamma}\bigg{)}\|\theta_{1}-\theta_{2}\|.$ (58)
+Define the constant $L_{c}:=\frac{2L_{q}L_{r}C_{d}\sqrt{|\mathcal{S}|\cdot|\mathcal{A}|}}{1-\gamma}+\frac{5L_{g}}{1-\gamma}$, we have the following inequality:
+$\|\nabla L(\theta_{1})-\nabla L(\theta_{2})\|\leq L_{c}\|\theta_{1}-\theta_{2}\|.$
+Therefore, we complete the proof of the inequality (39b) in Lemma A.1.
+A.4 Proof of Theorem 4.1
+In this section, we prove (20a) and (20b) respectively, to show the convergence of the lower-level problem and the upper-level problem.
+A.4.1 Proof of (20a)
+In this proof, we first show the convergence of the lower-level variable $\{\pi_{k}\}_{k\geq 0}$.
+Recall that we approximate the optimal policy $\pi_{\theta_{k}}$ by $\pi_{k+1}$ at each iteration $k$.
+We first analyze the approximation error between $\pi_{\theta_{k}}$ and $\pi_{k+1}$ as follows.
+For any $s\in\mathcal{S}$ and $a\in\mathcal{A}$, we have the following relation:
+$\big{|}\log\big{(}\pi_{k+1}(a|s)\big{)}-\log\big{(}\pi_{\theta_{k}}(a|s)\big{)}\big{|}$
+=$\overset{(i)}{=}\bigg{|}\log\bigg{(}\frac{\pi^{0}(a|s)\exp\big{(}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s,a)\big{)}}{\sum_{\tilde{a}}\exp\pi^{0}(\tilde{a}|s)\big{(}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s,\tilde{a})\big{)}}\bigg{)}-\log\bigg{(}\frac{\pi^{0}(a|s)\exp\big{(}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,a)\big{)}}{\sum_{\tilde{a}}\pi^{0}(\tilde{a}|s)\exp\big{(}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,\tilde{a})\big{)}}\bigg{)}\bigg{|}$
+$\overset{(ii)}{\leq}\big{|}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s,a)-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,a)\big{|}+\bigg{|}\log\bigg{(}\sum_{\tilde{a}}\pi^{0}(\tilde{a}|s)\exp\big{(}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s,\tilde{a})\big{)}\bigg{)}$
+$\quad-\log\bigg{(}\sum_{\tilde{a}}\pi^{0}(\tilde{a}|s)\exp\big{(}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,\tilde{a})\big{)}\bigg{)}\bigg{|}$ (59)
+where (i) follows (40); (ii) is by the triangle inequality. We further analyze the second term in (59).
+We first denote the operator $\log(\|w\exp(v)\|_{1}):=\log(\|\sum_{\tilde{a}\in\mathcal{A}}w\exp(v_{\tilde{a}})\|_{1})$, where the vector $w,v\in\mathbb{R}^{|\mathcal{A}|}$ and $v=[v_{1},v_{2},\cdots,v_{|\mathcal{A}|}],w=[w_{1},w_{2},\cdots,w_{|\mathcal{A}|}]$.
+Then for any $v^{\prime},v^{\prime\prime}\in\mathbb{R}^{|\mathcal{A}|}$, we have the following relation:
+$\big{|}\log\big{(}\|w^{\prime}\exp(v^{\prime})\|_{1}\big{)}-\log\big{(}\|w^{\prime\prime}\exp(v^{\prime\prime})\|_{1}\big{)}$ =$\overset{(i)}{=}\big{\langle}v^{\prime}-v^{\prime\prime},\nabla_{v}\log\big{(}\|w\exp(v)\|_{1}\big{)}|_{v=v^{c}}\big{\rangle}$
+$\leq\|v^{\prime}-v^{\prime\prime}\|_{\infty}\cdot\|\nabla_{v}\log\big{(}\|w\exp(v)\|_{1}\big{)}|_{v=v^{c}}\|_{1}$
+=$\overset{(ii)}{=}\|v^{\prime}-v^{\prime\prime}\|_{\infty}$ (60)
+where (i) follows the mean value theorem and $v_{c}$ is a convex combination of $v^{\prime}$ and $v^{\prime\prime}$; (ii) follows the following equalities:
+$[\nabla_{v}\log\big{(}\|w\exp(v)\|_{1}\big{)}]_{i}=\frac{w_{i}\exp(v_{i})}{\sum_{1\leq a\leq|\mathcal{A}|}w_{a}\exp(v_{a})},\quad\|\nabla_{v}\log\big{(}\|w\exp(v)\|_{1}\big{)}\|_{1}=1,\quad\forall v\in\mathbb{R}^{|\mathcal{A}|}.$
+Through plugging (60) into (59), it holds that
+$\big{|}\log\big{(}\pi_{k+1}(a|s)\big{)}-\log\big{(}\pi_{\theta_{k}}(a|s)\big{)}\big{|}$
+$\leq\big{|}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s,a)-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,a)\big{|}+\max_{\tilde{a}\in\mathcal{A}}\big{|}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s,\tilde{a})-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,\tilde{a})\big{|}$ (61)
+Taking the infinity norm over $\mathbb{R}^{|\mathcal{S}|\cdot|\mathcal{A}|}$, the following result holds:
+$\|\log\pi_{k+1}-\log\pi_{\theta_{k}}\|_{\infty}\leq 2\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}\|_{\infty}$ (62)
+where $\|\log\pi_{k+1}-\log\pi_{\theta_{k}}\|_{\infty}=\max_{s\in\mathcal{S},a\in\mathcal{A}}|\log\pi_{k+1}(a|s)-\log\pi_{\theta_{k}}(a|s)|$ and $\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}\|_{\infty}=\max_{s\in\mathcal{S},a\in\mathcal{A}}|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s,a)-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,a)|$.
+Based on the inequality (62), we analyze $\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}\|_{\infty}$ to show the convergence of the policy estimates.
+It leads to the following analysis:
+$\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}\|_{\infty}$
+$=\|Q_{r_{\theta_{k}},\pi_{k}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}+Q^{\text{soft}}_{r_{\theta_{k-1}},\pi_{\theta_{k-1}}}-Q^{\text{soft}}_{r_{\theta_{k-1}},\pi_{\theta_{k-1}}}+Q_{r_{\theta_{k-1}},\pi_{k}}^{\text{soft}}-Q_{r_{\theta_{k-1}},\pi_{k}}^{\text{soft}}\|_{\infty}$
+$\leq\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}-Q^{\text{soft}}_{r_{\theta_{k-1}},\pi_{\theta_{k-1}}}\|_{\infty}+\|Q_{r_{\theta_{k-1}},\pi_{k}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k-1}},\pi_{\theta_{k-1}}}\|_{\infty}+\|Q_{r_{\theta_{k}},\pi_{k}}^{\text{soft}}-Q_{r_{\theta_{k-1}},\pi_{k}}^{\text{soft}}\|_{\infty}$
+$\overset{(i)}{\leq}L_{q}\|\theta_{k}-\theta_{k-1}\|+\|Q_{r_{\theta_{k-1}},\pi_{k}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k-1}},\pi_{\theta_{k-1}}}\|_{\infty}+\|Q_{r_{\theta_{k}},\pi_{k}}^{\text{soft}}-Q_{r_{\theta_{k-1}},\pi_{k}}^{\text{soft}}\|_{\infty}$
+$\overset{(ii)}{\leq}\|Q_{r_{\theta_{k-1}},\pi_{k}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k-1}},\pi_{\theta_{k-1}}}\|_{\infty}+2L_{q}\|\theta_{k}-\theta_{k-1}\|$ (63)
+where (i) is from (39a) in Lemma A.1; (ii) follows (39a). Based on (63), we further analyze the two terms in (63) as below.
+Recall we have the “soft” Bellman operator expressed as below:
+$\mathcal{T}_{\theta}(Q)(s,a)=r(s,a;\theta)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s^{\prime},a^{\prime})}\bigg{[}\log\left(\sum_{a^{\prime}}\pi^{0}(a^{\prime}|s^{\prime})\exp\big{(}Q(s^{\prime},a^{\prime})\big{)}\right)\bigg{]}$ (64)
+According to the soft Bellman operator, it holds that
+$Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k+1}}(s,a)$ $=r(s,a;\theta_{k})+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[V^{\text{soft}}_{r_{\theta_{k}},\pi_{k+1}}(s^{\prime})]$
+$=r(s,a;\theta_{k})+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a),a^{\prime}\sim\pi_{k+1}(\cdot|s^{\prime})}[-\frac{\log\pi_{k+1}(a^{\prime}|s^{\prime})}{\log\pi_{0}(a^{\prime}|s^{\prime})}+Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k+1}}(s^{\prime},a^{\prime})]$
+$\overset{(i)}{\geq}r(s,a;\theta_{k})+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a),a^{\prime}\sim\pi_{k+1}(\cdot|s^{\prime})}[-\frac{\log\pi_{k+1}(a^{\prime}|s^{\prime})}{\log\pi_{0}(a^{\prime}|s^{\prime})}+Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s^{\prime},a^{\prime})]$
+=$\overset{(ii)}{=}r(s,a;\theta_{k})+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}\left[\log\bigg{(}\sum_{a^{\prime}}\pi^{0}(a^{\prime}|s^{\prime})\exp\big{(}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s^{\prime},a^{\prime})\big{)}\bigg{)}\right]$
+=$\overset{(iii)}{=}\mathcal{T}_{\theta_{k}}(Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}})(s,a)$ (65)
+where (i) follows the policy improvement result(ii) follows the definition $\pi_{k+1}(a|s):=\frac{\pi^{0}(a|s)\exp\big{(}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s,a)\big{)}}{\sum_{\tilde{a}}\pi^{0}(\tilde{a|s})\exp\big{(}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s,\tilde{a})\big{)}}$ (iii) follows the definition of the soft Bellman operator in (64).
+For any $s\in\mathcal{S}$ and $a\in\mathcal{A}$, it holds that
+$0\overset{(i)}{\leq}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,a)-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k+1}}(s,a)\overset{(ii)}{\leq}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,a)-\mathcal{T}_{\theta_{k}}(Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}})(s,a)$ (66)
+where (i) is due to the fact that $\pi_{\theta_{k}}$ is the optimal policy under reward parameter $\theta_{k}$; (ii) is from (65).
+Hence, it further leads to
+$\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k+1}}\|_{\infty}$ $\overset{(i)}{\leq}\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}-\mathcal{T}_{\theta_{k}}(Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}})\|_{\infty}$
+=$\overset{(ii)}{=}\|\mathcal{T}_{\theta_{k}}(Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}})-\mathcal{T}_{\theta_{k}}(Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}})\|_{\infty}$
+$\overset{(iii)}{\leq}\gamma\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}\|_{\infty}$ (67)
+where (i) is from (66); (ii) is from the fixed-point property in (83); (iii) is from the contraction property in (82). Therefore, we have the following result:
+$\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}\|_{\infty}$
+$\overset{(i)}{\leq}\|Q_{r_{\theta_{k-1}},\pi_{k}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k-1}},\pi_{\theta_{k-1}}}\|_{\infty}+2L_{q}\|\theta_{k}-\theta_{k-1}\|$
+$\overset{(ii)}{\leq}\gamma\|Q_{r_{\theta_{k-1}},\pi_{k-1}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k-1}},\pi_{\theta_{k-1}}}\|_{\infty}+2L_{q}\|\theta_{k}-\theta_{k-1}\|$ (68)
+where (i) is from (63); (ii) is from (67).
+To show the convergence of the soft Q-function based on (68), we further analyze the error between the reward parameters $\theta_{k}$ and $\theta_{k-1}$.
+Recall in Alg.1, the updates in reward parameters (4):
+$\theta_{k}=\theta_{k-1}+\alpha g_{k-1}$
+where we denote $\tau=\{(s_{t},a_{t})\}_{t=0}^{\infty}$, $h(\theta,\tau):=\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta)$ and $g_{k-1}$ is the stochastic gradient estimator at iteration $k-1$.
+Here, $\tau^{E}_{k-1}$ denotes the trajectory sampled from the expert’s dataset $D$ at iteration $k-1$ and $\tau^{A}_{k-1}$ denotes the trajectory sampled from the agent’s policy $\pi_{k}$ at time $k-1$,$\tau_{w},\tau_{i}$ denote the trajectory sampled from the preference dataset.
+Then according to the inequality (38a) in Assumption 2, we could show that
+$\|g_{k-1}$ $\|\leq\|h(\theta_{k-1},\tau^{E}_{k-1})-h(\theta_{k-1},\tau^{A}_{k-1})\|+\|h(\theta_{k-1},\tau^{W}_{k-1})-h(\theta_{k-1},\tau^{L}_{k-1})\|$
+$\leq\frac{2L_{r}}{1-\gamma}+\frac{2L_{r}}{1-\gamma}=4L_{q}$ (69)
+where the last equality follows the fact that we have defined the constant $L_{q}:=\frac{L_{r}}{1-\gamma}$. Then we could further show that
+$\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}\|_{\infty}$
+$\overset{(i)}{\leq}\gamma\|Q_{r_{\theta_{k-1}},\pi_{k-1}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k-1}},\pi_{\theta_{k-1}}}\|_{\infty}+4L_{q}\|\theta_{k}-\theta_{k-1}\|$
+=$\overset{(ii)}{=}\gamma\|Q_{r_{\theta_{k-1}},\pi_{k-1}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k-1}},\pi_{\theta_{k-1}}}\|_{\infty}+4\alpha L_{q}\|g_{k-1}\|$
+$\overset{(iii)}{\leq}\gamma\|Q_{r_{\theta_{k-1}},\pi_{k-1}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k-1}},\pi_{\theta_{k-1}}}\|_{\infty}+8\alpha L_{q}^{2}$ (70)
+where (i) is from (68); (ii) follows the reward update scheme; (iii) is from (69).
+Summing the inequality (70) from $k=1$ to $k=K$, it holds that
+$\sum_{k=1}^{K}\|Q_{r_{\theta_{k}},\pi_{k}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}\|_{\infty}\leq\gamma\sum_{k=0}^{K-1}\|Q_{r_{\theta_{k}},\pi_{k}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}\|_{\infty}+8\alpha KL_{q}^{2}$ (71)
+Rearranging the inequality (71) and divided (71) by $K$ on both sides, it holds that
+$\frac{1-\gamma}{K}\sum_{k=1}^{K}\|Q_{r_{\theta_{k}},\pi_{k}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}\|_{\infty}\leq\frac{\gamma}{K}\bigg{(}\|Q_{r_{\theta_{0}},\pi_{0}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{0}},\pi_{\theta_{0}}}\|_{\infty}-\|Q_{r_{\theta_{K}},\pi_{K}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{K}},\pi_{\theta_{K}}}\|_{\infty}\bigg{)}+8\alpha L_{q}^{2}$ (72)
+Dividing the constant $1-\gamma$ on both sides of (72), it holds that
+$\frac{1}{K}\sum_{k=1}^{K}\|Q_{r_{\theta_{k}},\pi_{k}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}\|_{\infty}\leq\frac{\gamma C_{0}}{K(1-\gamma)}+\frac{8L_{q}^{2}}{1-\gamma}\alpha$
+where we denote $C_{0}:=\|Q_{r_{\theta_{0}},\pi_{0}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{0}},\pi_{\theta_{0}}}\|_{\infty}$. We could also write the inequality above as
+$\frac{1}{K}\sum_{k=0}^{K-1}\|Q_{r_{\theta_{k}},\pi_{k}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}\|_{\infty}$
+$\leq\frac{\gamma C_{0}}{K(1-\gamma)}+\frac{C_{0}}{K}-\frac{\|Q_{r_{\theta_{K}},\pi_{K}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{K}},\pi_{\theta_{K}}}\|_{\infty}}{K}+\frac{8L_{q}^{2}}{1-\gamma}\alpha$
+$\leq\frac{C_{0}}{K(1-\gamma)}+\frac{8L_{q}^{2}}{1-\gamma}\alpha.$
+Recall the stepsize is defined as $\alpha=\frac{\alpha_{0}}{K^{\sigma}}$ where $\sigma>0$. Then we have the following result:
+$\frac{1}{K}\sum_{k=0}^{K-1}\|Q_{r_{\theta_{k}},\pi_{k}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}\|_{\infty}=\mathcal{O}(K^{-1})+\mathcal{O}(K^{-\sigma}).$ (73)
+With the inequality (62), it follows that
+$\frac{1}{K}\sum_{k=0}^{K-1}\|\log\pi_{k+1}-\log\pi_{\theta_{k}}\|_{\infty}\leq\frac{2}{K}\sum_{k=0}^{K-1}\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}\|_{\infty}=\mathcal{O}(K^{-1})+\mathcal{O}(K^{-\sigma}).$
+Therefore, we complete the proof of (20a) in Theorem 4.1.
+A.4.2 Proof of (20b)
+In this part, we prove the convergence of reward parameters $\{\theta_{k}\}_{k\geq 0}$.
+We have the following result of the objective function $L(\theta)$:
+$L(\theta_{k+1})$ $\overset{(i)}{\geq}L(\theta_{k})+\langle\nabla L(\theta_{k}),\theta_{k+1}-\theta_{k}\rangle-\frac{L_{c}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$
+=$\overset{(ii)}{=}L(\theta_{k})+\alpha\langle\nabla L(\theta_{k}),g_{k}\rangle-\frac{L_{c}\alpha^{2}}{2}\|g_{k}\|^{2}$
+$=L(\theta_{k})+\alpha\langle\nabla L(\theta_{k}),g_{k}-\nabla L(\theta_{k})\rangle+\alpha\|\nabla L(\theta_{k})\|^{2}-\frac{L_{c}\alpha^{2}}{2}\|g_{k}\|^{2}$
+$\overset{(iii)}{\geq}L(\theta_{k})+\alpha\langle\nabla L(\theta_{k}),g_{k}-\nabla L(\theta_{k})\rangle+\alpha\|\nabla L(\theta_{k})\|^{2}-8L_{c}L_{q}^{2}\alpha^{2}$ (74)
+where (i) is from the Lipschitz smooth property in (39b) of Lemma A.1; (ii) follows the update scheme (4); (iii) is from constant bound in (69).
+Taking an expectation over the both sides of (74), it holds that
+$\mathbb{E}\left[L(\theta_{k+1})\right]$
+$\geq\mathbb{E}\left[L(\theta_{k})\right]+\alpha\mathbb{E}\bigg{[}\langle\nabla L(\theta_{k}),g_{k}-\nabla L(\theta_{k})\rangle\bigg{]}+\alpha\mathbb{E}\bigg{[}\|\nabla L(\theta_{k})\|^{2}\bigg{]}-8L_{c}L_{q}^{2}\alpha^{2}$
+$=\mathbb{E}\left[L(\theta_{k})\right]+\alpha\mathbb{E}\bigg{[}\langle\nabla L(\theta_{k}),\mathbb{E}\big{[}g_{k}-\nabla L(\theta_{k})\big{|}\theta_{k}]\rangle\bigg{]}+\alpha\mathbb{E}\bigg{[}\|\nabla L(\theta_{k})\|^{2}\bigg{]}-8L_{c}L_{q}^{2}\alpha^{2}$
+${=}\mathbb{E}\left[L(\theta_{k})\right]+\alpha\mathbb{E}\bigg{[}\bigg{\langle}\nabla L(\theta_{k}),\mathbb{E}_{\tau\sim\pi_{\theta_{k}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{k})\bigg{]}-\mathbb{E}_{\tau\sim\pi_{k+1}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{k})\bigg{]}$
+$+\mathbb{E}_{(\tau_{l}\prec\tau_{w})\sim\pi^{P}}\bigg{[}\sum_{t\geq 0}(1-\sigma(\gamma^{t}r(s^{w}_{t},a^{w}_{t};\theta_{k})-\gamma^{t}r(s^{l}_{t},a^{l}_{t};\theta_{k}))(\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{k})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{k}))\bigg{]}\bigg{\rangle}\bigg{]}$
+$+\alpha\mathbb{E}\bigg{[}\|\nabla L(\theta_{k})\|^{2}\bigg{]}-8L_{c}L_{q}^{2}\alpha^{2}$
+$\overset{(i)}{\geq}\mathbb{E}\left[L(\theta_{k})\right]-4\alpha L_{q}\underbrace{\mathbb{E}\bigg{[}\bigg{\|}\mathbb{E}_{\tau\sim\pi_{\theta_{k}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{k})\bigg{]}-\mathbb{E}_{\tau\sim\pi_{k+1}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{k})}_{\rm term~{}A}\bigg{]}\bigg{\|}$
+$+\alpha\mathbb{E}\bigg{[}\|\nabla L(\theta_{k})\|^{2}\bigg{]}-8L_{c}L_{q}^{2}\alpha^{2}$ (75)
+(i) is due to the fact that $\|\nabla L(\theta)\|\leq 4L_{q}$ and $\mathbb{E}[g_{k,2}-\nabla_{\theta}L_{2}(\theta_{k})|\theta_{k}]=0$.
+Then we further analyze the term A as below:
+$\mathbb{E}\left[\bigg{\|}\mathbb{E}_{\tau\sim\pi_{\theta_{k}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{k})\bigg{]}-\mathbb{E}_{\tau\sim\pi_{k+1}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{k})\bigg{]}\bigg{\|}\right]$
+=$\overset{(i)}{=}\mathbb{E}\left[\bigg{\|}\frac{1}{1-\gamma}\mathbb{E}_{(s,a)\sim d(\cdot,\cdot;\pi_{\theta_{k}})}\big{[}\nabla_{\theta}r(s,a;\theta_{k})\big{]}-\frac{1}{1-\gamma}\mathbb{E}_{(s,a)\sim d(\cdot,\cdot;\pi_{k+1})}\big{[}\nabla_{\theta}r(s,a;\theta_{k})\big{]}\bigg{\|}\right]$
+$\overset{(ii)}{\leq}\frac{2}{1-\gamma}\cdot\max_{s\in\mathcal{S},a\in\mathcal{A}}\|\nabla_{\theta}r(s,a;\theta_{k})\|\cdot\mathbb{E}\big{[}\|d(\cdot,\cdot;\pi_{\theta_{k}})-d(\cdot,\cdot;\pi_{k+1})\|_{TV}\big{]}$
+$\overset{(iii)}{\leq}\frac{2L_{r}}{1-\gamma}\mathbb{E}\big{[}\|d(\cdot,\cdot;\pi_{\theta_{k}})-d(\cdot,\cdot;\pi_{k+1})\|_{TV}\big{]}$
+$\overset{(iv)}{\leq}2L_{q}C_{d}\mathbb{E}\left[\|\log\frac{\pi^{0}(a|s)\exp Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,a)}{\sum_{\tilde{a}}\pi^{0}(\tilde{a}|s)\exp Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,\tilde{a})}-\log\frac{\pi^{0}(a|s)\exp Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k+1}}(s,a)}{\sum_{\tilde{a}}\pi^{0}(\tilde{a}|s)\exp Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k+1}}(s,\tilde{a})}\|\right]$
+$\overset{(v)}{\leq}2L_{q}C_{d}\mathbb{E}\left[\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}\|+\|\log\sum_{a}\pi^{0}(\tilde{a}|s)\exp Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,\tilde{a})-\log\sum_{a}\pi^{0}(\tilde{a}|s)\exp Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k+1}}(s,\tilde{a})\|\right]$
+$\overset{(vi)}{\leq}2L_{q}C_{d}\sqrt{|\mathcal{S}|\cdot|\mathcal{A}|}\mathbb{E}\left[\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}\|_{\infty}+\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}\|_{\infty}\right]$
+$=4L_{q}C_{d}\sqrt{|\mathcal{S}|\cdot|\mathcal{A}|}\mathbb{E}\left[\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}\|_{\infty}\right]$ (76)
+where (i) follows the definition $d(s,a;\pi)=(1-\gamma)\pi(a|s)\sum_{t\geq 0}\gamma^{t}\mathcal{P}^{\pi}(s_{t}=s|s_{0}\sim\eta)$; (ii) is due to distribution mismatch between two visitation measures; (iii) follows the inequality (38a) in Assumption 2; the inequality (iv) follows Lemma A.2 and the fact that $\pi_{\theta_{k}}(\cdot|s)\propto\pi^{0}(\cdot|s)\exp\big{(}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,\cdot)\big{)}$, $\pi_{k+1}(\cdot|s)\propto\pi^{0}(\cdot|s)\exp\big{(}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s,\cdot)\big{)}$ and the constant $L_{q}:=\frac{L_{r}}{1-\gamma}$; (v) follows the (60);(vi) follows the conversion between Frobenius norm and infinity norm.
+Through plugging the inequality (76) into (75), it leads to
+$\mathbb{E}\left[L(\theta_{k+1})\right]$
+$\geq\mathbb{E}\left[L(\theta_{k})\right]-2\alpha L_{q}\mathbb{E}\left[\big{\|}\mathbb{E}_{\tau\sim\pi_{\theta_{k}}}\big{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{k})\big{]}-\mathbb{E}_{\tau\sim\pi_{k+1}}\big{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{k})\big{]}\big{\|}\right]$
+$\quad+\alpha\mathbb{E}\bigg{[}\|\nabla L(\theta_{k})\|^{2}\bigg{]}-8L_{c}L_{q}^{2}\alpha^{2}$
+$\overset{(i)}{\geq}\mathbb{E}\left[L(\theta_{k})\right]-8\alpha C_{d}L_{q}^{2}\sqrt{|\mathcal{S}|\cdot|\mathcal{A}|}\mathbb{E}\left[\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}\|_{\infty}\right]+\alpha\mathbb{E}\bigg{[}\|\nabla L(\theta_{k})\|^{2}\bigg{]}-8L_{c}L_{q}^{2}\alpha^{2}$
+where (i) follows the inequality (76).
+Rearranging the inequality above and denote $C_{1}:=8C_{d}L_{q}^{2}\sqrt{|\mathcal{S}|\cdot|\mathcal{A}|}$, it holds that
+$\alpha\mathbb{E}\big{[}\|\nabla L(\theta_{k})\|^{2}\big{]}\leq 8L_{c}L_{q}^{2}\alpha^{2}+\alpha C_{1}\mathbb{E}\left[\|Q_{r_{\theta_{k}},\pi_{\theta_{k}}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}\|_{\infty}\right]+\mathbb{E}\big{[}L(\theta_{k+1})-L(\theta_{k})\big{]}$
+Summing the inequality above from $k=0$ to $K-1$ and dividing both sides by $\alpha K$, it holds that
+$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\|\nabla L(\theta_{k})\|^{2}\right]\leq 8L_{c}L_{q}^{2}\alpha+\frac{C_{1}}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\|Q_{r_{\theta_{k}},\pi_{\theta_{k}}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}\|_{\infty}\right]+\mathbb{E}\left[\frac{L(\theta_{K})-L(\theta_{0})}{K\alpha}\right]$ (77)
+Note that the log-likelihood function $L(\theta_{K})$ is negative and $L(\theta_{0})$ is a bounded constant. Then we could plug (73) into (77), it holds that
+$\frac{1}{K}\sum_{K=0}^{K-1}\mathbb{E}\left[\|\nabla L(\theta_{K})\|^{2}\right]=\mathcal{O}(K^{-\sigma})+\mathcal{O}(K^{-1})+\mathcal{O}(K^{-1+\sigma})$ (78)
+which completes the proof for the inequality (20b).
+A.5 Auxiliary Lemmas
+Lemma A.2
+((Xu et al., 2020, Lemma 3)) Consider the initialization distribution $\eta(\cdot)$ and transition kernel $\mathcal{P}(\cdot|s,a)$.
+Under $\eta(\cdot)$ and $\mathcal{P}(\cdot|s,a)$, denote $d_{w}(\cdot,\cdot)$ as the state-action visitation distribution of MDP with the Boltzman policy parameterized by parameter $w$.
+Suppose Assumption 1 holds, for all policy parameter $w$ and $w^{\prime}$, we have
+$\|d_{w}(\cdot,\cdot)-d_{w^{\prime}}(\cdot,\cdot)\|_{TV}\leq C_{d}\|w-w^{\prime}\|$ (79)
+where $C_{d}$ is a positive constant.
+Next, to facilitate analysis for KL-regularized MDPs, we introduce a “soft” Bellman optimality operator $\mathcal{T}:\mathbb{R}^{|\mathcal{S}|\times|\mathcal{A}|}\to\mathbb{R}^{|\mathcal{S}|\times|\mathcal{A}|}$ as follows:
+$\mathcal{T}(Q)(s,a):=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}\left[\max_{\pi(\cdot|s)}~{}\mathbb{E}_{a^{\prime}\sim\pi(\cdot|s^{\prime})}\left[Q(s^{\prime},a^{\prime})-\frac{\log\pi(a^{\prime}|s^{\prime})}{\log\pi^{0}(a^{\prime}|s^{\prime})}\right]\right].$ (80)
+In the following lemma, the properties of KL-regularized MDPs are characterized.
+Lemma A.3
+(The operator $\mathcal{T}$ as defined in (80) satisfies the properties below:
+  • $\mathcal{T}$ has the following closed-form expression:
+    $\mathcal{T}(Q)(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}\left[\log\left(\sum_{a^{\prime}}\pi^{0}(a^{\prime}|s^{\prime})\exp\big{(}Q(s^{\prime},a^{\prime})\big{)}\right)\right].$ (81)
+  • $\mathcal{T}$ is a $\gamma$-contraction in the $\ell_{\infty}$ norm, namely, for any $Q_{1},Q_{2}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{A}|}$, it holds that
+    $\|\mathcal{T}(Q_{1})-\mathcal{T}(Q_{2})\|_{\infty}\leq\gamma\|Q_{1}-Q_{2}\|_{\infty}.$ (82)
+  • Under a given reward function $r(\cdot,\cdot)$, the corresponding optimal soft $Q$-function $Q^{\text{soft}}_{r,\pi^{*}}$ is a unique fixed point of the operator $\mathcal{T}$, namely,
+    $\mathcal{T}(Q^{\text{soft}}_{r,\pi^{*}})=Q^{\text{soft}}_{r,\pi^{*}}$ (83)
+We refine its analysis as below.
+We first show that
+$\mathbb{E}_{a\sim\pi(\cdot|s)}\bigg{[}Q(s,a)-\frac{\log\pi(a|s)}{\log\pi^{0}(a|s)}\bigg{]}=\sum_{a}\pi(a|s)\log\left(\frac{\pi^{0}(a|s)\exp(Q(s,a))}{\pi(a|s)}\right)\overset{(i)}{\leq}\log\bigg{(}\sum_{a}\pi^{0}(a|s)\exp\big{(}Q(s,a)\big{)}\bigg{)}$ (84)
+where (i) is from Jensen’s inequality.
+Moreover, the equality between both sides of (i) holds when the policy $\pi$ has the expression $\pi(\cdot|s)\propto\pi^{0}(a|s)\exp(Q(s,\cdot))$.
+Therefore, through applying the inequality (84) to (80), it obtains that
+$\mathcal{T}(Q)(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}\left[\log\left(\sum_{a^{\prime}}\pi^{0}(a|s)\exp\big{(}Q(s^{\prime},a^{\prime})\big{)}\right)\right],$ (85)
+which proves the equality (81).
+We define $\|Q_{1}-Q_{2}\|_{\infty}:=\max_{s\in\mathcal{S},a\in\mathcal{A}}|Q_{1}(s,a)-Q_{2}(s,a)|$ and $\epsilon=\|Q_{1}-Q_{2}\|_{\infty}$.
+Then for any $s\in\mathcal{S}$ and $a\in\mathcal{A}$, it follows that
+$\log\left(\sum_{a}\pi^{0}(a|s)\exp\big{(}Q_{1}(s,a)\big{)}\right)$ $\leq\log\left(\sum_{a}\pi^{0}(a|s)\exp\big{(}Q_{2}(s,a)+\epsilon\big{)}\right)$
+$=\log\left(\exp(\epsilon)\sum_{a}\pi^{0}(a|s)\exp\big{(}Q_{2}(s,a)\big{)}\right)$
+$=\epsilon+\log\left(\sum_{a}\pi^{0}(a|s)\exp\big{(}Q_{2}(s,a)\big{)}\right)$
+Similarly, it is easy to obtain that $\log\left(\sum_{a}\pi^{0}(a|s)\exp\big{(}Q_{1}(s,a)\big{)}\right)\geq-\epsilon+\log\left(\sum_{a}\pi^{0}(a|s)\exp\big{(}Q_{2}(s,a)\big{)}\right)$.
+Hence, it leads to the contraction property that
+$\|\mathcal{T}(Q_{1})-\mathcal{T}(Q_{2})\|_{\infty}\leq\gamma\epsilon=\gamma\|Q_{1}-Q_{2}\|_{\infty}$ (86)
+which proves the contraction property (82). Moreover, we have
+$\mathcal{T}(Q^{\text{soft}}_{r,\pi^{*}})(s,a)\overset{(i)}{=}r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}\left[\log\left(\sum_{a^{\prime}}\pi^{0}(a^{\prime}|s^{\prime})\exp\big{(}Q^{\text{soft}}_{r,\pi^{*}}(s^{\prime},a^{\prime})\big{)}\right)\right]\overset{(ii)}{=}Q^{\text{soft}}_{r,\pi^{*}}(s,a)$ (87)
+where (i) follows the equality (85). Based on the definition of the soft Q-function $Q^{\text{soft}}_{r,\pi^{*}}$, we have
+$Q^{\text{soft}}_{r,\pi^{*}}(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}\big{[}\mathbb{E}_{a^{\prime}\sim\pi^{*}(\cdot|s^{\prime})}[-\frac{\log\pi^{*}(a^{\prime}|s^{\prime})}{\log\pi^{0}(a^{\prime}|s^{\prime})}+Q^{\text{soft}}_{r,\pi^{*}}(s^{\prime},a^{\prime})]\big{]}.$ (88)
+We prove the equality (ii) in (87) through combining (88) and the fact that the optimal soft policy has the closed form $\pi^{*}(\cdot|s)\propto\pi^{0}(\cdot|s^{\prime})\exp\big{(}Q^{\text{soft}}_{r,\pi^{*}}(s,\cdot)\big{)}$.
+Suppose two different fixed points of the soft Bellman operator exist, then it contradicts with the contraction property in (86).
+Hence, we proved the uniqueness of the optimal soft $Q$-function $Q^{\text{soft}}_{r,\pi^{*}}$.
+Moreover, the optimal soft $Q$-function $Q^{\text{soft}}_{r,\pi^{*}}$ is a fixed point to the soft Bellman operator $\mathcal{T}$ in (83).